### v2

# Space spanned by characteristic exponents

Abstract.
We prove several rigidity results on multiplier spectrum and length spectrum. For example, we show that for every non-exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$, the $\mathbb{Q}$-vector space generated by all the (finite) characteristic exponents of periodic points of $f$ has infinite dimension. This answers a stronger version of a question of Levy and Tucker. Our result can also be seen as a generalization of recent results of Ji-Xie and of Huguin which proved Milnor’s conjecture about rational maps having integer multipliers. We also get a characterization of postcritically finite maps by using its length spectrum. Finally as an application of our result, we get a new proof of the Zariski-dense orbit conjecture for endomorphisms on $(\mathbb{P}^{1})^{N},N\geq 1$.
2020 Mathematics Subject Classification:
Primary 37P35; Secondary 37F10, 11G35
1. Introduction
Let $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$. Our aim is to study the $\mathbb{Q}$-vector space spanned by the characteristic exponents of periodic points of a rational map on $\mathbb{P}^{1}(\mathbb{C})$ and prove some rigidity results.
1.1. Multiplier, length, and characteristic exponent
Let $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ be a periodic point of $f$ with exact period $n$. Define $n_{f}(z_{0}):=n$ be this period. We write $n(z_{0})$ for simplicity when the map $f$ is clear. The multiplier $\rho_{f}(z_{0})$ of $f$ at $z_{0}$ is defined to be the differential $df^{n}(z_{0})\in\mathbb{C}$. We write $\rho(z_{0})$ for simplicity when the map $f$ is clear. The length of $f$ at $z_{0}$ is the norm $|\rho_{f}(z_{0})|$. The multiplier and the length are invariant under conjugacy. The characteristic exponent of $f$ at $z_{0}$ is defined to be $\chi_{f}(z_{0}):=n^{-1}\log\lvert\rho_{f}(z_{0})\rvert$ (when $\rho_{f}(z_{0})\neq 0$).
Denote by ${\rm Per}(f)(\mathbb{C})$ the set of all periodic points in $\mathbb{P}^{1}(\mathbb{C})$ of $f$ and define ${\rm Per}^{*}(f)(\mathbb{C}):=\{z_{0}\in{\rm Per}(f)(\mathbb{C}):\rho_{f}(z_{0})\neq 0\}$. When the base field $\mathbb{C}$ is clear, we also write ${\rm Per}(f)$ and ${\rm Per}^{*}(f)$ for simplicity.
The Lyapunov exponent (with respect to the maximal entropy measure) of $f$ is defined by
$\mathcal{L}_{f}:=\int_{\mathbb{P}^{1}(\mathbb{C})}\log|df|d\mu_{f},$
where $\mu_{f}$ is the unique maximal entropy measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ which is also the canonical measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ (see [14, 7, 15, 34]), and the norm of the differential is computed with respect to the spherical metric. In fact, the Lyapunov exponent dose not change with the spherical metric replaced by any Kähler metric on $\mathbb{P}^{1}_{\mathbb{C}}$ in the definition. It is known that $\mathcal{L}_{f}\geq\log(d)/2$, and the equality holds if and only if $f$ is a Lattès map in the sense of Definition 1.1 (see [34, 35]). The quantity $\exp(\mathcal{L}_{f})$ is considered as an average of $|df|$ on $\mathbb{P}^{1}(\mathbb{C})$, which measures the average rate of expansion of $f$ along a typical orbit.
1.2. Exceptional maps
In complex dynamics, the exceptional maps defined below are often considered as exceptional examples among all rational maps of degree $\geq 2$. We may view them as rational maps on $\mathbb{P}^{1}(\mathbb{C})$ related to algebraic groups.
Definition 1.1.
Let $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ be an endomorphism over $\mathbb{C}$ of degree $d\geq 2$.
  • The map $f$ is called Lattès if there is an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi$ is semi-conjugate to $f$. Here, for an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $g$ is semi-conjugate to $h$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$. Furthermore, the map $f$ is called flexible Lattès if there are an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that the multiplication $[n]$ on $E$ is semi-conjugate to $f$ via the quotient map modulo {±$\{\pm 1\}$ on $E$. Otherwise, it is called rigid Lattès.
  • We say that $f$ is of monomial type if it semi-conjugate to the map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$.
We call $f$ exceptional if it is Lattès or of monomial type. It is well-known that $f$ is exceptional if and only if some iterate $f^{k}$ is exceptional ($k\in\mathbb{Z}_{>0}$).
1.3. Statement of the main results
We fix an embedding of the algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ and consider all number fields as subfields of $\mathbb{C}$. Denote the usual absolute value on $\mathbb{C}$ by ||$\cdot$|| ⋅ |.
Our first result shows that the definition field of a rational map is determined by its length spectrum, provided that it is not a flexible Lattès map.
Theorem 1.2.
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree at least $2$. Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $|\rho_{f}(x)|\in\overline{\mathbb{Q}}$. Then $f$ is defined over $\overline{\mathbb{Q}}$.
Remark 1.3.
In Theorem 2.1, we prove a more general version of Theorem 1.2, in which $\overline{\mathbb{Q}}$ can be replaced by any algebraically closed subfield of $\mathbb{C}$ invariant under the complex conjugation. In fact, the actual proof shows that we can weaken the conditions in Theorems 1.2 and 2.1 by discarding the corresponding condition for finitely many periodic points of $f$, see Remark 2.2. For example, in Theorem 1.2, we can only require that $|\rho_{f}(x)|\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$, and the conclusion still holds.
McMullen’s rigidity of multiplier spectrum [16, Corollary 2.3] asserts that multipliers at all periodic points determine the conjugacy class of rational maps up to only finitely many choices, except for the flexible Lattès family. With a standard spread out argument, McMullen’s rigidity theorem implies that, for a rational map $f$ of degree at least $2$ which is not flexible Lattès, if its multipliers at periodic points are all algebraic, then $f$ is defined over $\overline{\mathbb{Q}}$. We sketch the argument as follows. Let $f$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$ that is not flexible Lattès with all multipliers in $\overline{\mathbb{Q}}$. The noetherianity implies that the set $D$ of conjugate classes $[g]$ (in the moduli space $\mathcal{M}_{d}$) such that $g$ has the same multiplier spectrum (i.e., all multipliers at periodic points) as $f$ is an algebraic subset of $\mathcal{M}_{d}$ defined over $\overline{\mathbb{Q}}$, since the multiplier spectrum morphism is defined over $\mathbb{Q}$ and all multipliers of $f$ are in $\overline{\mathbb{Q}}$. Then McMullen’s rigidity implies that $\dim(D)=0$, hence $D$ is a set of finitely many $\overline{\mathbb{Q}}$-points. Thus $f$ is defined over $\overline{\mathbb{Q}}$ as $[f]\in D$.
Theorem 1.2 is a generalization of this result from multiplier spectrum to length spectrum (which contains less information). The rigidity of length spectrum was proved by Ji-Xie [12, Theorem 1.5]. However, the spread out argument does not apply directly in this case as the length spectrum map (and its square) is not algebraic on the moduli space of rational maps. Indeed as shown in [12, §8.1], its square is not even real algebraic (but it is still semialgebraic). In §2.3, we introduce a method to do the spread out argument respecting the structure of $\mathcal{M}_{d}(\mathbb{C})$ as a real algebraic variety using Weil restriction. Another difficulty in the length spectrum case is the lack of noetherianity for semialgebraic subsets. We overcome this difficulty using the notion of admissible subsets introduced in [12].
The following two results concern the $\mathbb{Q}$-vector space spanned by all the (finite) characteristic exponents of periodic points.
Theorem 1.4.
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a rational map of degree $d\geq 2$. Suppose that $f$ is not exceptional. Then the $\mathbb{Q}$-vector space generated by $\{\chi_{f}(z):z\in{\rm Per}^{*}(f)\}$ in $\mathbb{R}$ has infinite dimension.
Theorem 1.5.
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a rational map of degree $d\geq 2$. Assume that there exists a number field $K$ such that
(1.1) $\forall z_{0}\in{\rm Per}(f),\ \exists n=n(z_{0})\in\mathbb{Z}_{>0},\ \lvert\rho_{f}(z_{0})\rvert^{n}\in K.$
Then $f$ is exceptional.
Remark 1.6.
With the help of Remark 1.3, the proof of Theorem 1.5 in §4 indeed shows that we can weaken the condition of Theorem 1.5 by only requiring that there is a number field $K$ such that for all but finitely many $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n(z_{0})}\in K$ for some $n(z_{0})\in\mathbb{Z}_{>0}$.
Finitely many nonzero elements $z_{1},\dots,z_{N}$ in a field $F$ are called multiplicatively independent if for all triples $(m_{1},\dots,m_{N})$ of integers, $z_{1}^{m_{1}}\cdots z_{N}^{m_{N}}=1$ if and only if $m_{1}=\cdots=m_{N}=0$. A sequence $(z_{n})_{n=1}^{\infty}$ in $F\setminus\{0\}$ is called multiplicatively independent if all its finite subsequences are multiplicatively independent. Theorem 1.4 immediately implies the existence of infinitely many multipliers for a non-exceptional $f$ whose absolute values are multiplicatively independent.
Corollary 1.7.
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a rational map of degree $d\geq 2$. Suppose that $f$ is not exceptional. Then there exists a sequence $(x_{j})_{j=1}^{\infty}$ in ${\rm Per}^{*}(f)$ such that the sequence $(\lvert\rho_{f}(x_{j})\rvert)_{j=1}^{\infty}$ is multiplicatively independent in $\mathbb{R}$.
1.4. Motivations and previous results
1.4.1. Milnor’s conjecture
Milnor [18] has showed that an exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ must have all its multipliers of periodic points in the ring of integers $\mathcal{O}_{K}$ for some imaginary quadratic number field $K$, and in fact in $\mathbb{Z}$ when $f$ is not a rigid Lattès map. Milnor conjectured that the converse is also true. Milnor’s conjecture was recently proved by Ji-Xie:
Theorem 1.8 ([12, Theorem 1.13]).
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a rational map of degree $d\geq 2$. Assume that there exists an imaginary quadratic field $K$ such that all multipliers of $f$ belong to $\mathcal{O}_{K}$. Then $f$ is exceptional.
See also [5] for a different proof. Recently Huguin generalized the above result using different approach:
Theorem 1.9 ([10, Theorem 7]).
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a rational map of degree $d\geq 2$. Assume that there exists a number field $K$ such that all multipliers of $f$ belong to $K$. Then $f$ is exceptional.
Since our assumption (1.1) in Theorem 1.5 is weaker than that of Theorem 1.9, Theorem 1.5 is a generalization of Theorem 1.9. Note that our assumption (1.1) is even weaker than the condition that there is a number field $K$ such that
(1.2) $\forall z_{0}\in{\rm Per}(f),\ \exists n\in\mathbb{Z}_{>0},\ (\rho_{f}(z_{0}))^{n}\in K.$
1.4.2. A question of Levy and Tucker
On the other hand, in the 2014 AIM workshop Postcritically Finite Maps In Complex And Arithmetic Dynamics, Levy [13] and Tucker [26] asked the following question independently:
Question 1.10.
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a non-exceptional rational map of degree $d\geq 2$ and let $S$ be the set of all multipliers of periodic points of $f$. Take the subgroup of $\mathbb{C}^{\ast}$ generated by $S\setminus\{0\}$. Is that true that the rank of this group is infinite?
It is not hard to see that our Corollary 1.7 gives a positive answer to (a generalized version of) Levy and Tucker’s question.
1.5. Sketch of the proofs
We have given the idea of Theorem 1.2 before. Here we explain the proofs of Theorem 1.4 and Theorem 1.5.
We first give the idea of the proof of Theorem 1.5. We argue by contradiction and suppose that $f$ is not exceptional. The first step is to reduce to the case where $f$ is defined over $\overline{\mathbb{Q}}$, which can be done using our Theorem 1.2. After enlarging $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ is the rational map obtained from $f$ via replacing the coefficients by their complex conjugations. In the second step, we combine an arithmetic equidistribution theorem with a result of Zdunik [35] (see [10]) on the Lyapunov exponent to get a contradiction. This argument is inspired by Huguin’s proof of Theorem 1.9. Contrary to the case of Theorem 1.9, we can not apply the equidistribution theorem to the one-dimensional dynamical system $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ directly. Our idea is to consider the two-dimensional system $F:=f\times\overline{f}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ instead. More precisely, applying a result of Zdunik [35], we get a sequence $(x_{n})_{n=1}^{\infty}$ of distinct periodic points of $f$ such that
$\lim\limits_{n\to+\infty}\chi_{f}(x_{n})=a>\mathcal{L}_{f}.$
Consider the endomorphism $F:=f\times\overline{f}$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and $\Gamma:=\overline{\{p_{n}=(x_{n},\overline{x_{n}})\}}^{\rm Zar}\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$. Without loss of generality, we may assume that $\Gamma$ is irreducible and that the sequence $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$. Since the Dynamical Manin-Mumford problem is solved for such an endomorphism $F$ by [8] and $\Gamma$ contains a Zariski-dense subset $\{p_{n}:n\geq 1\}$ of $F$-periodic points, the subscheme $\Gamma$ is $F$-periodic itself. Hence we may assume that $\Gamma$ is $F$-invariant after replacing $F$ by a suitable iteration.
Let $\nu_{n}$ be the discrete probability measure equally supported at the union of Galois orbits of iterates of $p_{n}$ under $F$. Then $\nu_{n}$ converges weakly to the canonical measure $\mu$ on $\Gamma$ with respect to $F$ by an equidistribution-type theorem (Theorem 3.1), which is a reformulation of [32, Theorem 3.1], see §3 for details. Applying $\nu_{n}\to\mu$ to the continuous test functions $\max\{\log\lvert\det(dF)\rvert,A\}$ ($A\in\mathbb{R}$) and letting $A\to-\infty$, we get
[…]
which is impossible since the right hand side equals to $2\mathcal{L}_{f}<2a$ by a direct computation.
Next we sketch the proof of Theorem 1.4. According to [6], postcritically finite (PCF) maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. So it suffices to consider the following two cases: 1). $f$ is defined over $\overline{\mathbb{Q}}$, and 2). $f$ is not PCF. For the first case the conclusion follows from Theorem 1.5. For the second case, we need to develop some new techniques, which are presented in §5. In §5, we consider some pseudo-linear algebra (which means that the domain of a map may not be the whole vector space), and the vector space $\mathbb{D}(k)_{\mathbb{Q}}=k^{*}\otimes_{\mathbb{Z}}\mathbb{Q}$ for a field $k$ of characteristic zero. We will actually prove a theorem (Theorem 5.7) stronger than the non-PCF case of Theorem 1.4, see §5 and §6 for details. To prove Theorem 5.7, in §6.1 we first deal with the case that $f$ is defined over $\overline{\mathbb{Q}}$. A key ingredient in this step is [3, Lemma 4.1] which is a consequence of Siegel’s theorem on $S$-integral points. The existence of a non-preperiodic critical point of $f$ is essentially used here. In §6.2, we consider the general case and finish the proof. This is achieved by reducing to the case that $f$ is defined over $\overline{\mathbb{Q}}$ via an algebraic-geometric argument and techniques developed in §5.
1.6. Applications
1.6.1. The Zariski-dense orbit conjecture
By applying Corollary 1.7 we can give a new proof of a special case of the Zariski-dense orbit conjecture.
Zariski-dense orbit conjecture (=ZDO).
Let $k$ be an algebraically closed field of characteristic 00. Given an irreducible quasi-projective variety $X$ over $k$ and a dominant rational self-map $f$ on $X$. If we have $\{g\in k(X):g\circ f=g\}=k$ where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
Remark 1.11.
The converse of ZDO holds and is easy. For some progress on ZDO, see [1, 2, 17, 27, 29] for example.
As an application of Corollary 1.7, we give a new proof of (the most difficult part of) a special case of ZDO, which was first proved by Xie [29, Theorem 1.16].
Theorem 1.12.
Let $X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}$ be the variety of product of $N$ copies of projective line over an algebraically closed field $k$ of characteristic 00. Suppose that $f:X\to X$ is an endomorphism of the form $f_{1}\times\cdots\times f_{N}$ where $f_{j}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ is a non-constant rational map for $1\leq j\leq N$. The ZDO holds for $(X,f)$.
Remark 1.13.
We note that every dominant endomorphism $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ over an algebraically closed field $k$ of characteristic zero must have the form $f_{1}\times\cdots\times f_{N}$, after replacing $f$ by a suitable iterate.
The original proof of Theorem 1.12 in [29] relies on the solution of the (adelic) Zariski-dense orbit conjecture on smooth projective surfaces [29, Theorem 1.15], the notion of adelic topology introduced in [29, §3] and a classification result [29, Proposition 9.2] on invariant subvarieties of $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ (see also [17, 8]). When $n=2$, Pakovich gave another proof [22] of Theorem 1.12 using his classification of invariant curves in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and some height argument. We give a new proof in the case where all the $f_{i}$’s are non-exceptional of degree at least $2$, and this proof does not need the ingredients mentioned above. Other cases of Theorem 1.12 are relatively easy or can be reduced to the case where $f_{i}$ are all non-exceptional, as shown in [29, §9.3].
1.6.2. A characterization of PCF maps
We also show that one can decide whether a rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ is PCF with the information of its multiplier spectrum or length spectrum on periodic points.
Theorem 1.14.
Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a rational map of degree $d\geq 2$. Then the following statements are equivalent:
$(1)$ $f$ is PCF;
$(2)$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $V=V(f)$ of $\mathbb{R}$ is of finite dimension, where $V$ is generated over $\mathbb{Q}$ by
$\{\log\lvert N_{K_{x}/\mathbb{Q}}(\rho_{f}(x))\rvert:x\in{\rm Per}^{*}(f)(\mathbb{C})\};$
$(3)$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $W=W(f)$ of $\mathbb{R}$ is of finite dimension, where $W$ is generated over $\mathbb{Q}$ by
$\{\log\lvert N_{L_{x}/\mathbb{Q}}(\lvert\rho_{f}(x)\rvert)\rvert:x\in{\rm Per}^{*}(f)(\mathbb{C})\}.$
Here $K_{x}$ (resp. $L_{x}$) is any number field containing $\rho_{f}(x)$ (resp. $\lvert\rho_{f}(x)\rvert$) and $N_{K_{x}/\mathbb{Q}}$ (resp. $N_{L_{x}/\mathbb{Q}}$) is the norm map for the extension $K_{x}/\mathbb{Q}$ (resp. $L_{x}/\mathbb{Q}$), i.e., the determinant of the $\mathbb{Q}$-linear transformation induced by multiplication by $\rho_{f}(x)$ (resp. $|\rho_{f}(x)|$).
Clearly, the subspaces $V,W$ above are independent of the choices of the fields $K_{x},L_{x}$, respectively. We also remark that the $\mathbb{Q}$-vector spaces $V$ in $(2)$ and $W$ in $(3)$ are the same space.
Remark 1.15.
Similarly, the actual proof shows that we can weaken the conditions $(2)$ and $(3)$ in Theorem 1.14, in the spirit of Remarks 1.3 and 1.6. Precisely, we can add two more equivalent statements in Theorem 1.14:
$(2)^{*}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all but finitely $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $V^{*}=V^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $V^{*}$ is generated over $\mathbb{Q}$ by
$\{\log\lvert N_{K_{x}/\mathbb{Q}}(\rho_{f}(x))\rvert:x\in{\rm Per}^{*}(f)(\mathbb{C}),\rho_{f}(x)\in\overline{\mathbb{Q}}\};$
$(3)^{*}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all but finitely $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $W^{*}=W^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $W^{*}$ is generated over $\mathbb{Q}$ by
$\{\log\lvert N_{L_{x}/\mathbb{Q}}(\lvert\rho_{f}(x)\rvert)\rvert:x\in{\rm Per}^{*}(f)(\mathbb{C}),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}\}.$
The proofs of Theorems 1.12 and 1.14 will be given in §6.
2. Rational maps with algebraic lengths
Let $K$ be an algebraically closed subfield of $\mathbb{C}$ which is invariant under the complex conjugation $\tau$, i.e., $\tau(K)=K$. The aim of this section is to prove the following theorem.
Theorem 2.1.
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$. Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $\lvert\rho_{f}(x)\rvert\in K$. Then $f$ is defined over $K$.
Remark 2.2.
In fact, we can weaken the condition of Theorem 2.1 by only requiring that $f$ is not flexible Lattès and $\lvert\rho_{f}(x)\rvert\in K$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$. To show this, we need to replace $A$ in the proof in §2.7 by $A=(A_{n}=(RL^{*}(f)_{n})\cap K))_{n}$ and then the original proof works. Here, $A_{n}=(RL^{*}(f)_{n})\cap K$ is the sub-multiset of the multiset $RL^{*}(f)_{n}$ of all elements also in $K$, see §2.6 and §2.7.
Applying Theorem 2.1 to the case $K=\overline{\mathbb{Q}}$, we get Theorem 1.2. We will introduce the notions of Weil restriction, admissible subsets, and transcendental points in §2.1-2.3. Then we construct certain moduli spaces and study the length map on them in §2.4-2.6. After these preparations, we finish the proof of Theorem 2.1 in §2.7.
2.1. Weil restriction
Recall that $K$ is an algebraically closed subfield of $\mathbb{C}$ such that $\tau(K)=K$. Set $L:=K^{\tau}=K\cap\mathbb{R}$. For example, if $K=\mathbb{C}$, then $L=\mathbb{R}$. We need the following easy lemma.
Lemma 2.3.
[…]
Proof of Lemma 2.3.
Since $K$ is algebraically closed, $i\in K$. In particular, $K\neq L$. For every $u\in K$, we may write
$u=\frac{u+\tau(u)}{2}+\frac{u-\tau(u)}{2i}i$
and both $\frac{u+\tau(u)}{2}$ and $\frac{u-\tau(u)}{2i}$ are contained in $L$. This concludes the proof. ∎
In order to prove Theorem 2.1, we need to analyze the length map respecting the structure of $\mathcal{M}_{d}(K)$ as an algebraic variety over $L$ using Weil restriction, due to the appearance of the absolute value in the condition. Hence, the notion of Weil restriction, which we will briefly recall, is necessary. See [23, §4.6] and [4, §7.6] for more information.
Denote by ${\rm Var}_{/K}$ (resp. ${\rm Var}_{/L}$) the category of varieties over $K$ (resp. $L$). For every variety $X$ over $K$, there is a unique variety $R(X)$ over $L$ which represents the functor ${\rm Var}_{/L}\to{\rm Sets}$ sending $V\in{\rm Var}_{/L}$ to ${\rm Hom}(V\otimes_{L}K,X)$. It is called the Weil restriction of $X$ with respect to $K/L$. The functor $X\mapsto R(X)$ is called the Weil restriction. One has the canonical bijection $\psi_{K}:X(K)\to R(X)(L)$ between sets. When $K=\mathbb{C}$, this map is a real analytic diffeomorphism. One may view $X(K)$ as the set of $L$-points of an $L$-algebraic variety $R(X)$ via $\psi_{K}$.
Definition 2.4.
The $L$-Zariski topology on $X(K)$ is the restriction of the Zariski topology on $R(X)$ to $R(X)(L)$ via the identification $\psi_{K}$. A subset $Y$ of $X(K)$ is called $L$-algebraic if it is closed in the $L$-Zariski topology. When $K=\mathbb{C}$, the $L$-Zariski topology is exactly the real Zariski topology as in [12, §8.1.1].
By the last statement of Proposition 2.6 below, the $L$-Zariski topology is stronger than the Zariski topology on $X(K)$.
When $K=\mathbb{C}$, roughly speaking, the Weil restriction is just constructed by splitting a complex variable $z$ into two real variables $x,y$ via $z=x+iy$. For the convenience of the reader, in the following example, we show the concrete construction of $R(X)$ when $X$ is affine.
Example 2.5.
First assume that $X=\mathbb{A}^{N}_{K}$. Then $R(X)=\mathbb{A}^{2N}_{L}$. The map
$\psi_{K}:\mathbb{A}^{N}_{K}(K)=K^{N}\to\mathbb{A}^{2N}_{L}(L)=L^{2N}$
sends $(z_{1},\dots,z_{N})$ to $(x_{1},y_{1},x_{2},y_{2},\dots,x_{N},y_{N})$, where $z_{j}=x_{j}+iy_{j}$ for $1\leq j\leq N$.
Consider the algebra
(2.3) $\mathbb{B}:=L[I]/(I^{2}+1)\simeq L\oplus IL\simeq K$
by Lemma 2.3, where the variable $I$ satisfying $I^{2}=-1$ corresponds to $i$ in $z_{j}=x_{j}+iy_{j}$. Via (2.3), every $f\in K[z_{1},\dots,z_{N}]\simeq\mathbb{B}[z_{1},\dots,z_{N}]$ defines an element
$F:=f(x_{1}+Iy_{1},\dots,x_{N}+Iy_{N})\in\mathbb{B}[x_{1},y_{1},\dots,x_{N},y_{N}].$
Since
$\mathbb{B}[x_{1},y_{1},\dots,x_{N},y_{N}]=L[x_{1},y_{1},\dots,x_{N},y_{N}]\oplus IL[x_{1},y_{1},\dots,x_{N},y_{N}],$
$F$ can be uniquely decomposed to
$F=r(f)+Ii(f)$
where $r(f),i(f)\in L[x_{1},y_{1},\dots,x_{N},y_{N}]$ are determined by $f$.
More generally, if $X$ is the closed subvariety of $\mathbb{A}^{N}_{K}={\rm Spec}K[z_{1},\dots,z_{N}]$ defined by the ideal $(f_{1},\dots,f_{s})$ of $K[z_{1},\dots,z_{N}]$, then $R(X)$ is the closed subvariety of
$R(\mathbb{A}^{N}_{K})=\mathbb{A}^{2N}_{L}={\rm Spec}\,L[x_{1},y_{1},\dots,x_{N},y_{N}]$
defined by the ideal $(r(f_{1}),i(f_{1}),\dots,r(f_{s}),i(f_{s}))$ of $L[x_{1},y_{1},\dots,x_{N},y_{N}]$.
We list some basic properties of Weil restriction without proof.
Proposition 2.6.
Let $X,Y\in{\rm Var}_{/K}$, then we have the following properties:
  • if $X$ is irreducible, then $R(X)$ is irreducible;
  • $\dim R(X)=2\dim X;$
  • if $f:Y\to X$ is a closed (resp. open) immersion, then the induced morphism $R(f):R(Y)\to R(X)$ is a closed (resp. open) immersion.
We still denote by $\tau$ the restriction of $\tau$ to $K$. Denote by $X^{\tau}$ the base change of $X$ by the homomorphism $\tau:K\to K$. This induces a morphism of schemes (over $\mathbb{Z}$) $\tau:X^{\tau}\to X$. It is not a morphism of schemes over $K$. It is clear that $(X^{\tau})^{\tau}=X$.
Example 2.7.
If $X$ is the subvariety of $\mathbb{A}^{N}_{K}={\rm Spec}K[z_{1},\dots,z_{N}]$ defined by the equations $\sum_{I}a_{i,I}z^{I}=0$, $i=1,\dots,s$. Then $X^{\tau}$ is the subvariety of $\mathbb{A}^{N}_{K}$ defined by $\sum_{I}\tau(a_{i,I})z^{I}=0$, $i=1,\dots,s$. The map $\tau:X=(X^{\tau})^{\tau}\to X^{\tau}$ sends a point $(z_{1},\dots,z_{N})\in X(K)$ to $(\tau(z_{1}),\dots,\tau(z_{N}))\in X^{\tau}(K)$.
The following result due to Weil is useful for computing the Weil restriction.
Proposition 2.8.
[23, Exercise 4.7] We have a canonical isomorphism
$R(X)\otimes_{L}K\simeq X\times X^{\tau}.$
Under this isomorphism,
$R(X)(L)=\{(z_{1},z_{2})\in X(K)\times X^{\tau}(K)|\,\,z_{2}=\tau(z_{1})\}$
and $\psi_{K}$ sends $z\in X(K)$ to $(z,\tau(z))\in R(X)(L)$.
2.2. Admissible subsets
In this subsection, we recall the notion of admissible subsets of real algebraic varieties introduced in [12], which satisfies the important decreasing chain condition.
Let $X$ be a variety over $\mathbb{R}$.
Definition 2.9.
[12, §8.2] A closed subset $V$ of $X(\mathbb{R})$ (with the real topology) is called admissible if there is a morphism $f:Y\to X$ of real algebraic varieties and a Zariski closed subset $V^{\prime}\subseteq Y$ such that $V=f(V^{\prime}(\mathbb{R}))$ and $f$ is étale at every point in $V^{\prime}(\mathbb{R})$.
In particular, every algebraic subset of $X(\mathbb{R})$ is admissible.
Remark 2.10.
Denote by $J$ the non-étale locus for $f$ in $Y$. We have $J\cap V^{\prime}(\mathbb{R})=\emptyset$. Since we may replace $(Y,V^{\prime})$ by $(Y\setminus J,V^{\prime}\cap(Y\setminus J))$, in the above definition we may further assume that $f$ is étale.
Proposition 2.11.
[12, Remarks 8.14, 8.15, and Proposition 8.16] We have the following basic properties:
  • (1)
    Let $Y$ be a Zariski closed subset of $X$. If $V$ is admissible as a subset of $X(\mathbb{R})$, then $V\cap Y$ is admissible as a subset of $Y(\mathbb{R})$.
  • (2)
    An admissible subset is semialgebraic.
  • (3)
    Let $V_{1}$ and $V_{2}$ be two admissible closed subsets of $X(\mathbb{R})$. Then $V_{1}\cap V_{2}$ is admissible.
The following theorem shows that admissible subsets satisfy the descending chain condition.
Theorem 2.12.
[12, Theorem 8.17] Let $(V_{n})_{n\geq 1}$ be a sequence of decreasing admissible subsets of $X(\mathbb{R})$. Then there is $N\geq 1$ such that $V_{n}=V_{N}$ for all $n\geq N$.
2.3. Transcendental points
Let $X_{K}$ be a variety over $K$ and $X:=X_{K}\otimes_{K}\mathbb{C}$. We consider $X_{K}$ as a model of $X$ over $K$.
Denote by $\pi_{K}:X\to X_{K}$ the natural projection. For every point $x\in X(\mathbb{C})$, define $Z(x)_{K}$ to be the Zariski closure of $\pi_{K}(x)$ in $X_{K}$ and $Z(x):=\pi_{K}^{-1}(Z(x)_{K})\subseteq X$. It is clear that $Z(x)$ is irreducible. We call $Z(x)$ the $\mathbb{C}/K$-closure of $x$ w.r.t the model $X_{K}$. We say that $x$ is transcendental if $\dim Z(x)\geq 1$ and call $\dim Z(x)$ the transcendental degree of $x$.
The notion of transcendental points plays important role in the proof of some cases of the geometric Bombieri-Lang conjecture by Xie-Yuan [30, 31] and the proof of the dynamical André-Oort conjecture for curves by Ji-Xie [11]. Roughly speaking, a very general point in $Z(x)$ satisfies the same algebraic properties as $x$. In this paper, we study lengths of periodic points in whose definition (see §1.1) we need the norm map $|\cdot|:\mathbb{C}\to\mathbb{R}_{\geq 0}$ which is not algebraic. However $|\cdot|^{2}:\mathbb{C}\to\mathbb{R}$ is real algebraic. For this reason, we need to generalize the notion of transcendental points to respect the structure of $X_{K}(K)$ as an algebraic variety over $L$ and the structure of $X(\mathbb{C})$ as a real algebraic variety, and the generalized notion is important in the proof of Theorem 2.1.
The Weil restriction $R(X)$ of $X$ w.r.t. $\mathbb{C}/\mathbb{R}$ is an algebraic variety over $\mathbb{R}$ and the Weil restriction $R(X_{K})$ of $X_{K}$ w.r.t. $K/L$ is an algebraic variety over $L$. We have $R(X)=R(X_{K})\otimes_{L}\mathbb{R}$. Denote by $\pi_{L}:R(X)\to R(X_{K})$ the natural projection. For every $x\in X(\mathbb{C})$, let $Y(x)_{L}$ be the Zariski closure of $\pi_{L}(\psi_{\mathbb{C}}(x))$ in $R(X_{K})$ and $Y(x):=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, where $\psi_{\mathbb{C}}:X(\mathbb{C})\to R(X)(\mathbb{R})$ is the identification in §2.1. Set $Z^{\mathbb{R}}(x):=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$ which is a real Zariski closed subset of $X(\mathbb{C})$.
We now give a more concrete description of $Z(x)$ and $Z^{\mathbb{R}}(x)$. Let $U_{K}$ be an irreducible affine open neighborhood of $\pi_{K}(x)$ in $X_{K}$ (over $K$). Set $U:=\pi^{-1}_{K}(U_{K})=U_{K}\otimes_{K}\mathbb{C}$. We have a natural embedding $\pi_{K}^{*}:\mathcal{O}(U_{K})\hookrightarrow\mathcal{O}(U)$. We view the elements in $\mathcal{O}^{K}(U):=\pi_{K}^{*}(\mathcal{O}(U_{K}))\subseteq\mathcal{O}(U)$ as the algebraic functions on $U(\mathbb{C})$ defined over $K$. Then we have
$Z(x)\cap U=\{y\in U|\,\,h(y)=0\text{ for every }h\in\mathcal{O}^{K}(U)\text{ with }h(x)=0\}$
and $Z(x)$ is the Zariski closure of $Z(x)\cap U$.
As $\mathcal{O}(R(U)_{\mathbb{C}})=\mathcal{O}(R(U))\otimes_{\mathbb{R}}\mathbb{C}$, every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ can be viewed as a $\mathbb{C}$-valued algebraic function on $R(U)(\mathbb{R})$. Every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ induces a function $h\circ\psi_{\mathbb{C}}$ on $U(\mathbb{C})$. The functions of this form are exactly the $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$. Denote by $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$ the $\mathbb{R}$-algebra of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$. Since algebraic functions are real algebraic, we have a natural embedding $\mathcal{O}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$. By Proposition 2.8, we have
$\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)\simeq\mathcal{O}(U)\otimes_{\mathbb{C}}\tau(\mathcal{O}(U)).$
Let $\mathcal{O}^{L}(R(U)):=\pi_{L}^{*}(\mathcal{O}(R(U_{K})))$ be the set of algebraic functions defined over $L$ on $R(U)$. Let $\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$ the image of $\mathcal{O}^{L}(R(U))\otimes_{L}K$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$, which is the set of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$ defined over $L$. It is clear that $\mathcal{O}^{K}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$. By Proposition 2.8, we have
$\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)\simeq\mathcal{O}^{K}(U)\otimes_{K}\tau(\mathcal{O}^{K}(U)).$
We have
$Z^{\mathbb{R}}(x)\cap U(\mathbb{C})=\{y\in U(\mathbb{C})|\,\,h(y)=0\text{ for every }h\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)\text{ with }h(x)=0\}$
and $Z^{\mathbb{R}}(x)$ is the real Zariski closure of $Z^{\mathbb{R}}(x)\cap U(\mathbb{C})$. This implies the following lemma.
Lemma 2.13.
Let $f_{K}:X^{\prime}_{K}\to X_{K}$ be a morphism between $K$-varieties. Set $X^{\prime}:=X^{\prime}_{K}\otimes_{K}\mathbb{C}$ and let $f:X^{\prime}\to X$ be the morphism induced by $f$. Let $x^{\prime}\in X^{\prime}(\mathbb{C})$ and $x:=f(x^{\prime})\in X(\mathbb{C})$. Then we have $f(Z^{\mathbb{R}}(x^{\prime}))\subseteq Z^{\mathbb{R}}(x)$.
Lemma 2.14.
We have $Z^{\mathbb{R}}(x)\subseteq Z(x)$ and $Z^{\mathbb{R}}(x)$ is Zariski dense in $Z(x)$. In particular, if $x$ is transcendental, then $\dim_{\mathbb{R}}Z^{\mathbb{R}}(x)\geq 1$.
Proof.
It is clear that $Z^{\mathbb{R}}(x)\subseteq Z(x)$. After replacing $X_{K}$ by an irreducible affine open neighborhood of $\pi_{K}(x)$, we may assume that both $X_{K}$ and $X$ are affine. Let $h\in\mathcal{O}(X)$ such that $h(Z^{\mathbb{R}}(x))=0$. Then we have
$h\otimes_{\mathbb{C}}1\in\mathcal{O}(X)\otimes_{\mathbb{C}}\tau(\mathcal{O}(X))=\mathcal{C}^{\mathbb{R}-alg}(X)=\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}.$
View $\mathbb{C}$ as a vector space over $K$ and let $(e_{j})_{j\in J}$ be a $K$-basis of $\mathbb{C}$. We may assume that $0\in J$ and $1=e_{0}$. Write $h\otimes_{\mathbb{C}}1=\sum_{j\in J}g_{j}e_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}$, where $g_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)$. We have $g_{j}(x)=0$ for every $j\in J$.
Let $(f_{n})_{n\in N}$ be a $K$-basis of $\mathcal{O}^{K}(X)$. We may assume that $0\in N$ and $1=f_{0}$. Write
$g_{j}=\sum_{m,n\in N}b_{j,m,n}f_{m}\otimes_{K}\tau(f_{n}).$
The we get
$h\otimes_{\mathbb{C}}1=\sum_{j\in J,m,n\in N}b_{j,m,n}e_{j}f_{m}\otimes_{K}\tau(f_{n}).$
As $(e_{j}f_{m}\otimes\tau(f_{n}))_{j\in J,m,n\in N}$ forms a $K$-basis of
$\mathcal{C}^{\mathbb{R}-{\rm alg}}(X)=\mathcal{O}^{K}(X)\otimes_{K}\tau(\mathcal{O}^{K}(X))\otimes_{K}\mathbb{C},$
and $h\otimes_{\mathbb{C}}1$ is contained in the image of $\mathcal{O}(X)=\mathcal{O}^{K}(X)\otimes_{K}\mathbb{C}$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(X).$ We have
$b_{j,m,n}=0$
for every $n\neq 0$. So $g_{j}=\sum_{m\in N}b_{j,m,0}f_{m}\otimes 1\in\mathcal{O}^{K}(X)$. Since $g_{j}(x)=0$, $g_{j}|_{Z(x)}=0$. Then we have $h|_{Z(x)}=0$ which concludes the proof. ∎
Lemma 2.15.
Assume that $X_{K}$ is affine. Let $h\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)$. For $x\in X(\mathbb{C})$, if $h(x)\in K$, then $h$ is constant on $Z^{\mathbb{R}}(x)$.
Proof.
Write $h=g\circ\psi_{\mathbb{C}}$, where $g\in\mathcal{O}^{L}(R(X))\otimes_{L}K$. Write $g=\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i$, where $g_{1},g_{2}\in\mathcal{O}(R(X_{K}))$. Since $h(x)\in K$, we see that both $\pi_{L}^{*}(g_{1})(\psi_{\mathbb{C}}(x))$ and $\pi_{L}^{*}(g_{2})(\psi_{\mathbb{C}}(x))$ are in $L$. The map $\pi_{L}|_{\psi_{\mathbb{C}}(x)}:\psi_{\mathbb{C}}(x)\to Y(x)_{L}$ induces an embedding
$(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}:\mathcal{O}(Y(x)_{L})\hookrightarrow\mathbb{R}.$
We view elements of $L$ as constant functions contained $\mathcal{O}(Y(x)_{L})$ and note that $(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}(L)=L\subseteq\mathbb{R}.$ For $i\in\{1,2\}$, we have $g_{i}|_{Y(x)_{L}}\in\mathcal{O}(Y(x)_{L})$. As
$(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}(g_{i}|_{Y(x)_{L}})=\pi_{L}^{*}(g_{i})(\psi_{\mathbb{C}}(x))\in L,$
the injectivity of $(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}$ implies that $g_{i}|_{Y(x)_{L}}\in L$, for $i=1,2$. So both $\pi_{L}^{*}(g_{1})$ and $\pi_{L}^{*}(g_{2})$ are constants on $Y(x)=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, with values in $L$. We conclude that $h=g\circ\psi_{\mathbb{C}}=(\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i)\circ\psi_{\mathbb{C}}$ is constant on $Z^{\mathbb{R}}(x)$, since $Z^{\mathbb{R}}(x)=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$. ∎
2.4. Moduli space of rational maps
For $d\geq 2$, let ${\rm Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$. It is a smooth irreducible affine variety of dimension $12d+$ 1 [25]. The group ${\rm PGL}_{2}={\rm Aut}(\mathbb{P}^{1})$ acts on ${\rm Rat}_{d}$ by conjugacy. The geometric quotient
$\mathcal{M}_{d}:={\rm Rat}_{d}/{\rm PGL}_{2}={\rm Spec}(\mathcal{O}({\rm Rat}_{d})^{{\rm PGL}_{2}})$
is the (coarse) moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$, which is an irreducible affine variety of dimension $22d-$ 2, see [25, 24] for a detailed treatment. Let
$\Psi:{\rm Rat}_{d}\to\mathcal{M}_{d}$
be the quotient morphism. Let ${\rm FL}_{d}\subseteq{\rm Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}$. Set $[{\rm FL}_{d}]:=\Psi({\rm FL}_{d})$. The flexible-Lattès locus $[{\rm FL}_{d}]$ is non-empty if and only if $d$ is a square, and in this case $[{\rm FL}_{d}]$ is equidimensional of dimension $1$.
The above construction works over any algebraically closed field of characteristic 00 and commutes with base changes.
In the proof of Theorem 2.1, we need to consider only repelling periodic points of rational maps and the notion of moduli spaces with certain marked periodic points is necessary. We introduce such moduli spaces here. See [11, §8] for more information.
For every $n\in\mathbb{Z}_{>0}$, we construct the space ${\rm Rat}_{d}[n]$ of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with a single marked $n$-periodic point as follows. It is the closed subvariety of ${\rm Rat}_{d}\times\mathbb{P}^{1}$ defined by $f_{{\rm Rat}_{d}}^{n}(g,z)=(g,z)$, where
$f_{{\rm Rat}_{d}}:{\rm Rat}_{d}\times\mathbb{P}^{1}\to{\rm Rat}_{d}\times\mathbb{P}^{1},(g,z)\mapsto(g,g(z))$
is the universal rational map. The first projection
$\phi_{n}:{\rm Rat}_{d}[n]\to{\rm Rat}_{d}$
is a finite morphism of degree $d^{n}+1$. Let
$\lambda_{n}:{\rm Rat}_{d}[n]\to\mathbb{A}^{1},(f_{t},x)\mapsto df_{t}^{n}(x)\in\mathbb{A}^{1}$
be the morphism given by the multiplier at the marked periodic point.
Next we introduce the moduli space of endomorphisms of degree $d$ with a finite sequence of marked periodic points whose periods satisfy certain conditions. Let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. We construct the space ${\rm R_{d}}(s_{1},\dots,s_{n})$ of rational maps of degree $d$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already $(n-1)!$-periodic, $\dots$, such that among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), and we always count with multiplicity, as follows. An $m$-periodic point of $f$ just means a fixed point of $f^{m}$ and we do not require its exact period to be $m$, hence an $m!$-periodic point is automatically a $k!$-periodic point for $k\geq m$. Consider the fiber product $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$ of $s_{n}$ copies of ${\rm Rat}_{d}[n!]$ over ${\rm Rat}_{d}$. For $i\neq j\in\{1,\dots,s_{n}\}$, let
$\pi_{i,j}:({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}\to({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$
be the projection to the $i,j$ coordinates. The diagonal $\Delta\subseteq({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$ is an irreducible component of $({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$. Consider the open subset
$U:=({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}\setminus(\cup_{i\neq j\in\{1,\dots,s_{n}\}}\pi_{i,j}^{-1}(\Delta))$
in which the $s_{n}$ marked points are pairwise distinct. Let $U^{\prime}$ be the subset of $U$ of points $(f,x_{1},\dots,x_{s_{n}})$ satisfying $f^{m!}(x_{i})=x_{i}$ for all integers $1\leq m\leq n$ and $1\leq i\leq s_{m}$. It means that in $U^{\prime}$ the first $s_{m}$ points in the whole sequence of $s_{n}$ marked $n!$-periodic points are indeed $m!$-periodic, for all integers $1\leq m\leq n$. The set $U^{\prime}$ is open and closed in $U$. We then define ${\rm R_{d}}(s_{1},\dots,s_{n})$ to be the Zariski closure of $U^{\prime}$ in $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$. For all integers $0\leq m\leq n$, define
$\phi_{n,m}:{\rm R_{d}}(s_{1},\dots,s_{n})\to{\rm R_{d}}(s_{1},\dots,s_{m}):(f,x_{1},\dots,x_{s_{n}})\mapsto(f,x_{1},\dots,x_{s_{m}})$
to be the restriction morphism, where $s_{0}:=0$ and when $m=0$ the morphism is given by
$\phi_{n,0}:{\rm R_{d}}(s_{1},\dots,s_{n})\to{\rm Rat}_{d}:(f,x_{1},\dots,x_{s_{n}})\mapsto f.$
For all integers $0\leq m_{1}\leq m_{2}\leq n$, we have
$\phi_{m_{2},m_{1}}\circ\phi_{n,m_{2}}=\phi_{n,m_{1}}.$
We define the multiplier morphism on ${\rm R_{d}}(s_{1},\dots,s_{n})$ to be the morphism
$\lambda_{s_{1},\dots,s_{n}}:{\rm R_{d}}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}:(f,x_{1},\dots,x_{s_{n}})\mapsto(df^{n!}(x_{1}),\dots,df^{n!}(x_{s_{n}})).$
Recall that for a rational map $g(z)\in\mathbb{C}(z)$ of degree $\geq 2$ and a fixed point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ of $g$, the multiplicity of the fixed point $z_{0}$ of $g$ is the order of $g(z)-z=0$ at $z_{0}$ when $z_{0}\neq\infty$, and we can define the multiplicity using a conjugation to move $z_{0}$ to $\mathbb{C}$ when $z_{0}=\infty$. For every $k\geq 1$, it is clear that a fixed point $x$ of $f_{t}^{k}$ has multiplicity $1$ iff
(2.4) $\lambda_{k}(f_{t},x)=df_{t}^{k}(x)=\rho_{f^{k}}(x)\neq 1.$
Then the projection $\phi_{k}:{\rm Rat}_{d}[k]\to{\rm Rat}_{d}$ is étale at every point $(f_{t},x)\in{\rm Rat}_{d}[k]\setminus\lambda_{k}^{-1}(1)$. Thus we conclude that $\phi_{n,0}:{\rm R_{d}}(s_{1},\dots,s_{n})\to{\rm Rat}_{d}$ is étale at every point in $(\lambda_{s_{1},\dots,s_{n}})^{-1}((\mathbb{A}^{1}\setminus\{1\})^{s_{n}})$.
Define
$\mathcal{M}_{d}(s_{1},\dots,s_{n}):={\rm R_{d}}(s_{1},\dots,s_{n})/{\rm PGL}_{2}$
to be the moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already $(n-1)!$-periodic, $\dots$, such that in these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), and we always count with multiplicity. The moduli space $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ is affine, in fact, so the construction is relatively easy. The morphisms $\phi_{n,m}$ and $\lambda_{s_{1},\dots,s_{n}}$ descend to
$[\phi_{n,m}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d}(s_{1},\dots,s_{m})$
when $1\leq m\leq n$,
$[\phi_{n,0}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d},$
and
$[\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}.$
Then $[\phi_{n,0}]$ is étale at every point $x\in[\lambda_{s_{1},\dots,s_{n}}]^{-1}((\mathbb{A}^{1}\setminus\{1\})^{s_{n}})$.
All the schemes ${\rm Rat}_{d}$, $\mathcal{M}_{d}$, ${\rm R_{d}}(s_{1},\dots,s_{n})$, and $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ which we consider are defined over $\mathbb{Q}$.
2.5. Length maps
Recall that $K$ is an algebraically closed subfield of $\mathbb{C}$ which is invariant under the complex conjugation. We use the terminology and notations in §2.1-2.4. For $d\geq 2$, let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. Let
$|\lambda_{s_{1},\dots,s_{n}}|:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{R}_{\geq 0}^{s_{n}}$
be the composition of
$[\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{C}^{s_{n}}$
and the norm map
$(a_{1},\dots,a_{s_{n}})\in\mathbb{C}^{s_{n}}\mapsto(|a_{1}|,\dots,|a_{s_{n}}|)\in\mathbb{R}_{\geq 0}^{s_{n}}.$
Define
$q_{s_{1},\dots,s_{n}}:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{R}_{\geq 0}^{s_{n}}$
to be the composition of
$|\lambda_{s_{1},\dots,s_{n}}|:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{R}_{\geq 0}^{s_{n}}$
and the map
$(a_{1},\dots,a_{s_{n}})\in\mathbb{R}_{\geq 0}^{s_{n}}\mapsto(a_{1}^{2},\dots,a_{s_{n}}^{2})\in\mathbb{R}_{\geq 0}^{s_{n}}.$
It is clear that
$q_{s_{1},\dots,s_{n}}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(\mathcal{M}_{d}(s_{1},\dots,s_{n})_{\mathbb{C}}).$
Here the model of
$\mathcal{M}_{d}(s_{1},\dots,s_{n})_{\mathbb{C}}:=\mathcal{M}_{d}(s_{1},\dots,s_{n})\times_{{\rm Spec}(\mathbb{Q})}{\rm Spec}(\mathbb{C})$
over $K$ is taken to be
$\mathcal{M}_{d}(s_{1},\dots,s_{n})_{K}:=\mathcal{M}_{d}(s_{1},\dots,s_{n})\times_{{\rm Spec}(\mathbb{Q})}{\rm Spec}(K).$
By Lemma 2.15, for every $x\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$, if $q_{s_{1},\dots,s_{n}}(x)\in L^{s_{n}}$, then $q_{s_{1},\dots,s_{n}}|_{Z^{\mathbb{R}}(x)}$ is constant. Hence for every $x\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$, if $|\lambda_{s_{1},\dots,s_{n}}|(x)\in L^{s_{n}}$, then $|\lambda_{s_{1},\dots,s_{n}}||_{Z^{\mathbb{R}}(x)}$ is constant.
2.6. Rigidity of length spectrum
In this subsection, we recall the rigidity of length spectrum proved by Ji-Xie [12].
Let $f$ be an endomorphism of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$. As in [12, §8.3] the length spectrum
$L(f)=\{L(f)_{n},n\geq 1\}$
of $f$ is a sequence of finite multisets [^1] A multiset is a set except allowing multiple instances for each of its elements. The number of the instances of an element is called the multiplicity. For example: $\{a,a,b,c,c,c\}$ is a multiset of cardinality $6$, the multiplicities for $a,b,c$ are 2,1,3, respectively., where $L(f)_{n}:=L_{n}(f)$ is the multiset of norms of multipliers (i.e., lengths) of all fixed points of $f^{n}$. In particular, $L(f)_{n}$ is a multiset of non-negative real numbers of cardinality $d^{n}+1$. For every $n\geq 1$, let $RL(f)_{n}$ be the sub-multiset of $L(f)_{n}$ consisting of all elements $>1$. We call
$RL(f):=\{RL(f)_{n},n\geq 1\}$
the repelling length spectrum of $f$ and its subsequence
$RL^{*}(f):=\{RL^{*}(f)_{n}:=RL(f)_{n!},n\geq 1\}$
over the factorials $n!$ the main repelling length spectrum of $f$. By definition, We have $d^{n}+1\geq\#RL(f)_{n}$ for every $n\geq 1$. Observe that the sequence of differences
$(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$
is increasing, since every $n!$-periodic point is $(n+1)!$-periodic and the multiplicity of a repelling fixed point of $f^{n!}$ is $1$ by (2.4). Also, the sequence $(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$ is bounded from above, since every rational map on $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ has only finitely many periodic cycles whose multiplier has norm $\leq 1$ (see [19, Theorem 13.1]). As $L(f),RL(f)$, and $RL^{*}(f)$ are invariant under conjugacy, they descend on $\mathcal{M}_{d}(\mathbb{C})$. For every $[f]\in\mathcal{M}_{d}(\mathbb{C})$, define
$L([f]):=L(f),RL([f]):=RL(f),RL^{*}([f]):=RL^{*}(f)$
for any $f$ in the class $[f]$.
Let $\Omega$ be the set of sequences $\{A_{n},n\geq 1\}$ of multisets consisting of real numbers of norm strictly larger than $1$ satisfying #$\#A_{n}\leq d^{n!}+1$ for $n\geq 1$, and such that for every $a\in A_{n}$ with multiplicity $m\geq 1$, we have $a^{n+1}\in A_{n+1}$ with multiplicity $\geq m$. For $A,B\in\Omega$, we write $A\subseteq B$ if $A_{n}\subseteq B_{n}$ as multisets for every $n\geq 1$. An element $A=\{A_{n}\}\in\Omega$ is called big if $(d^{n!}+1-\#A_{n})_{n=1}^{\infty}$ is bounded. For every endomorphism $f$ of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d$, we have $RL^{*}(f)\in\Omega$ and it is big.
For a big $A\in\Omega$, Ji-Xie proved the following rigidity result on length spectrum, which is a main ingredient of the proof of Theorem 2.1.
Theorem 2.16.
[12, Theorem 8.25] If $A\in\Omega$ is big, then the set
$\{[f]\in\mathcal{M}_{d}(\mathbb{C})\setminus[{\rm FL}_{d}]:A\subseteq RL^{*}(f)\}$
is finite.
2.7. Proof of Theorem 2.1
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$. Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $|\rho_{f}(x)|\in K$. We want to show that $[f]\in\mathcal{M}_{d}(\mathbb{C})$ is not transcendental over $K$ for the model $(\mathcal{M}_{d})_{K}$. Now assume that $[f]$ is transcendental.
Set $A:=RL^{*}(f)\in\Omega$, which is big. Set $s_{n}:=\#A_{n}$ for $n\geq 1$. We may pick a sequence of periodic points $(x_{i})_{i=1}^{\infty}$ such that for every $n\geq 1$, the first $s_{n}$ points $x_{1},\dots,x_{s_{n}}$ are fixed by $f^{n!}$, and we have
$A_{n}=\{|\rho(x_{i})|^{n!}:i=1,\dots,s_{n}\}.$
Let $[f_{n}]\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$ be the point presented by $(f,x_{1},\dots,x_{s_{n}})$. It is clear that $[\phi_{n,0}]([f_{n}])=[f]$ for every $n\geq 1$.
Since $[f]$ is transcendental, for every $n\geq 1$, $[f_{n}]$ is transcendental as well. By Lemma 2.14, we get $\dim_{\mathbb{R}}(Z^{\mathbb{R}}([f_{n}]))\geq 1$ for every $n\geq 1$. Our assumption implies that
$|\lambda_{s_{1},\dots,s_{n}}|([f_{n}])\in L^{s_{n}},n\geq 1.$
Then from the last paragraph of §2.5, we see that $|\lambda_{s_{1},\dots,s_{n}}|$ is constant on $Z^{\mathbb{R}}(f_{n})$. As $|\lambda_{s_{1},\dots,s_{n}}|([f_{n}])\in(1,+\infty)^{s_{n}}$, the morphism $[\phi_{n,0}]$ is étale in a neighborhood of $Z^{\mathbb{R}}(f_{n})$. Since $[\phi_{n,0}]$ is a finite morphism, the subset
$V_{n}:=[\phi_{n,0}](Z^{\mathbb{R}}([f_{n}]))$
is closed in $\mathcal{M}_{d}(\mathbb{C})$ and it is an admissible subset of $\mathcal{M}_{d}(\mathbb{C})$. Moreover, by Lemma 2.13, the sequence $(V_{n})_{n\geq 1}$ is decreasing. By Theorem 2.12, there is $N\geq 1$ such that $V_{n}=V_{N}$ for all $n\geq N$.
Then for every $g\in V_{N}$, we have $A\subseteq RL^{*}([g])$. Since $[f]\not\in[{\rm FL}_{d}]$, the set $Z^{\mathbb{R}}([f_{N}])$ is real irreducible with $\dim_{\mathbb{R}}Z^{\mathbb{R}}([f_{N}])\geq 1$, hence $V_{N}\cap(\mathcal{M}_{d}(\mathbb{C})\setminus[{\rm FL}_{d}])$ is infinite. This contradicts to Theorem 2.16, which concludes the proof. ∎
3. An equidistribution theorem
Results of equidistribution-type play an important role in the study of arithmetic geometry and arithmetic dynamics in past decades. In this section, we state and prove a reformulation (Theorem 3.1) of Yuan’s equidistribution theorem [32, Theorem 3.1]. The reason that we present the theorem is that the equidistribution theorems in the literature (for example, [32]) are only dealing with the Galois orbit of a single small point at one time, which is not sufficient for our proof of Theorem 1.5. Our statement is slightly stronger than [32, Theorem 3.1] as we work with a sequence of (weighted) countable Galois invariant subsets, and this is useful for applications in dynamics. The presented proof of Theorem 3.1 is by applying Yuan’s original equidistribution theorem [32, Theorem 3.1]. We follow the terminology in [32].
For the convenience of readers, we present a short review of some basic notions in Arakelov geometry. See [36, 32, 20, 33] for more details.
Let $K$ be a number field, $X$ a projective variety of dimension $n-1$ over $K$, and $\mathcal{L}$ a line bundle on $X$. We fix an embedding $\overline{K}\to\mathbb{C}_{v}$ for each place $v\in\mathcal{M}_{K}$ of $K$. An adelic metric on $\mathcal{L}$ is a $\mathbb{C}_{v}$-norm $\|\cdot\|_{v}$ on the fiber $\mathcal{L}_{\mathbb{C}_{v}}(x)$ for every $x\in X(\overline{K})$ and $v\in\mathcal{M}_{K}$, satisfying certain continuity and coherence conditions, see [36, §1] for a precise definition.
An adelic metric on $\mathcal{L}$ is called an algebraic metric (or a model metric) if it is induced by an $\mathcal{O}_{K}$-model $(\mathcal{X},\tilde{\mathcal{L}})$ of $(X,\mathcal{L}^{e})$ for some $e\in\mathbb{Z}_{>0}$ (i.e., $\mathcal{X}$ is an integral scheme projective and flat over $\mathcal{O}_{K}$ and $\tilde{\mathcal{L}}$ is a Hermitian line bundle over $\mathcal{X}$ such that the generic fiber of $(\mathcal{X},\tilde{\mathcal{L}})$ is $(X,\mathcal{L}^{e})$). Here the induced metric at an archimedean place is induced by the Hermitian metric of $\tilde{\mathcal{L}}$, and for a non-archimedean place $v$ and $x\in X(\overline{K})$, the induced metric on $\mathcal{L}_{\mathbb{C}_{v}}(x)$ is determined by the lattice $(\tilde{x}^{*}\tilde{\mathcal{L}}_{\mathcal{O}_{\mathbb{C}_{v}}})^{1/e}$ in $\mathcal{L}_{\mathbb{C}_{v}}(x)$, where $\tilde{x}:{\rm Spec}(\mathcal{O}_{\mathbb{C}_{v}})\to\mathcal{X}_{\mathcal{O}_{\mathbb{C}_{v}}}$ is the extension of $x$. The metric induced by such a model always satisfies the conditions of an adelic metric, see [36, §1]. The algebraic metric is called semipositive if $\tilde{\mathcal{L}}$ has semipositive curvature at all archimedean places and non-negative degree on every vertical curve of $\mathcal{X}$.
An (adelic) metric on $\mathcal{L}$ is called semipositive if it is the uniform limit of some sequence of semipositive algebraic metrics on $\mathcal{L}$. Assume that $\mathcal{L}$ is ample and equip it with a semipositive (adelic) metric (such a metric always exists). The resulting metrized line bundle is denoted by $\overline{\mathcal{L}}$. Then for every closed subvariety $Y$ of $X_{\overline{K}}$, we can define its $\overline{\mathcal{L}}$-height $h_{\overline{\mathcal{L}}}(Y)$ by
$h_{\overline{\mathcal{L}}}(Y)=\frac{\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}}{(\dim(Y)+1)\deg_{\mathcal{L}}(\overline{Y})},$
where $\overline{Y}$ is the image of $Y\to X_{\overline{K}}\to X$ (i.e., the closed subvariety of $X$ corresponding to the Galois orbit of $Y$), and $\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}$ is the arithmetic self-intersection number of ¯$\overline{\mathcal{L}}|_{\overline{Y}}$ on $\overline{Y}$. In particular, for every algebraic point $x\in X(\overline{K})$, we get its height $h_{\overline{\mathcal{L}}}(x)$ and the function $h_{\overline{\mathcal{L}}}:{X(\overline{K})}\to\mathbb{R}$ is a Weil height associated to $\mathcal{L}$ (up to a normalizing factor $[K:\mathbb{Q}]$). For an algebraic point $x\in X(\overline{K})$, we have
$h_{\overline{\mathcal{L}}}(x)=\frac{1}{\deg(x)}\sum_{v\in\mathcal{M}_{K}}\sum_{z\in O(x)}(-\log\|s(z)\|_{v}),$
where $O(x)$ is the Galois orbit of $x$ and $s\in\Gamma(X,\mathcal{L})$ is a section non-vanishing on $O(x)$.
Theorem 3.1.
Let $K\subset\overline{\mathbb{Q}}$ be a number field and $X$ be a projective variety over $K$. Fix an embedding of $\overline{\mathbb{Q}}$ into $\mathbb{C}$. Let $\overline{\mathcal{L}}$ be a metrized line bundle on $X$ such that $\mathcal{L}$ is ample and the metric is semipositive. Let $\mu:=\deg_{\mathcal{L}}(X)^{-1}c_{1}(\overline{\mathcal{L}})^{\dim(X)}_{\mathbb{C}}$ be the canonical probability measure on $X(\mathbb{C})$ associated to $\overline{\mathcal{L}}$. For $n\in\mathbb{Z}_{>0}$, let $S_{n}$ be a countable subset of $X(\overline{\mathbb{Q}})$ which is ${\rm Gal}(\overline{\mathbb{Q}}/K)$-invariant. For every $n\geq 1$ and $y\in S_{n}$, fix a real number $a_{n,y}>0$ such that $\sum_{y\in S_{n}}a_{n,y}=1$ and $a_{n,y}=a_{n,\sigma(y)}$, for all $n\geq 1$, $y\in S_{n}$, and $\sigma\in{\rm Gal}(\overline{\mathbb{Q}}/K)$. Assume that the following two conditions hold:
(1) (small) $\sum_{y\in S_{n}}a_{n,y}h_{\overline{\mathcal{L}}}(y)\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$;
(2) (generic) for every proper subvariety $V\subsetneqq X$ of $X$, $\sum_{y\in S_{n}\cap V}a_{n,y}\to 0$ as $n\to+\infty$.
Then the measure $\mu_{n}:=\sum_{y\in S_{n}}a_{n,y}\delta_{y}$ converges weakly to $\mu$ on $X(\mathbb{C})$ as $n\to+\infty$, where $\delta_{y}$ denotes the Dirac measure at the point $y$, i.e., for all continuous function $g$ on $X(\mathbb{C})$ (with the complex topology), we have
[…]
Proof.
For every $n\geq 1$, write $S_{n}=\sqcup_{i=1}^{m_{n}}O_{n,i}$, where $O_{n,i}$ are the distinct Galois orbits contained in $S_{n}$ and $m_{n}\in\mathbb{Z}_{>0}\cup\{\infty\}$.
In [32, Theorem 3.1], Yuan prove the theorem for the case where every $S_{n}$ ($n\geq 1$) is consisting of a single Galois orbit, i.e., $m_{n}=1$ for every $n\geq 1$.
[…]
for every $n\geq 1$.
As $a_{n,y}$ is constant for $y$ in its Galois orbit, we can define $a_{n,i}:=a_{n,y}$ where $y$ is an arbitrary point in $O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$. For all $n\geq 1$ and $1\leq i\leq m_{n}$, set $\mu_{n,i}:=\sum_{y\in O_{n,i}}\delta_{y}$. Then we have
$\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})=1\text{ and }\mu_{n}=\sum_{i=1}^{m_{n}}a_{n,i}\mu_{n,i}$
for every $n\geq 1$. As $h_{\overline{\mathcal{L}}}(y)$ is constant for $y$ in its Galois orbit, we can define $h_{n,i}:=h_{\overline{\mathcal{L}}}(y)$ where $y$ is an arbitrary point in $O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
The small condition $(1)$ asserts that $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$. Taking a suitable subsequence, we may assume that
(3.7) $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\leq h_{\overline{\mathcal{L}}}(X)+\frac{1}{n^{2}}$
for every $n\geq 1$.
As $X$ has only countably many subvarieties defined over $\overline{\mathbb{Q}}$, there is a sequence $V_{1}\subseteq V_{2}\subseteq\cdots\subseteq V_{n}\subseteq\cdots\subsetneqq X$ of increasing proper subvarieties of $X$ defined over $K$, such that for every proper subvariety $V$ of $X$ defined over $\overline{\mathbb{Q}}$, there is $N=N(V)\in\mathbb{Z}_{>0}$ such that $V\subseteq V_{n}$ for all integers $n\geq N$. (By considering the Galois orbits of subvarieties, we can assume that the $V_{n}$’s are defined over $K$.) By Zhang’s theorem of successive minimum [36, Theorem 1.10], we may further assume that
(3.8) $h_{\overline{\mathcal{L}}}(y)\geq h_{\overline{\mathcal{L}}}(X)-\frac{1}{n^{2}}$
for every $n\geq 1$ and $y\in X(\overline{\mathbb{Q}})\setminus V_{n}$. Set
$G_{n}:=\{i\in[1,m_{n}]\cap\mathbb{Z}:O_{n,i}\cap V_{n}=\emptyset\}$
and
$B_{n}:=([1,m_{n}]\cap\mathbb{Z})\setminus G_{n}=\{i\in[1,m_{n}]\cap\mathbb{Z}:O_{n,i}\subseteq V_{n}\},$
for $n\geq 1$. (Note that for every subvariety $V$ of $X$ defined over $K$, either $O_{n,i}\cap V=\emptyset$ or $O_{n,i}\subseteq V$.)
The generic condition $(2)$ shows that after replacing $(S_{n})$ by a suitable subsequence, we may assume that
$1-\frac{1}{n^{2}}\leq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})\leq 1,$
for every $n\geq 1$. Equivalently, we have
(3.9) $0\leq\sum_{i\in B_{n}}a_{n,i}(\#O_{n,i})\leq\frac{1}{n^{2}},$
for every $n\geq 1$.
Since $\mathcal{L}$ is ample, there is a constant $M\in\mathbb{R}_{\geq 0}$ such that
(3.10) $h_{\overline{\mathcal{L}}}(y)\geq h_{\overline{\mathcal{L}}}(X)-M$
for every $y\in X(\overline{\mathbb{Q}})$.
By (3.7), (3.10), and (3.9), we get
$\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})h_{n,i}$ $\leq h_{\overline{\mathcal{L}}}(X)+\frac{1}{n^{2}}-\sum_{i\in B_{n}}a_{n,i}(\#O_{n,i})h_{n,i}$
$\leq h_{\overline{\mathcal{L}}}(X)+\frac{1}{n^{2}}-\sum_{i\in B_{n}}a_{n,i}(\#O_{n,i})(h_{\overline{\mathcal{L}}}(X)-M)$
$=(\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i}))h_{\overline{\mathcal{L}}}(X)+\frac{1+Mn^{2}\sum_{i\in B_{n}}a_{n,i}(\#O_{n,i})}{n^{2}}$
$\leq(\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i}))h_{\overline{\mathcal{L}}}(X)+\frac{1+M}{n^{2}}.$
Set $H_{n}:=\{i\in G_{n}:h_{n,i}\leq h_{\overline{\mathcal{L}}}(X)+1/n\}$ and $I_{n}:=G_{n}\setminus H_{n}$ for $n\geq 1$. By (3.8), we have
$\frac{1+M}{n^{2}}$ $\geq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})(h_{n,i}-h_{\overline{\mathcal{L}}}(X))\geq\frac{-\sum_{i\in H_{n}}a_{n,i}(\#O_{n,i})}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}$
$\geq-\frac{1}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}.$
Therefore, we get
$\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})\leq\frac{2+M}{n}$
for every $n\geq 1$. It follows that
$\sum_{i\in H_{n}}a_{n,i}(\#O_{n,i})\geq 1-\frac{1}{n^{2}}-\frac{2+M}{n}\to 1$
as $n\to+\infty$.
[…]
$h_{\overline{\mathcal{L}}}(X)-\frac{1}{n^{2}}\leq h_{n,i_{n}}\leq h_{\overline{\mathcal{L}}}(X)+\frac{1}{n}$
for every $n\geq 1$, i.e., the small condition $(1)$ holds for $(O_{n,i_{n}})_{n}$. Applying Yuan’s equidistribution Theorem [32, Theorem 3.1] to $(O_{n,i_{n}})_{n}$, we have $\mu_{n,i_{n}}\to\mu$ as $n\to+\infty$. Thus we have
[…]
which is a contradiction. ∎
Remark 3.2.
For a place $v\in\mathcal{M}_{K}$, we can consider two canonical analytic $v$-spaces of $X$, i.e., the $\mathbb{C}_{v}$-analytic space $X^{an}_{\mathbb{C}_{v}}$ and the $K_{v}$-analytic space $X^{an}_{K_{v}}$. We call the former the geometric case and the latter the algebraic case, see [32, §3.1]. Theorem 3.1 deals with the geometric case at an archimedean place. The same idea also applies for a non-archimedean place or the algebraic case, which gives the full analogy of [32, Theorems 3.1 and 3.2].
In order to check the “generic” condition in Theorem 3.1, we need the following lemma. The proof uses ergodic theory with respect to the constructible topology (on algebraic varieties) introduced by Xie [28].
Lemma 3.3.
Let $K$ be a number field and $X$ be a projective variety over $K$. Given a dominant endomorphism $f:X\to X$, and a sequence $(x_{n})_{n=1}^{\infty}$ of $f$-periodic points in $X(\overline{K})$ which are pairwise distinct. Assume that $(x_{n})_{n=1}^{\infty}$ is generic in $X$, i.e., every proper Zariski closed subset $Z\subsetneqq X$ contains only finitely many $x_{n}$. Then for every proper subvariety $V\subsetneqq X$, we have
(3.11) #$\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}\to 0,\ \text{as}\ n\to+\infty,$
where $O_{f}(x_{n})$ is the (forward) orbit of $x_{n}$ under $f$.
Proof.
Clearly, it suffices to show that for every subsequence $(n_{k})_{k}$ of $(n)_{n=1}^{\infty}$, there exists a subsubsequence $(n_{k_{l}})_{l}$ such that
#$\frac{\#(V\cap O_{f}(x_{n_{k_{l}}}))}{\#O_{f}(x_{n_{k_{l}}})}\to 0,\ \text{as}\ l\to+\infty.$
Fix a proper subvariety $V$ in $X$. Let $\lvert X\rvert$ be $X$ equipped with the constructible topology (i.e., the topology of $X$ generated by all its Zariski closed and open subsets) and $\mathcal{M}^{1}(\lvert X\rvert)$ be the space of all probability Radon measures on $\lvert X\rvert$ with the topology of weak convergence relative to all continuous functions on $\lvert X\rvert$. Then $\mathcal{M}^{1}(\lvert X\rvert)$ is sequentially compact (cf. [28, Corollary 1.14]). For $n\in\mathbb{Z}_{>0}$, set
$m_{n}=(\#O_{f}(x_{n}))^{-1}\sum_{z\in O_{f}(x_{n})}\delta_{z}.$
Denote the characteristic function of $V\subsetneqq\lvert X\rvert$ by $1_{V}$, then $1_{V}$ is continuous with respect to the constructible topology. By the sequential compactness of $\mathcal{M}^{1}(\lvert X\rvert)$, it suffices to show that for every subsequence $(n_{k})_{k}$ of $(n)_{n=1}^{\infty}$ with $m_{n_{k}}\to m\ \text{as}\ k\to+\infty$ in $\mathcal{M}^{1}(\lvert X\rvert)$ for some $m\in\mathcal{M}^{1}(\lvert X\rvert)$, we have
[…]
Without loss of generality, we may assume that $(m_{n})$ itself converges to a measure $m\in\mathcal{M}^{1}(\lvert X\rvert)$; and it suffices to show (3.12) in this case. As $f_{*}m_{n}=m_{n}$ for all $n\geq 1$, we see that $f_{*}m=m$ as well. Then according to [28, Lemma 5.3], $m$ must have the form $m=\sum_{y\in S}a_{y}\delta_{O_{f}(y)}$, where $S$ is a countable set of $f$-periodic points in $\lvert X\rvert$ (note that the points in $S$ may not be closed with respect to the Zariski topology of $X$), $a_{y}\in\mathbb{R}_{>0}$ with $\sum_{y\in S}a_{y}=1$, and $\delta_{O_{f}(y)}=(\#O_{f}(y))^{-1}\sum_{z\in O_{f}(y)}\delta_{z}$ for $y\in S$. As $m_{n}\to m$, we get
[…]
Suppose that (3.12) fails. Then there must be a scheme point $y\in S$ with $a_{y}>0$ and $V\cap O_{f}(y)\neq\emptyset$. Denote the exact period of $y$ under $f$ by $k$. Let $Y$ be the Zariski closure of $\{y\}$. Then $Y\subseteq\cup_{j=0}^{k-1}f^{j}(V)$, hence $Y$ is also a proper Zariski closed subset of $X$. Note that
[…]
Hence for every sufficiently large integer $n\gg 1$, we have
(3.13) $x_{n}\in\cup_{j=0}^{\infty}f^{j}(Y)=\cup_{j=0}^{k-1}f^{j}(Y),$
where the last equality follows from $f^{k}(y)=y$ and $Y=\overline{\{y\}}^{{\rm Zar}}$. Note that
$\dim(\cup_{j=0}^{k-1}f^{j}(Y))=\dim(Y)<\dim(X),$
hence $\cup_{j=0}^{k-1}f^{j}(Y)$ contains only finitely many $x_{n}$ because $(x_{n})$ is generic, contradicting to (3.13). ∎
4. Proofs of Theorem 1.5 and the defined over $\overline{\mathbb{Q}}$ case of Theorem 1.4
Proof of Theorem 1.5.
Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $d\geq 2$ and there is a number field $K$ such that for every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n}\in K$ for some integer $n=n(z_{0})\geq 1$. We need to show that such a $f$ must be exceptional (Definition 1.1).
Assume by contradiction that $f$ is not exceptional. In particularly, $f$ is not a flexible Lattès map. For every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert\in\overline{\mathbb{Q}}$, thus $f$ is defined over $\overline{\mathbb{Q}}$ (hence over some number field) by Theorem 1.2. After replacing $K$ by a suitable finite extension of $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$. Here we denote by $\overline{f}$ the rational map obtained from $f$ via replacing the coefficients by their complex conjugations. According to [10, Theorem 9 and Lemma 11], which are essentially proved by Zdunik [35], there exists a sequence $(x_{n})_{n=1}^{\infty}$ of distinct points in ${\rm Per}^{*}(f)$ such that
$a:=\lim\limits_{n\to\infty}\chi_{f}(x_{n})>\mathcal{L}_{f},$
where the limit exists and is finite.
Clearly, $\mathcal{L}_{f}=\mathcal{L}_{\overline{f}}$. For an arbitrary $x\in{\rm Per}(f)$, we have
$\overline{x}\in{\rm Per}(\overline{f}),n_{f}(x)=n_{\overline{f}}(\overline{x}),\text{ and }\rho_{f}(x)=\overline{\rho_{\overline{f}}(\overline{x})},$
hence $\chi_{f}(x)=\chi_{\overline{f}}(\overline{x})$.
Consider the morphism
$F:=f\times\overline{f}:\mathbb{P}^{1}\times\mathbb{P}^{1}\to\mathbb{P}^{1}\times\mathbb{P}^{1}$
over $K$. For $n\in\mathbb{Z}_{>0}$, set $p_{n}=(x_{n},\overline{x_{n}})\in{\rm Per}(F)$. Let $\Gamma$ be the Zariski closure of $\{p_{n}\mid n\in\mathbb{Z}_{>0}\}$ in $\mathbb{P}^{1}\times\mathbb{P}^{1}$, which is defined over $\overline{\mathbb{Q}}$ since the $p_{n}$ are over $\overline{\mathbb{Q}}$. As the points $p_{n}$ are pairwise distinct, we have $\dim\ (\Gamma)\geq 1$. After replacing $(p_{n})$ by a suitable subsequence, we may assume that $\Gamma$ is irreducible and that $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$, i.e., no proper subvariety of $\Gamma$ contains an infinite number of $p_{n}$. Let $\pi_{j}:\Gamma\to\mathbb{P}^{1}$ be the $j$-th projection on $\Gamma\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$ to the $j$-th factor of $\mathbb{P}^{1}\times\mathbb{P}^{1}$, for $j=1,2$. It is clear that $\pi_{1}\circ F=f\circ\pi_{1}$ and $\pi_{2}\circ F=\overline{f}\circ\pi_{2}$.
There are 2 cases to consider, i.e., $\dim(\Gamma)=2$ or $\dim(\Gamma)=1$.
When $\dim(\Gamma)=2$, then $\Gamma=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and the canonical $F$-invariant probability measure on the projective complex manifold $\Gamma(\mathbb{C})=\mathbb{P}^{1}(\mathbb{C})\times\mathbb{P}^{1}(\mathbb{C})$ is $\mu:=\mu_{f}\times\mu_{\overline{f}}$, where $\mu_{f}$ and $\mu_{\overline{f}}$ are the canonical probability measures on $\mathbb{P}^{1}(\mathbb{C})$ which are $f$-invariant and $\overline{f}$-invariant, respectively. (Equivalently, $\mu_{f}$ and $\mu_{\overline{f}}$ are the unique maximal entropy measures of $f$ and $\overline{f}$, respectively.)
When $\dim(\Gamma)=1$, by the solution [8] of the dynamical Manin-Mumford problem for $F$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$, the irreducible curve $\Gamma$ is periodic under $F$. After replacing $f$ by $f^{m}$ for some suitable $m\in\mathbb{Z}_{>0}$, we may assume that the irreducible curve $\Gamma$ is $F$-invariant. Still denote by $\mu$ the canonical $F$-invariant probability measure on the projective complex analytic space $\Gamma(\mathbb{C})$. Recall that if $g:X\to X$ is a polarized surjective endomorphism on a projective variety $X$ over $\mathbb{C}$ (hence $g$ is finite), then the canonical $g$-invariant probability measure $\mu_{g}$ on the projective complex analytic space $X(\mathbb{C})$ is exactly the unique probability measure with continuous potentials on $X(\mathbb{C})$ such that $g_{*}\mu_{g}=\mu_{g}$ and $g^{*}\mu_{g}=\deg(g)^{\dim(X)}\mu_{g}$. Since $f\circ\pi_{1}=\pi_{1}\circ F$ and $\overline{f}\circ\pi_{2}=\pi_{2}\circ F$ are commuting finite morphisms, via the Tate’s limiting construction of canonical measures, it is easy to deduce that
$\deg(\pi_{1})\mu=\pi_{1}^{*}(\mu_{f})\text{ and }\deg(\pi_{2})\mu=\pi_{2}^{*}(\mu_{\overline{f}}).$
[…]
For $n\in\mathbb{Z}_{>0}$, we consider the probability measure
$\nu_{n}=\frac{1}{n_{f}(x_{n})[K_{n}:K]}\sum_{j=0}^{n_{f}(x_{n})-1}\sum_{\tau\in{\rm Gal}(K_{n}/K)}\delta_{F^{j}(\tau(p_{n}))},$
where $K_{n}$ is the Galois closure of $K(x_{n})$ over $K$ in $\overline{\mathbb{Q}}$.
Claim: $\nu_{n}$ converges weakly to $\mu$ as $n\to+\infty$.
We prove the claim using Theorem 3.1. Let $\mathcal{L}$ be the ample line bundle
$\pi_{1}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)\otimes\pi_{2}^{*}\mathcal{O}_{\mathbb{P}^{1}}(1)$
on $\Gamma$. Then $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$. By [36, §2], there exists a unique semipositive metric over $\mathcal{L}$ making $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$ an isometry; denote $\mathcal{L}$ with this metric by $\overline{\mathcal{L}}$. The canonical measure $\mu_{\overline{\mathcal{L}}}$ associated with the metrized line bundle $\overline{\mathcal{L}}$ on $\Gamma(\mathbb{C})$ is exactly the canonical $F$-invariant probability measure $\mu$ on $\Gamma(\mathbb{C})$ considered before, since the metric of ${\overline{\mathcal{L}}}$ can be obtained by Tate’s limit process like $\mu$, see [36, §2] and [32, §3.5]. We need to check the conditions (1) (small) and (2) (generic) in Theorem 3.1. The condition (1) is trivial, since $h_{\overline{\mathcal{L}}}(\Gamma)=0$ and the height of every periodic algebraic point relative to $\overline{\mathcal{L}}$ is zero, in the setting of a polarized algebraic dynamical system (cf. [32]). For the condition (2), let $V$ be an arbitrary proper subvariety of $\Gamma$ and fix $V$. After replacing $V$ by its Galois orbit, we may assume that $V$ is defined over $K$. Then the generic condition (2) follows from Lemma 3.3 immediately. Thus the claim is true.
Let $n\in\mathbb{Z}_{>0}$, take $m\in\mathbb{Z}_{>0}$ such that $\lvert\rho_{f}(x_{n})\rvert^{m}\in K$ by the assumption, and write $l=n_{f}(x_{n})$. For every $\tau\in{\rm Gal}(\overline{\mathbb{Q}}/K)$ and $10\leq j\leq l-$ 1, we have
$\det(dF^{l}(F^{j}(\tau(p_{n}))))^{m}=\det(dF^{l}(\tau(p_{n})))^{m}$
$=$ $\tau(\det(dF^{l}(p_{n}))^{m})=\tau(\rho_{f}(x_{n})^{m}\rho_{\overline{f}}(\overline{x_{n}})^{m})$
$=$ $\tau(\lvert\rho_{f}(x_{n})\rvert^{2m})$
$=$ $\lvert\rho_{f}(x_{n})\rvert^{2m},$
where the first equality holds because both $\tau(p_{n})$ and $F^{j}(\tau(p_{n}))$ are in the same $F$-periodic orbit of length $l$ and the appearance $dF^{l}=d(F^{l})$ in the expressions, hence $\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert=\lvert\rho_{f}(x_{n})\rvert^{2}$. Then by the definition of $\nu_{n}$, we have
[…]
$=$ $\frac{1}{l^{2}[K_{n}:K]}\sum_{j=0}^{l-1}\sum_{\tau\in{\rm Gal}(K_{n}/K)}\log\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert$
$=$ $\frac{1}{l^{2}[K_{n}:K]}\sum_{j=0}^{l-1}\sum_{\tau\in{\rm Gal}(K_{n}/K)}\log\lvert\rho_{f}(x_{n})\rvert^{2}$
$=$ $\frac{2}{l}\log\lvert\rho_{f}(x_{n})\rvert=2\chi_{f}(x_{n}).$
For every $A\in\mathbb{R}$, since the function $\max\{\log\lvert\det(dF)\rvert,A\}$ is continuous, we have
[…]
Let $A\to-\infty$, by the monotone convergence theorem, we have
[…]
We will show that (4.14) leads to a contradiction, by considering the case where $\dim(\Gamma)=2$ and the case where $\dim(\Gamma)=1$ separately.
When $\dim(\Gamma)=2$, it is clear that
[…]
contradicting (4.14).
When $\dim(\Gamma)=1$, then
[…]
$=$ $\mathcal{L}_{f}+\mathcal{L}_{\overline{f}}=2\mathcal{L}_{f},$
contradicting (4.14).
Therefore, $f$ must be exceptional. We have finished the proof. ∎
Proof of Theorem 1.4 when $f$ is defined over $\overline{\mathbb{Q}}$.
Assume that $f$ is defined over $\overline{\mathbb{Q}}$, hence $f$ is defined over a number field $K$. Suppose that the Theorem 1.4 does not hold for $f$. Let $V$ be the $\mathbb{Q}$-span of $\chi_{f}({\rm Per}^{*}(f))$ in $\mathbb{R}$, then $\dim_{\mathbb{Q}}V<\infty$. We can take $M\in\mathbb{Z}_{>0}$ and $x_{1},\dots,x_{M}\in{\rm Per}^{*}(f)$ such that $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ generate $V$ over $\mathbb{Q}$. By enlarging $K$, we may assume that $\lvert\rho_{f}(x_{1})\rvert,\dots,\lvert\rho_{f}(x_{M})\rvert\in K$. Then for every $z_{0}\in{\rm Per}^{*}(f)$, its characteristic exponent $\chi_{f}(z_{0})$ is a linear combination of $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ over $\mathbb{Q}$; in other words, there exist $n\in\mathbb{Z}_{>0}$ and $n_{1},\dots,n_{M}\in\mathbb{Z}$ such that
$\lvert\rho_{f}(z_{0})\rvert^{n}=\lvert\rho_{f}(x_{1})\rvert^{n_{1}}\cdots\lvert\rho_{f}(x_{M})\rvert^{n_{M}}\in K,$
which contradicts Theorem 1.5 since $f$ is not exceptional by the assumption. ∎
Remark 4.1.
We will finish the whole proof of Theorem 1.4 in §6.3.
According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. Therefore, the above proof of Theorem 1.4 for $f$ defined over $\overline{\mathbb{Q}}$ includes the special case where $f$ is PCF (and non-exceptional).
5. Pseudo-linear algebra for field multiplication
5.1. Pseudo-linear algebra
Let $V$ and $W$ be two $\mathbb{R}$-linear spaces. A pseudo-morphism $f:V\to W$ is a pair $(V_{f},f)$ where $V_{f}$ is a linear subspace of $V$ and $f:V_{f}\to W$ is an $\mathbb{R}$-linear map. If $x\in V\setminus V_{f}$, we write $f(x)=\infty$. When $W=\mathbb{R}$, we say that $f$ is a pseudo-function.
Denote by ${\rm PHom}(V,W)$ the set of pseudo-morphisms from $V$ to $W$. For $f,g\in{\rm PHom}(V,W)$, we define $f+g$ to be the pair $(V_{f}\cap V_{g},f|_{V_{f}\cap V_{g}}+g|_{V_{f}\cap V_{g}})$. Then ${\rm PHom}(V,W)$ is a commutative semigroup with $+$ as the operation. We denote by 00 the pair $(V,0)$. We have $0+f=f$ for all $f\in{\rm PHom}(V,W)$. For every $a\in\mathbb{R}$, we define $af$ to be the pair $(V_{f},af)$. We note that $f+(-f)=(V_{f},0)$, which is not 00 if $V_{f}\neq V$. We have a natural embedding ${\rm Hom}(V,W)\hookrightarrow{\rm PHom}(V,W)$.
For $f\in{\rm PHom}(U,V)$ and $g\in{\rm PHom}(V,W)$, we define their composition $g\circ f$ to be $(U_{f}\cap f^{-1}(V_{g}),g\circ f|_{U_{f}\cap f^{-1}(V_{g})})\in{\rm PHom}(U,W)$. Observe that if $f(v)=\infty$, then $g\circ f(v)=\infty$.
Fix a subset $O$ of $V$.
Definition 5.1.
A sequence $(f_{i})_{i=1}^{\infty}$ in ${\rm PHom}(V,\mathbb{R})$ is said to be an $O$-sequence if the following conditions are satisfied:
(i) $f_{i}(O)\subseteq\mathbb{R}_{\geq 0}\cup\{\infty\}$ for $i\geq 1$; (ii) for every $\lambda\in O$, the set $\{i\geq 1:f_{i}(\lambda)\neq 0\}$ is finite.
Clearly, an infinite subsequence of an $O$-sequence is still an $O$-sequence.
Definition 5.2.
Let $(\lambda_{i})_{i=1}^{\infty}$ be a sequence in $O$ and $(f_{i})_{i=1}^{\infty}$ be a sequence in ${\rm PHom}(V,\mathbb{R})$. We say that $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ is an upper triangle $O$-system (resp. weak upper triangle $O$-system) if the following conditions hold:
(i) $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence;
(ii) $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}$ (resp. $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}\cup\{\infty\})$ for $i\geq 1$;
(iii) $f_{j}(\lambda_{i})=0$ for $j>i\geq 1$.
Clearly, an upper triangle $O$-system is a weak upper triangle $O$-system.
Lemma 5.3.
Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be a weak upper triangle $O$-system. Then $(\lambda_{i})_{i=1}^{\infty}$ are linearly independent over $\mathbb{R}$.
Proof.
Since $f_{1}(\lambda_{1})\neq 0$, we see that $\lambda_{1}\neq 0$. Then we only need to show that for all $l\geq 2$, $\lambda_{l}$ is not contained in ${\rm span}_{\mathbb{R}}\{\lambda_{i}:i\leq l-1\}$. Otherwise, $\lambda_{l}=\sum_{i=1}^{l-1}a_{i}\lambda_{i}$ for some $l\geq 2,a_{i}\in\mathbb{R},1\leq i\leq l-1$. Then $f_{l}(\lambda_{l})=\sum_{i=1}^{l-1}a_{i}f_{l}(\lambda_{i})=0$, which contradicts to our assumption. ∎
Let $\tau:V\to V$ be an involution (here we just means that $\tau^{2}=\mathrm{id}$ and $\tau$ may be the identity on $V$).
Lemma 5.4.
Assume that $\tau(O)\subseteq O$. Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangle $O$-system. Then there exists a strictly increasing sequence $(m_{i})_{i=1}^{\infty}$ in $\mathbb{Z}_{>0}$ such that the pair $((\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))_{i=1}^{\infty},(f_{m_{i}})_{i=1}^{\infty})$ is a weak upper triangle $O^{\prime}$-system, where $O^{\prime}=\{\lambda_{m_{i}}+\tau(\lambda_{m_{i}}):i\in\mathbb{Z}_{>0}\}$.
Proof.
It is clear that $(f_{i})_{i=1}^{\infty}$ is also an $O^{\prime}$-sequence. We construct $(m_{i})_{i=1}^{\infty}$ recursively. Set $m_{1}:=1$. As $\tau(\lambda_{1})\in O$, we have $f_{1}(\tau(\lambda_{1}))\in\mathbb{R}_{\geq 0}\cup\{\infty\}$. Since $f_{1}(\lambda_{1})\in\mathbb{R}_{>0}$, we have
$f_{m_{1}}(\lambda_{m_{1}}+\tau(\lambda_{m_{1}}))=f_{1}(\lambda_{1})+f_{1}(\tau(\lambda_{1}))\in\mathbb{R}_{>0}\cup\{\infty\}.$
Assume that we have constructed $m_{1},\dots,m_{l}$ satisfying the conditions for weak upper triangle $O^{\prime}$-systems. Since $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence and $\tau(\lambda_{1}),\cdots,\tau(\lambda_{l})\in O$, there exists $m_{l+1}>m_{l}$ such that $f_{m_{l+1}}(\tau(\lambda_{i}))=0$ for all $i=1,\dots,l$. Then for all $i=1,\dots,l$, we have
$f_{m_{l+1}}(\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))=f_{m_{l+1}}(\lambda_{m_{i}})+f_{m_{l+1}}(\tau(\lambda_{m_{i}}))=0.$
Also,
$f_{m_{l+1}}(\lambda_{m_{l+1}}+\tau(\lambda_{m_{l+1}}))=f_{m_{l+1}}(\lambda_{m_{l+1}})+f_{m_{l+1}}(\tau(\lambda_{m_{l+1}}))\in\mathbb{R}_{>0}\cup\{\infty\}$
and
$f_{m_{i}}(\lambda_{m_{l+1}}+\tau(\lambda_{m_{l+1}}))=f_{m_{i}}(\lambda_{m_{l+1}})+f_{m_{i}}(\tau(\lambda_{m_{l+1}}))\in\mathbb{R}_{\geq 0}\cup\{\infty\}.$
We conclude the proof. ∎
By Lemma 5.4 and Lemma 5.3 we get the following result.
Corollary 5.5.
Assume that $\tau(O)\subseteq O$. Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangle $O$-system. Then $\dim_{\mathbb{R}}{\rm span}_{\mathbb{R}}\{2^{-1}(\lambda_{i}+\tau(\lambda_{i})):i\geq 1\}=\infty$.
Remark 5.6.
Note that the discussion in this subsection also applies with $\mathbb{R}$ replaced by any ordered field $F$ (for example, $\mathbb{Q}$ with the usual order).
5.2. Linear algebra for multiplication
For every field $k$ of characteristic 00, denote by $\mu_{k}$ the subgroup of roots of unity in $k$. Denote by
${\rm rog}:k^{*}\to\mathbb{D}(k):=k^{*}/\mu_{k}$
the quotient map. Extend ${\rm rog}$ to a map ${\rm rog}:k\to\mathbb{D}(k)\cup\{\infty\}$ by sending 00 to $\infty$. Here we use the notation “${\rm rog}$” since it is an analogy of the classical $\log$ function to some extent. The embedding $k\hookrightarrow\overline{k}$ gives a natural embedding $\mathbb{D}(k)\hookrightarrow\mathbb{D}(\overline{k})$ as multiplicative abelian groups. Set $\mathbb{D}(k)_{\mathbb{Q}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{Q}$, where $\mathbb{D}(k)$ is viewed as a multiplicative abelian group, hence a $\mathbb{Z}$-module. Then $\mathbb{D}(k)_{\mathbb{Q}}$ is the subspace of $\mathbb{D}(\overline{k})$ spanned by $\mathbb{D}(k)$ over $\mathbb{Q}$. Set $\mathbb{D}(k)_{\mathbb{R}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{R}$.
Let $A\subseteq k$ be an integral domain with ${\rm Frac}(A)=k$. Define
$\mathbb{D}(A):={\rm rog}(A\setminus\{0\})\subseteq\mathbb{D}(k),$
which is a subsemigroup of $\mathbb{D}(k)$. For every prime ideal $p$ of $A$, the surjective projection $A\to A/p$ induces a surjective map
$s_{p}:\mathbb{D}(A)\cup\{\infty\}\to\mathbb{D}(A/p)\cup\{\infty\}.$
In fact, we may view $s_{p}$ as a pseudo-morphism
$s_{p}:\mathbb{D}(k)_{\mathbb{R}}\to\mathbb{D}({\rm Frac}(A/p))_{\mathbb{R}}$
with domain $V_{s_{p}}:=(A\setminus p)\otimes_{\mathbb{Z}}\mathbb{R}$.
5.3. Norms
Let $k$ be a field of characteristic 00. For every finite field extension $\widetilde{k}$ of $k$, denote by $N_{\widetilde{k}/k}:\widetilde{k}\to k$ the norm map. The norm map induces a map
$\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{Q}}\to\mathbb{D}(k)_{\mathbb{Q}},{\rm rog}(x)\mapsto[l:k]^{-1}{\rm rog}(N_{l/k}(x)),$
where $l$ is any finite field extension of $k$ containing $x$. One can check that $\mathbf{n}_{k}$ is well defined and is $\mathbb{Q}$-linear. We also denote by $\mathbf{n}_{k}$ its $\mathbb{R}$-linear extension $\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{R}}\to\mathbb{D}(k)_{\mathbb{R}}$. When the field $k$ is clear, we also write $\mathbf{n}$ for $\mathbf{n}_{k}$.
5.4. Valuations
Assume that $K$ is a number field. Denote by $\mathcal{M}_{K}$ the set of all places of $K$. For every $v\in\mathcal{M}_{K}$, we define a map $\phi_{v}:\mathbb{D}(K)_{\mathbb{R}}\to\mathbb{R}$, by $\phi_{v}({\rm rog}(x))=-\log(\lvert x\rvert_{v})$ for $x\in K^{*}$ and extending $\mathbb{R}$-linearly, where $\left|\cdot\right|_{v}$ is a fixed multiplicative norm on $K$ corresponding to $v$. It is easy to check that the map $\phi_{v}$ is well-defined and $\mathbb{R}$-linear. We denote the restriction of $\phi_{v}$ to $\mathbb{D}(K)_{\mathbb{Q}}$ by $\phi_{v}:\mathbb{D}(K)_{\mathbb{Q}}\to\mathbb{R}$ as well. For every $a\in\mathbb{D}(K)_{\mathbb{R}}$, it is clear that the set $\{v\in\mathcal{M}_{K}:\phi_{v}(a)\neq 0\}$ is finite.
Let $S$ be a finite subset of $\mathcal{M}_{K}$ containing all the archimedean places. Let $\mathcal{O}_{K,S}$ be the ring of $S$-integers in $K$. Let $\mathcal{O}$ be the integral closure of $\mathcal{O}_{K,S}$ in $\overline{K}$. For every $v\in\mathcal{M}_{K}\setminus S$ and $\lambda\in\mathcal{O}$, we have $\phi_{v}\circ\mathbf{n}_{K}(\lambda)\geq 0$. Write $\mathcal{M}_{K}\setminus S=\{v_{1},v_{2},\dots\}$. Then $(\phi_{v_{i}})_{i=1}^{\infty}\subseteq{\rm Hom}(\mathbb{D}(K)_{\mathbb{R}},\mathbb{R})$ is an $\mathcal{O}_{K,S}$-sequence and $(\phi_{v_{i}}\circ\mathbf{n}_{K})_{i=1}^{\infty}\subseteq{\rm Hom}(\mathbb{D}(\overline{K})_{\mathbb{R}},\mathbb{R})$ is an $\mathcal{O}$-sequence.
5.5. Complex conjugation and absolute value
Denote by $\tau:\mathbb{C}\to\mathbb{C}$ the complex conjugation. Then $\mathbb{R}$ is the invariant subfield $\mathbb{C}^{\tau}$ of $\mathbb{C}$ under $\tau$. As $\mathbb{Q}$-vector spaces, we have an identification
$\mathbb{D}(\mathbb{R})=\mathbb{R}^{*}/\{\pm 1\}\to\mathbb{R},{\rm rog}(a)\mapsto\log\lvert a\rvert,$
where the latter $\log$ is the classical one on $\mathbb{R}_{>0}$. Using this identification, the absolute value map on $\mathbb{C}$ can be viewed as the norm map
$\mathbf{n}_{\mathbb{R}}:\mathbb{D}(\mathbb{C})\to\mathbb{D}(\mathbb{R}),{\rm rog}(x)\mapsto\frac{{\rm rog}(x)+{\rm rog}(\tau(x))}{2}.$
Let $\mathbf{k}$ be an algebraically closed subfield of $\mathbb{C}$ stable under the complex conjugation $\tau$. Still denote by $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ the restriction of the complex conjugation on $\mathbf{k}$, which is an involution of $\mathbf{k}$. Denote by $\mathbf{k}^{\tau}=\mathbf{k}\cap\mathbb{R}$ the $\tau$-invariant subfield of $\mathbf{k}$. Then the restriction of the absolute value $\mathbf{n}_{\mathbb{R}}$ on $\mathbf{k}$ is
$\mathbf{n}_{\mathbf{k}^{\tau}}:\mathbb{D}(\mathbf{k})\to\mathbb{D}(\mathbf{k}^{\tau}),{\rm rog}(x)\mapsto\frac{{\rm rog}(x)+{\rm rog}(\tau(x))}{2}.$
We shall prove the following theorem involving an involution $\tau$ (which may not be the complex conjugation and may be the identity) in the next section.
Theorem 5.7.
Let $\mathbf{k}$ be an algebraically closed field of characteristic 00. Let $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ be an element with $\tau^{2}=\mathrm{id}$. If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace in $\mathbb{D}(\mathbf{k}^{\tau})_{\mathbb{Q}}$ spanned by $\{\mathbf{n}_{\mathbf{k}^{\tau}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
Applying Theorem 5.7 to the case where $\mathbf{k}=\mathbb{C}$ and $\tau$ is the complex conjugation, we deduce Theorem 1.4 in the case where $f$ is not PCF.
Remark 5.8.
Setting $\tau=\mathrm{id}$, then from Theorem 5.7 we get the following result:
Assume that $\mathbf{k}$ is an algebraically closed field of characteristic 00. If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace in $\mathbb{D}(\mathbf{k})_{\mathbb{Q}}$ spanned by $\{{\rm rog}(\rho_{f}(x)):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
6. Proofs of Theorem 5.7 and Theorem 1.4
6.1. Proof of Theorem 5.7: the case $\mathbf{k}=\overline{\mathbb{Q}}$
Assume that $\mathbf{k}=\overline{\mathbb{Q}}$ and let $\tau$ be an element in ${\rm Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ with $\tau^{2}=\mathrm{id}$.
Denote by $\mathcal{C}_{f}$ the set of critical points of $f$. Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ such that the (forward) $f$-orbit $O_{f}(o)$ of $o$ is infinite. We fix such a critical point $o$. Let $X$ be the union of all (forward) orbits of periodic critical points of $f$. Then $X$ is finite.
Pick a number field $K$ satisfying $\tau(K)=K$ such that $f,o$, and all points in $X$ are defined over $K$.
Denote by $\mathcal{M}_{K}$ the set of all places of $K$. Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all the archimedean places, satisfying $\tau(B)=B$, such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$. Then we have $\tau(\mathcal{O}_{K,B})=\mathcal{O}_{K,B}$. For $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$, set $\lambda(x)=(n_{f}(x))^{-1}{\rm rog}(\rho_{f}(x))\in\mathbb{D}(\overline{\mathbb{Q}})_{\mathbb{R}}\cup\{\infty\}$. Recall that $n_{f}(x)\in\mathbb{Z}_{>0}$ is the exact period of $x$ and $\rho_{f}(x)=df^{n_{f}(x)}(x)\in\overline{\mathbb{Q}}$ is the multiplier of $x$. For all $x\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$, we have $\mathbf{n}_{K}(\lambda(x))\in O:=\mathbb{D}(\mathcal{O}_{K,B})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ by [24, Corollary 2.23 (a)]. Note that $O=\tau(O)$ is invariant under $\tau$, where $\tau$ descends to $\mathbb{D}(K)$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(K)_{\mathbb{Q}}$.
Denote by $\mathbb{C}_{v}$ the completion of the algebraic closure of $K_{v}$ for $v\in\mathcal{M}_{K}$. Every embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ gives a bijection $\sigma:{\rm Per}(f)(\overline{\mathbb{Q}})\to{\rm Per}(f)(\mathbb{C}_{v})$. Observe that for every $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$, we have $\sigma(\lambda(x))=\lambda(\sigma(x))$.
For every $v\in\mathcal{M}_{K}\setminus B$ and $x\in\mathbb{P}^{1}(\mathbb{C}_{v})$, denote by $\tilde{x}\in\mathbb{P}^{1}(\overline{\widetilde{K_{v}}})$ the reduction of $x$ in the special fiber at $v$ and $f_{v}:\mathbb{P}^{1}_{\overline{\widetilde{K_{v}}}}\to\mathbb{P}^{1}_{\overline{\widetilde{K_{v}}}}$ the reduction of $f$. (Here, $\overline{\widetilde{K_{v}}}$ is a fixed algebraic closure of the residue field $\widetilde{K_{v}}$ of $K_{v}$.) After enlarging $B$, we can assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$, where $X_{v}$ is the reduction of $X$ contained in $\mathbb{P}^{1}(\widetilde{K_{v}})\subseteq\mathbb{P}^{1}(\overline{\widetilde{K_{v}}})$.
Observe that for every $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$ of exact period $n\geq 1$ and every $v\in\mathcal{M}_{K}\setminus B$, we have $\phi_{v}(\mathbf{n}_{K}(\lambda(x)))\geq 0$ by [24, Corollary 2.23 (a)]. Moreover, the following statements are equivalent by Hensel’s lemma:
(i) $\phi_{v}(\mathbf{n}_{K}(\lambda(x)))>0$;
(ii) there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ such that $(f^{n}_{v})^{\prime}(\widetilde{\sigma(x)})=0$;
(iii) there exist an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$, $q\in\mathcal{C}_{f}$, and $m\in\mathbb{Z}_{\geq 0}$, such that $\widetilde{\sigma(q)}$ is $f_{v}$-periodic with $\widetilde{\sigma(x)}=f_{v}^{m}(\widetilde{\sigma(q)})$.
For every $v\in\mathcal{M}_{K}\setminus B$, denote by $P_{v}$ the union of all $f_{v}$-orbits of periodic critical points of $f_{v}$. The set $P_{v}$ is finite. For every $v\in\mathcal{M}_{K}\setminus B$ and $q\in P_{v}$, by Hensel’s lemma there exists a unique periodic point $y=y(q)\in{\rm Per}(f)(\mathbb{C}_{v}\cap\overline{\mathbb{Q}})$ such that $\tilde{y}=q$. Then there exists a unique ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit $O(q)$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ such that for some (then every) $x\in O(q)$, there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ satisfying $\widetilde{\sigma(x)}=q$ (here $O(q)$ is the ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit of $y=y(q)$). In particular, we have $X_{v}\subseteq P_{v}$ and $\cup_{q\in X_{v}}O(q)=X$. It follows that the set
$Q_{v}:=\{x\in{\rm Per}(f)(\overline{\mathbb{Q}}):\phi_{v}(\mathbf{n}_{K}(\lambda(x)))>0\}=\bigcup_{q\in P_{v}}O(q)\subseteq{\rm Per}(f)(\overline{\mathbb{Q}})$
is finite and $X\subseteq Q_{v}$. Moreover, $Q_{v}=X$ if and only if $P_{v}=X_{v}$.
Lemma 6.1.
The set $S:=\{v\in\mathcal{M}_{K}\setminus B:P_{v}\setminus X_{v}\neq\emptyset\}$ is infinite.
Proof.
Recall that $o\in\mathcal{C}_{f}$ is a critical point of $f$ which is not $f$-preperiodic and we assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$. By [3, Lemma 4.1], there are infinitely may $v\in\mathcal{M}_{K}\setminus B$, such that there exists $n\in\mathbb{Z}_{>0}$ with $f_{v}^{n}(\tilde{o})=\tilde{o}$. For such a place $v$, we have $\tilde{o}\in P_{v}\setminus X_{v}$, which concludes the proof. ∎
Lemma 6.2.
There exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{i})))>0$ for all $i\geq 1$ and $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))=0$ for $j\neq i$. In particular, $((\mathbf{n}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangle $O$-system for $\mathbb{D}(K)_{\mathbb{Q}}$.
Proof.
We construct these two sequences recursively.
By Lemma 6.1, $S$ is infinite. Pick $v_{1}\in S$, then there exists $x_{1}\in Q_{v_{1}}\setminus X\subseteq$ ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$. We have $\phi_{v_{1}}(\mathbf{n}(\lambda(x_{1})))>0$.
Assume that we have constructs $x_{1},\dots,x_{m}\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and $v_{1},\dots,v_{m}\in$ $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))\geq 0$ and the equality holds if and only if $j\neq i$. The set $\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ is finite. Then there exists a finite set $T_{m}\subseteq\mathcal{M}_{K}$ such that for all $x\in\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ and $v\in\mathcal{M}_{K}\setminus T_{m}$, we have $\phi_{v}(\mathbf{n}(\lambda(x)))=0$. By Lemma 6.1, we can take a place $v_{m+1}\in$ $S\setminus(\{v_{1},\dots,v_{m}\}\cup T_{m})$. Then we have $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{i})))=0$ for all $1\leq i\leq m$. Pick $x_{m+1}\in Q_{v_{m+1}}\setminus X$. We have $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{m+1})))>0$. It follows that $x_{m+1}\notin\cup_{i=1}^{m}Q_{v_{i}}$. Thus $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{m+1})))=0$ for all $1\leq i\leq m$. We conclude the proof of Lemma 6.2. ∎
Then we conclude the proof by Corollary 5.5 and Remark 5.6.∎
6.2. Proof of Theorem 5.7: the general case
Denote by $\mathcal{C}_{f}$ the set of critical points of $f$. Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ which is not $f$-preperiodic. We fix such a critical point $o$. Fix a subfield $K$ of $\mathbf{k}$ such that $K/\mathbb{Q}$ is finite generated, $\tau(K)=K$, and $o,f$ are defined over $K$. Without loss of generality, we may assume that $\mathbf{k}=\overline{K}$.
Take a finite generated $\mathbb{Z}$-subalgebra $A$ of $K$ with ${\rm Frac}(A)=K$ and $\tau(A)=A$. After shrinking ${\rm Spec}(A)$, we may assume that there exists an endomorphism $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ over $A$ whose restriction $f_{K}:\mathbb{P}^{1}_{K}\to\mathbb{P}^{1}_{K}$ on the generic fiber $\mathbb{P}^{1}_{K}$ satisfies $f=f_{K}\otimes_{K}\mathbf{k}$.
For every $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$, denote by $f_{c}$ the specialization of $f_{A}$ at $c$, and $o_{c}$ the specialization of $o$ at $c$. Then $o_{c}$ is a critical point of $f_{c}$. By [9, Lemma 3.3], we can take $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$ such that the $f_{c}$-orbit of $o_{c}$ is infinite, i.e., $o_{c}$ is not $f_{c}$-preperiodic. In particular, the specialization map $f_{c}$ is not PCF as well. We take a number field $L\subseteq\overline{K}$ such that $c$ is defined over $L$ and $\tau(L)=L$. Denote by $A_{1}$ the $\mathbb{Z}$-algebra generated by $A$, $\mathcal{O}_{L}$, and $\tau(\mathcal{O}_{L})$ in $A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}}$. We may replace ${\rm Spec}(A)$ by a suitable non-empty affine Zariski open set of ${\rm Spec}(A_{1})$ such that $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ is still everywhere well-defined and $c\in{\rm Spec}(A\otimes_{\mathcal{O}_{L}}L)(L)$, after the replacement. We view ${\rm Spec}(A)$ as an $\mathcal{O}_{L}$-scheme. After shrinking ${\rm Spec}(A)$, the Zariski closure of $c$ in ${\rm Spec}(A)$ is isomorphic to ${\rm Spec}(\mathcal{O}_{L,S})$ for a finite set $S\subseteq\mathcal{M}_{L}$ of places of $L$ containing all archimedean places, which corresponds to a prime ideal $p$ of $A$.
Denote by $s_{p}:\mathbb{D}(K)_{\mathbb{R}}\dashrightarrow\mathbb{D}(L)_{\mathbb{R}}$ the pseudo-morphism associated with $p$ as in §5.2. We have $s_{p}(\mathbb{D}(A))\subseteq\mathbb{D}(\mathcal{O}_{L,S})\cup\{\infty\}$. Then for every $v\in\mathcal{M}_{L}\setminus S$ and $\lambda\in\mathbb{D}(A)$, we have $\phi_{v}(s_{p}(\lambda))\in\mathbb{R}_{\geq 0}\cup\{\infty\}$. Moreover, for every $\lambda\in\mathbb{D}(A)$ with $s_{p}(\lambda)\neq\infty$, there are only finitely many $v\in\mathcal{M}_{L}\setminus S$ such that $\phi_{v}(s_{p}(\lambda))\neq 0$.
For every $y\in{\rm Per}(f)(\overline{K})$, denote by $y_{c}$ the set of $x\in{\rm Per}(f_{c})(\overline{L})$ whose image is contained in the Zariski closure of the image of $y$ in $\mathbb{P}_{A}^{1}$. For every $y\in{\rm Per}(f)(\overline{K})$, $y_{c}$ is finite and nonempty. On the other hand, for every $x\in{\rm Per}(f_{c})(\overline{L})$, the set $\{y\in{\rm Per}(f)(\overline{K}):x\in y_{c}\}$ is finite and nonempty. Moreover, if $x\in y_{c}$, then
$s_{p}(\mathbf{n}_{K}(\lambda(y)))=\mathbf{n}_{L}(\lambda(x)).$
Since the set of $x\in{\rm Per}(f_{c})(\overline{L})$ with $\mathbf{n}_{L}(\lambda(x))=\infty$ is finite, the set
$W_{c}:=\{y\in{\rm Per}(f)(\overline{K}):s_{p}(\mathbf{n}_{K}(\lambda(y)))=\infty\}$
is also finite. Similarly, $W_{\tau(c)}:=\{y\in{\rm Per}(f)(\overline{K}):s_{\tau(p)}(\mathbf{n}_{K}(\lambda(y)))=\infty\}$ is finite.
By Lemma 6.2, there exist $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}(f_{c})(\overline{L})$ and $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{L}\setminus S$ such that $((\mathbf{n}_{L}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangle $(\mathbb{D}(\mathcal{O}_{L,S})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(L)_{\mathbb{Q}}$. For every $i\in\mathbb{Z}_{>0}$, we can take $y_{i}\in{\rm Per}(f)(\overline{K})$ such that $x_{i}\in(y_{i})_{c}$, and hence
$s_{p}(\mathbf{n}_{K}(\lambda(y_{i})))=\mathbf{n}_{L}(\lambda(x_{i})).$
After removing finitely many terms, we can assume that $y_{i}\notin W_{c}\cup W_{\tau(c)}$ for all $i\geq 1$. It follows that $\mathbf{n}_{K}(\lambda(y_{i}))\in{\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ for $i\geq 1$. Observe that $(\phi_{v_{i}}\circ s_{p})_{i=1}^{\infty}$ is a $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-sequence. It follows that
$((\mathbf{n}_{K}(\lambda(y_{i})))_{i=1}^{\infty},(\phi_{v_{i}}\circ s_{p})_{i=1}^{\infty})$
is an upper triangle $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(K)_{\mathbb{Q}}$. Since
${\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$
is invariant under $\tau$, we conclude the proof by Corollary 5.5 and Remark 5.6.∎
6.3. Proof of Theorem 1.4
There are two cases:
1. The case $f$ is PCF. According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. We conclude that $f$ is defined over $\overline{\mathbb{Q}}$ since it is non-exceptional, and Theorem 1.4 was already proved in §4.
2. The case $f$ is not PCF. Then Theorem 1.4 is a direct consequence of Theorem 5.7.
We finish the proof of Theorem 1.4.∎
7. Proofs of the Applications
7.1. Proof of Theorem 1.12
Recall that our goal is to prove the Zariski-dense orbit conjecture for $f=f_{1}\times\cdots\times f_{N}:X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}\to X$ where $f_{1},\dots,f_{N}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ are non-constant endomorphisms on the projective line over an algebraically closed field $k$ of characteristic 00.
Without loss of generality, we may assume that $k$ is of finite transcendence degree over $\mathbb{Q}$. Fix an embedding of $k$ into $\mathbb{C}$. We view $f$ as an endomorphism on $X$ defined over $\mathbb{C}$. According to [29, Theorem 3.34], we may assume that all $f_{j}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ has degree at least $2$ for $1\leq j\leq N$.
Assume first that all $f_{j}$ are not exceptional, $1\leq j\leq N$. Corollary 1.7 implies that we can take $x_{j}\in{\rm Per}^{*}(f_{j})(\mathbb{C})$ for $1\leq j\leq N$ such that the multipliers $\rho_{f_{1}}(x_{1}),\cdots,\rho_{f_{N}}(x_{N})$ are multiplicatively independent in $\mathbb{C}$. After replacing $f$ by an iterate, we may assume that $f_{j}(x_{j})=x_{j}$ for $1\leq j\leq N$, and the multipliers $(\rho_{f_{j}}(x_{j})=f_{j}^{\prime}(x_{j}))_{j=1}^{N}$ are still multiplicatively independent. Denote $x=(x_{1},\dots,x_{N})\in X(k)$. Then $x$ is a fixed point of $f$ (smooth in the fixed locus of $f$) such that the eigenvalues of $df|_{x}$ are nonzero and multiplicatively independent. Then the conclusion follows from [1].
Assume that all $f_{j}$ are exceptional, $1\leq j\leq N$. This case is easy, and we just refer to the proof in the first several paragraphs of [29, §9.3].
We may assume that $0\leq s\leq N$ such that $f_{1},\cdots,f_{s}$ are not exceptional and $f_{s+1},\cdots,f_{N}$ are exceptional. Let $l(f)=\min\{s,N-s\}\geq 0$. Then we have done in the case $l(f)=0$. Then an induction on $l(f)\in\mathbb{Z}_{\geq 0}$ proves Theorem 1.12, as shown in the last several paragraphs of [29, §9.3].∎
7.2. Proof of Theorem 1.14
Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $\geq 2$. We aim to study the condition $(1)$ $f$ is PCF. Our goal is to prove that $(1)$ is equivalent to each of the conditions $(2)$ and $(3)$ in Theorem 1.14, which are determined by the multiplier spectrum and the length spectrum of $f$, respectively.
Using the terminology and notations in §5, it is clear that $(2)$ and $(3)$ are equivalent to the following conditions $(2)^{\prime}$ and $(3)^{\prime}$, respectively.
$(2)^{\prime}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x)))$ for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
$(3)^{\prime}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $\mathbf{n}_{\mathbb{Q}}({\rm rog}(\lvert\rho_{f}(x)\lvert))$ for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
Now we prove that the three conditions $(1),(2)^{\prime}$, and $(3)^{\prime}$ are equivalent.
$(1)\Rightarrow$ $(2)^{\prime}$ and $(3)^{\prime}$:
Suppose that $f$ is PCF. By [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in $\mathcal{M}_{d}$, except for the family of flexible Lattès maps. If $f$ is flexible Lattès, then according to [18, Lemma 5.6], we have $\rho_{f}(x)\in\mathbb{Z}$ for all $x\in{\rm Per}(f)(\mathbb{C})$. If $f$ is defined over $\overline{\mathbb{Q}}$, then clearly $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})={\rm Per}(f)(\mathbb{Q})$. Thus, we always have $\rho_{f}(x),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
Suppose that $(2)^{\prime}$ is false, then
$\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}=\infty.$
By [18, Corollary 3.9], $f$ cannot be a flexible Lattès map, hence $f$ is defined over $\overline{\mathbb{Q}}$, and over a number field $K$. We use the notation and ideas from §6.1 in the case where $\tau=\mathrm{id}$. Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all the archimedean places such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$. For every $v\in\mathcal{M}_{K}\setminus B$, the reduction $f_{v}$ is still PCF and its critical orbits come from those of $f$. Arguing similarly as in §6.1, it is easy to see that the set
$\mathcal{W}:=\{x\in{\rm Per}^{*}(f)(\mathbb{C}):\phi_{v}(\mathbf{n}_{K}(\lambda(x)))=0,\forall v\in\mathcal{M}_{K}\setminus B\}$
is co-finite in ${\rm Per}^{*}(f)(\mathbb{C})$, i.e., ${\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}$ is finite. It is well-known that ${\rm rank}(\mathcal{O}_{K,B}^{\times})=|B|-1<\infty$ (cf. [21, Theorem 3.12]). Note that
$\mathbf{n}_{K}({\rm rog}(\rho_{f}(x)))\in\mathbb{D}(\mathcal{O}_{K,B})$
for all $x\in{\rm Per}^{*}(f)(\mathbb{C})$ by [24, Corollary 2.23 (a)]. We deduce that
$\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{K}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}\leq|{\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}|+|B|-1<\infty,$
which implies $\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}<\infty$, contradicting the assumption. Thus $(2)^{\prime}$ must hold.
$(3)^{\prime}$ follows from an argument similar to the above paragraph, corresponding to the case where $\tau$ is the complex conjugation of §6.1.
$(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$:
Let $\tau$ be the identity on $\mathbb{C}$ in the case $(2)^{\prime}$ and be the complex conjugation in the case $(3)^{\prime}$. We will show $(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$ simultaneously. Suppose that $f$ is not PCF. In particularly, $f$ is not flexible Lattès. By $(2)^{\prime}$ (or $(3)^{\prime}$) and Theorem 1.2, the map $f$ is defined over $\overline{\mathbb{Q}}$, hence over a number field $K$. We use the notation and ideas from §6.1 for $\tau$. Arguing similarly as in §6.1, after enlarging the number field $K$, we can take a finite set $B\subseteq\mathcal{M}_{K}$ of places containing all the archimedean places satisfying the following conditions:
  • $K/\mathbb{Q}$ is Galois and $\tau(K)=K$, hence $\tau|_{K}\in{\rm Gal}(K/\mathbb{Q})$;
  • $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$;
  • $B$ is invariant under every $\sigma\in{\rm Gal}(K/\mathbb{Q})$;
  • $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$, where $o$ is a critical point of $f$ that is not $f$-preperiodic and $X$ is the union of all $f$-orbits of periodic critical points of $f$. (Here $o$ and all points in $X$ are assumed to be defined over $K$.)
A small modification of the proof of Lemma 6.2 shows that there exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ satisfying the following conditions:
(7.15) $\phi_{v_{i}}(\mathbf{n}_{K}(\lambda(x_{i})))>0\text{ for all }i\geq 1;$
(7.16) $\phi_{\sigma(v_{j})}(\mathbf{n}_{K}(\lambda(x_{i})))=0\text{ for all }i\neq j\text{ and }\sigma\in{\rm Gal}(K/\mathbb{Q}).$
For $i\geq 1$, let $p_{i}$ be the rational prime below $v_{i}$ and $\widetilde{B}\subseteq\mathcal{M}_{\mathbb{Q}}$ be the set of places of $\mathbb{Q}$ lying below the places in $B$. From (7.15) and (7.16), it is easy to see that the pair $((\mathbf{n}_{\mathbb{Q}}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{p_{i}})_{i=1}^{\infty})$ is an upper triangle $(\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$. Note that $\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ is $\tau$-invariant, where $\tau$ descends to $\mathbb{D}(\mathbb{Q})$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$. By Corollary 5.5 and Remark 5.6, this contradicts $(2)^{\prime}$ and $(3)^{\prime}$, respectively. Thus, $f$ must be PCF.
Acknowledgement
The second-named author Junyi Xie would like to thank Thomas Gauthier, Gabriel Vigny, Charles Favre, and Serge Cantat for helpful discussions. The authors would like to thank the anonymous referee for many thoughtful comments, which help to improve the manuscript.
The first-named author would like to thank Beijing International Center for Mathematical Research in Peking University for the invitation. The second and third-named authors Junyi Xie and Geng-Rui Zhang are supported by NSFC Grant (No.12271007).
Declarations
Data availability
Data sharing is not applicable to this article because no data were generated or analyzed during the current study.
Conflict of interest
The authors declare that they have no conflict of interest.

### v3

@@
 For example, we show that for every non-exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$, the $\mathbb{Q}$-vector space generated by all the (finite) characteristic exponents of periodic points of $f$ has infinite dimension.
-This answers a stronger version of a question of Levy and Tucker.
+This answers a stronger version of a question by Levy and Tucker.
 Our result can also be seen as a generalization of recent results of Ji-Xie and of Huguin which proved Milnor’s conjecture about rational maps having integer multipliers.
-We also get a characterization of postcritically finite maps by using its length spectrum.
+We also get a characterization of postcritically finite maps by using their length spectra.
 Finally as an application of our result, we get a new proof of the Zariski-dense orbit conjecture for endomorphisms on $(\mathbb{P}^{1})^{N},N\geq 1$.
@@
 Let $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ be a periodic point of $f$ with exact period $n$.
-Define $n_{f}(z_{0}):=n$ be this period.
+Define $n_{f}(z_{0}):=n$.
 We write $n(z_{0})$ for simplicity when the map $f$ is clear.
@@
 We write $\rho(z_{0})$ for simplicity when the map $f$ is clear.
-The length of $f$ at $z_{0}$ is the norm $|\rho_{f}(z_{0})|$.
-The multiplier and the length are invariant under conjugacy.
+The length of $f$ at $z_{0}$ is the absolute value $|\rho_{f}(z_{0})|$.
+The multiplier and length are invariant under conjugacy.
 The characteristic exponent of $f$ at $z_{0}$ is defined to be $\chi_{f}(z_{0}):=n^{-1}\log\lvert\rho_{f}(z_{0})\rvert$ (when $\rho_{f}(z_{0})\neq 0$).
@@
 where $\mu_{f}$ is the unique maximal entropy measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ which is also the canonical measure of $f$ on $\mathbb{P}^{1}(\mathbb{C})$ (see [14, 7, 15, 34]), and the norm of the differential is computed with respect to the spherical metric.
-In fact, the Lyapunov exponent dose not change with the spherical metric replaced by any Kähler metric on $\mathbb{P}^{1}_{\mathbb{C}}$ in the definition.
+In fact, the value of the Lyapunov exponent does not change with the spherical metric replaced by any Kähler metric on $\mathbb{P}^{1}_{\mathbb{C}}$ in the definition.
 It is known that $\mathcal{L}_{f}\geq\log(d)/2$, and the equality holds if and only if $f$ is a Lattès map in the sense of Definition 1.1 (see [34, 35]).
-The quantity $\exp(\mathcal{L}_{f})$ is considered as an average of $|df|$ on $\mathbb{P}^{1}(\mathbb{C})$, which measures the average rate of expansion of $f$ along a typical orbit.
+The quantity $\exp(\mathcal{L}_{f})$ is considered as an average of $|df|$ on $\mathbb{P}^{1}(\mathbb{C})$, which measures the average expansion rate of $f$ along a typical orbit.
 1.2. Exceptional maps
@@
   • The map $f$ is called Lattès if there is an endomorphism $\phi$ on an elliptic curve $E$ such that $\phi$ is semi-conjugate to $f$.
-Here, for an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $g$ is semi-conjugate to $h$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$.
-Furthermore, the map $f$ is called flexible Lattès if there are an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that the multiplication $[n]$ on $E$ is semi-conjugate to $f$ via the quotient map modulo {±$\{\pm 1\}$ on $E$.
+Here, for an endomorphism $g$ on an algebraic variety $X$ and an endomorphism $h$ on an algebraic variety $Y$, we say $h$ is semi-conjugate to $g$ if there is a dominant morphism $\pi:X\to Y$ such that $\pi\circ g=h\circ\pi$.
+Furthermore, the map $f$ is called flexible Lattès if there are an elliptic curve $E$ and $n\in\mathbb{Z}\setminus\{0,\pm 1\}$ such that $f$ is semi-conjugate to the multiplication-by-$n$ map $[n]$ on $E$ via the quotient map modulo $\{\pm 1\}$ on $E$.
 Otherwise, it is called rigid Lattès.
-  • We say that $f$ is of monomial type if it semi-conjugate to the map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$.
+  • We say that $f$ is of monomial type if it is semi-conjugate to the map $z\mapsto z^{n}$ on $\mathbb{P}^{1}$ for some integer $n\neq 0,\pm 1$.
 We call $f$ exceptional if it is Lattès or of monomial type. It is well-known that $f$ is exceptional if and only if some iterate $f^{k}$ is exceptional ($k\in\mathbb{Z}_{>0}$).
@@
 We fix an embedding of the algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ and consider all number fields as subfields of $\mathbb{C}$.
-Denote the usual absolute value on $\mathbb{C}$ by ||$\cdot$|| ⋅ |.
+Denote the usual absolute value on $\mathbb{C}$ by $|\cdot|$.
 Our first result shows that the definition field of a rational map is determined by its length spectrum, provided that it is not a flexible Lattès map.
@@
 Remark 1.3.
-In Theorem 2.1, we prove a more general version of Theorem 1.2, in which $\overline{\mathbb{Q}}$ can be replaced by any algebraically closed subfield of $\mathbb{C}$ invariant under the complex conjugation.
-In fact, the actual proof shows that we can weaken the conditions in Theorems 1.2 and 2.1 by discarding the corresponding condition for finitely many periodic points of $f$, see Remark 2.2.
-For example, in Theorem 1.2, we can only require that $|\rho_{f}(x)|\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$, and the conclusion still holds.
+In Theorem 2.1, we prove a more general version of Theorem 1.2, in which $\overline{\mathbb{Q}}$ can be replaced by any algebraically closed subfield of $\mathbb{C}$ invariant under complex conjugation.
+In fact, the actual proof shows that we can weaken the conditions in Theorems 1.2 and 2.1 by discarding the corresponding condition for finitely many periodic points of $f$; see Remark 2.2.
+For example, in Theorem 1.2, we can require only that $|\rho_{f}(x)|\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$, and the conclusion still holds.
 McMullen’s rigidity of multiplier spectrum [16, Corollary 2.3] asserts that multipliers at all periodic points determine the conjugacy class of rational maps up to only finitely many choices, except for the flexible Lattès family.
-With a standard spread out argument, McMullen’s rigidity theorem implies that, for a rational map $f$ of degree at least $2$ which is not flexible Lattès, if its multipliers at periodic points are all algebraic, then $f$ is defined over $\overline{\mathbb{Q}}$.
+With a standard spread-out argument, McMullen’s rigidity theorem implies that, for a rational map $f$ of degree at least $2$ which is not flexible Lattès, if its multipliers at periodic points are all algebraic, then $f$ is defined over $\overline{\mathbb{Q}}$.
 We sketch the argument as follows.
 Let $f$ be a rational map over $\mathbb{C}$ of degree $d\geq 2$ that is not flexible Lattès with all multipliers in $\overline{\mathbb{Q}}$.
-The noetherianity implies that the set $D$ of conjugate classes $[g]$ (in the moduli space $\mathcal{M}_{d}$) such that $g$ has the same multiplier spectrum (i.e., all multipliers at periodic points) as $f$ is an algebraic subset of $\mathcal{M}_{d}$ defined over $\overline{\mathbb{Q}}$, since the multiplier spectrum morphism is defined over $\mathbb{Q}$ and all multipliers of $f$ are in $\overline{\mathbb{Q}}$.
+Noetherianity implies that the set $D$ of conjugacy classes $[g]$ (in the moduli space $\mathcal{M}_{d}$) such that $g$ has the same multiplier spectrum (i.e., all multipliers at periodic points) as $f$ is an algebraic subset of $\mathcal{M}_{d}$ defined over $\overline{\mathbb{Q}}$, since the multiplier spectrum morphism is defined over $\mathbb{Q}$ and all multipliers of $f$ are in $\overline{\mathbb{Q}}$.
 Then McMullen’s rigidity implies that $\dim(D)=0$, hence $D$ is a set of finitely many $\overline{\mathbb{Q}}$-points.
-Thus $f$ is defined over $\overline{\mathbb{Q}}$ as $[f]\in D$.
+Thus $f$ is defined over $\overline{\mathbb{Q}}$ as $f]\in D$.
 Theorem 1.2 is a generalization of this result from multiplier spectrum to length spectrum (which contains less information).
 The rigidity of length spectrum was proved by Ji-Xie [12, Theorem 1.5].
-However, the spread out argument does not apply directly in this case as the length spectrum map (and its square) is not algebraic on the moduli space of rational maps.
-Indeed as shown in [12, §8.1], its square is not even real algebraic (but it is still semialgebraic).
-In §2.3, we introduce a method to do the spread out argument respecting the structure of $\mathcal{M}_{d}(\mathbb{C})$ as a real algebraic variety using Weil restriction.
+However, the spread-out argument does not apply directly in this case as the length spectrum map (and its square) is not algebraic on the moduli space of rational maps.
+Indeed, as shown in [12, §8.1], its square is not even real algebraic (though it is still semialgebraic).
+In §2.3, we introduce a method to do the spread-out argument respecting the structure of $\mathcal{M}_{d}(\mathbb{C})$ as a real algebraic variety using Weil restriction.
 Another difficulty in the length spectrum case is the lack of noetherianity for semialgebraic subsets.
-We overcome this difficulty using the notion of admissible subsets introduced in [12].
+We overcome this difficulty by using the notion of admissible subsets introduced in [12].
 The following two results concern the $\mathbb{Q}$-vector space spanned by all the (finite) characteristic exponents of periodic points.
@@
 Remark 1.6.
-With the help of Remark 1.3, the proof of Theorem 1.5 in §4 indeed shows that we can weaken the condition of Theorem 1.5 by only requiring that there is a number field $K$ such that for all but finitely many $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n(z_{0})}\in K$ for some $n(z_{0})\in\mathbb{Z}_{>0}$.
-Finitely many nonzero elements $z_{1},\dots,z_{N}$ in a field $F$ are called multiplicatively independent if for all triples $(m_{1},\dots,m_{N})$ of integers, $z_{1}^{m_{1}}\cdots z_{N}^{m_{N}}=1$ if and only if $m_{1}=\cdots=m_{N}=0$.
+With the help of Remark 1.3, the proof of Theorem 1.5 in §4 indeed shows that we can weaken the condition of Theorem 1.5 by requiring only that there is a number field $K$ such that for all but finitely many $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n(z_{0})}\in K$ for some $n(z_{0})\in\mathbb{Z}_{>0}$.
+Finitely many nonzero elements $z_{1},\dots,z_{N}$ in a field $F$ are called multiplicatively independent if for every $N$-tuple $(m_{1},\dots,m_{N})$ of integers, $z_{1}^{m_{1}}\cdots z_{N}^{m_{N}}=1$ if and only if $m_{1}=\cdots=m_{N}=0$.
 A sequence $(z_{n})_{n=1}^{\infty}$ in $F\setminus\{0\}$ is called multiplicatively independent if all its finite subsequences are multiplicatively independent.
@@
 1.4.1. Milnor’s conjecture
-Milnor [18] has showed that an exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ must have all its multipliers of periodic points in the ring of integers $\mathcal{O}_{K}$ for some imaginary quadratic number field $K$, and in fact in $\mathbb{Z}$ when $f$ is not a rigid Lattès map.
+Milnor [19] showed that an exceptional rational map $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ must have all its multipliers of periodic points in the ring of integers $\mathcal{O}_{K}$ for some imaginary quadratic number field $K$, and in fact in $\mathbb{Z}$ when $f$ is not a rigid Lattès map.
 Milnor conjectured that the converse is also true.
@@
 Then $f$ is exceptional.
-See also [5] for a different proof. Recently Huguin generalized the above result using different approach:
+See also [5] for a different proof. Recently, Huguin generalized the above result using a different approach:
 Theorem 1.9 ([10, Theorem 7]).
@@
 1.4.2. A question of Levy and Tucker
-On the other hand, in the 2014 AIM workshop Postcritically Finite Maps In Complex And Arithmetic Dynamics, Levy [13] and Tucker [26] asked the following question independently:
+On the other hand, at the 2014 AIM workshop Postcritically Finite Maps In Complex And Arithmetic Dynamics, Levy [13] and Tucker [26] independently asked the following question:
 Question 1.10.
-Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a non-exceptional rational map of degree $d\geq 2$ and let $S$ be the set of all multipliers of periodic points of $f$.
+Let $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ be a non-exceptional rational map of degree $d\geq 2$, and let $S$ be the set of all multipliers of periodic points of $f$.
 Take the subgroup of $\mathbb{C}^{\ast}$ generated by $S\setminus\{0\}$.
-Is that true that the rank of this group is infinite?
+Is it true that the rank of this group is infinite?
 It is not hard to see that our Corollary 1.7 gives a positive answer to (a generalized version of) Levy and Tucker’s question.
 1.5. Sketch of the proofs
-We have given the idea of Theorem 1.2 before. Here we explain the proofs of Theorem 1.4 and Theorem 1.5.
+We have already given the idea of Theorem 1.2. Here we explain the proofs of Theorem 1.4 and Theorem 1.5.
 We first give the idea of the proof of Theorem 1.5.
@@
 The first step is to reduce to the case where $f$ is defined over $\overline{\mathbb{Q}}$, which can be done using our Theorem 1.2.
-After enlarging $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ is the rational map obtained from $f$ via replacing the coefficients by their complex conjugations.
+After enlarging $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ is the rational map obtained from $f$ by replacing the coefficients with their complex conjugates.
 In the second step, we combine an arithmetic equidistribution theorem with a result of Zdunik [35] (see [10]) on the Lyapunov exponent to get a contradiction.
@@
 Without loss of generality, we may assume that $\Gamma$ is irreducible and that the sequence $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$.
-Since the Dynamical Manin-Mumford problem is solved for such an endomorphism $F$ by [8] and $\Gamma$ contains a Zariski-dense subset $\{p_{n}:n\geq 1\}$ of $F$-periodic points, the subscheme $\Gamma$ is $F$-periodic itself.
-Hence we may assume that $\Gamma$ is $F$-invariant after replacing $F$ by a suitable iteration.
+Since the Dynamical Manin-Mumford problem is solved for such an endomorphism $F$ by [8] and $\Gamma$ contains a Zariski-dense subset $\{p_{n}:n\geq 1\}$ of $F$-periodic points, the subscheme $\Gamma$ itself is $F$-periodic.
+Hence, after replacing $F$ by a suitable iteration, we may assume that $\Gamma$ is $F$-invariant.
 Let $\nu_{n}$ be the discrete probability measure equally supported at the union of Galois orbits of iterates of $p_{n}$ under $F$.
-Then $\nu_{n}$ converges weakly to the canonical measure $\mu$ on $\Gamma$ with respect to $F$ by an equidistribution-type theorem (Theorem 3.1), which is a reformulation of [32, Theorem 3.1], see §3 for details.
+Then $\nu_{n}$ converges weakly to the canonical measure $\mu$ on $\Gamma$ with respect to $F$ by an equidistribution-type theorem (Theorem 3.1), which is a reformulation of [32, Theorem 3.1]; see §3 for details.
 Applying $\nu_{n}\to\mu$ to the continuous test functions $\max\{\log\lvert\det(dF)\rvert,A\}$ ($A\in\mathbb{R}$) and letting $A\to$-$\infty$, we get
-[…]
-which is impossible since the right hand side equals to $2\mathcal{L}_{f}<2a$ by a direct computation.
-Next we sketch the proof of Theorem 1.4.
+$2a\leq\int\log\lvert\det(dF)\rvert d\mu,$
+which is impossible since the right-hand side equals $2\mathcal{L}_{f}<2a$ by a direct computation.
+Next, we sketch the proof of Theorem 1.4.
 According to [6], postcritically finite (PCF) maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps.
@@
 $f$ is not PCF.
-For the first case the conclusion follows from Theorem 1.5.
+For the first case, the conclusion follows from Theorem 1.5.
 For the second case, we need to develop some new techniques, which are presented in §5.
-In §5, we consider some pseudo-linear algebra (which means that the domain of a map may not be the whole vector space), and the vector space $\mathbb{D}(k)_{\mathbb{Q}}=k^{*}\otimes_{\mathbb{Z}}\mathbb{Q}$ for a field $k$ of characteristic zero.
-We will actually prove a theorem (Theorem 5.7) stronger than the non-PCF case of Theorem 1.4, see §5 and §6 for details.
+In §5, we consider some pseudo-linear algebra (meaning that the domain of a map may not be the whole vector space) and the vector space $\mathbb{D}(k)_{\mathbb{Q}}=k^{*}\otimes_{\mathbb{Z}}\mathbb{Q}$ for a field $k$ of characteristic zero.
+We will actually prove a theorem (Theorem 5.7) stronger than the non-PCF case of Theorem 1.4; see §5 and §6 for details.
 To prove Theorem 5.7, in §6.1 we first deal with the case that $f$ is defined over $\overline{\mathbb{Q}}$.
-A key ingredient in this step is [3, Lemma 4.1] which is a consequence of Siegel’s theorem on $S$-integral points.
+A key ingredient in this step is [3, Lemma 4.1], which is a consequence of Siegel’s theorem on $S$-integral points.
 The existence of a non-preperiodic critical point of $f$ is essentially used here.
-In §6.2, we consider the general case and finish the proof.
-This is achieved by reducing to the case that $f$ is defined over $\overline{\mathbb{Q}}$ via an algebraic-geometric argument and techniques developed in §5.
+In §6.2, we consider the general case and complete the proof.
+This is achieved by reducing to the case that $f$ is defined over $\overline{\mathbb{Q}}$ via an algebraic-geometric argument and the techniques developed in §5.
 1.6. Applications
 1.6.1. The Zariski-dense orbit conjecture
-By applying Corollary 1.7 we can give a new proof of a special case of the Zariski-dense orbit conjecture.
+By applying Corollary 1.7, we can give a new proof of a special case of the Zariski-dense orbit conjecture.
 Zariski-dense orbit conjecture (=ZDO).
-Let $k$ be an algebraically closed field of characteristic 00.
+Let $k$ be an algebraically closed field of characteristic 0.
 Given an irreducible quasi-projective variety $X$ over $k$ and a dominant rational self-map $f$ on $X$.
-If we have $\{g\in k(X):g\circ f=g\}=k$ where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
+If $\{g\in k(X):g\circ f=g\}=k$, where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
 Remark 1.11.
-The converse of ZDO holds and is easy. For some progress on ZDO, see [1, 2, 17, 27, 29] for example.
+The converse of ZDO holds and is easy to prove. For some progress on ZDO, see for example [1, 2, 17, 27, 29].
 As an application of Corollary 1.7, we give a new proof of (the most difficult part of) a special case of ZDO, which was first proved by Xie [29, Theorem 1.16].
 Theorem 1.12.
-Let $X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}$ be the variety of product of $N$ copies of projective line over an algebraically closed field $k$ of characteristic 00.
-Suppose that $f:X\to X$ is an endomorphism of the form $f_{1}\times\cdots\times f_{N}$ where $f_{j}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ is a non-constant rational map for $1\leq j\leq N$.
-The ZDO holds for $(X,f)$.
+Let $X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}$ be the product of $N$ copies of the projective line over an algebraically closed field $k$ of characteristic 0.
+Suppose that $f:X\to X$ is an endomorphism of the form $f_{1}\times\cdots\times f_{N}$, where $f_{j}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ is a non-constant rational map for $1\leq j\leq N$.
+Then the ZDO holds for $(X,f)$.
 Remark 1.13.
 We note that every dominant endomorphism $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ over an algebraically closed field $k$ of characteristic zero must have the form $f_{1}\times\cdots\times f_{N}$, after replacing $f$ by a suitable iterate.
-The original proof of Theorem 1.12 in [29] relies on the solution of the (adelic) Zariski-dense orbit conjecture on smooth projective surfaces [29, Theorem 1.15], the notion of adelic topology introduced in [29, §3] and a classification result [29, Proposition 9.2] on invariant subvarieties of $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ (see also [17, 8]).
-When $n=2$, Pakovich gave another proof [22] of Theorem 1.12 using his classification of invariant curves in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and some height argument.
-We give a new proof in the case where all the $f_{i}$’s are non-exceptional of degree at least $2$, and this proof does not need the ingredients mentioned above.
-Other cases of Theorem 1.12 are relatively easy or can be reduced to the case where $f_{i}$ are all non-exceptional, as shown in [29, §9.3].
+The original proof of Theorem 1.12 in [29] relies on the solution of the (adelic) Zariski-dense orbit conjecture for smooth projective surfaces [29, Theorem 1.15], the notion of adelic topology introduced in [29, §3], and a classification result [29, Proposition 9.2] on invariant subvarieties of $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ (see also [17, 8]).
+When $N=2$, Pakovich gave another proof [22] of Theorem 1.12 using his classification of invariant curves in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and a height argument.
+We give a new proof in the case where all the $f_{i}$’s are non-exceptional of degree at least $2$; this proof does not require the ingredients mentioned above.
+Other cases of Theorem 1.12 are relatively easy or can be reduced to the case where all $f_{i}$ are non-exceptional, as shown in [29, §9.3].
 1.6.2. A characterization of PCF maps
@@
 $L_{x}$) is any number field containing $\rho_{f}(x)$ (resp.
-$\lvert\rho_{f}(x)\rvert$) and $N_{K_{x}/\mathbb{Q}}$ (resp.
+$\lvert\rho_{f}(x)\rvert$), and $N_{K_{x}/\mathbb{Q}}$ (resp.
 $N_{L_{x}/\mathbb{Q}}$) is the norm map for the extension $K_{x}/\mathbb{Q}$ (resp.
@@
 $|\rho_{f}(x)|$).
-Clearly, the subspaces $V,W$ above are independent of the choices of the fields $K_{x},L_{x}$, respectively.
-We also remark that the $\mathbb{Q}$-vector spaces $V$ in $(2)$ and $W$ in $(3)$ are the same space.
+Clearly, the subspaces $V$ and $W$ above are independent of the choices of the fields $K_{x}$ and $L_{x}$, respectively.
+We also remark that the $\mathbb{Q}$-vector spaces $V$ in (2) and $W$ in (3) are the same.
 Remark 1.15.
-Similarly, the actual proof shows that we can weaken the conditions $(2)$ and $(3)$ in Theorem 1.14, in the spirit of Remarks 1.3 and 1.6.
-Precisely, we can add two more equivalent statements in Theorem 1.14:
-$(2)^{*}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all but finitely $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $V^{*}=V^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $V^{*}$ is generated over $\mathbb{Q}$ by
+Similarly, the actual proof shows that we can weaken conditions (2) and (3) in Theorem 1.14, in the spirit of Remarks 1.3 and 1.6. Precisely, we can add two more equivalent statements to Theorem 1.14:
+$(2)^{*}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $V^{*}=V^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $V^{*}$ is generated over $\mathbb{Q}$ by
 $\{\log\lvert N_{K_{x}/\mathbb{Q}}(\rho_{f}(x))\rvert:x\in{\rm Per}^{*}(f)(\mathbb{C}),\rho_{f}(x)\in\overline{\mathbb{Q}}\};$
-$(3)^{*}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all but finitely $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $W^{*}=W^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $W^{*}$ is generated over $\mathbb{Q}$ by
+$(3)^{*}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $W^{*}=W^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $W^{*}$ is generated over $\mathbb{Q}$ by
 $\{\log\lvert N_{L_{x}/\mathbb{Q}}(\lvert\rho_{f}(x)\rvert)\rvert:x\in{\rm Per}^{*}(f)(\mathbb{C}),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}\}$.
@@
 2. Rational maps with algebraic lengths
-Let $K$ be an algebraically closed subfield of $\mathbb{C}$ which is invariant under the complex conjugation $\tau$, i.e., $\tau(K)=K$. The aim of this section is to prove the following theorem.
+Let $K$ be an algebraically closed subfield of $\mathbb{C}$ which is invariant under complex conjugation $\tau$, i.e., $\tau(K)=K$. The aim of this section is to prove the following theorem.
 Theorem 2.1.
 Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$.
-Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $\lvert\rho_{f}(x)\rvert\in K$.
+Assume that $f$ is not a flexible Lattès map and that for every $x\in{\rm Per}(f)(\mathbb{C})$, $\lvert\rho_{f}(x)\rvert\in K$.
 Then $f$ is defined over $K$.
 Remark 2.2.
-In fact, we can weaken the condition of Theorem 2.1 by only requiring that $f$ is not flexible Lattès and $\lvert\rho_{f}(x)\rvert\in K$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$.
-To show this, we need to replace $A$ in the proof in §2.7 by $A=(A_{n}=(RL^{*}(f)_{n})\cap K))_{n}$ and then the original proof works.
-Here, $A_{n}=(RL^{*}(f)_{n})\cap K$ is the sub-multiset of the multiset $RL^{*}(f)_{n}$ of all elements also in $K$, see §2.6 and §2.7.
+The condition in Theorem 2.1 can be weakened: it suffices to require that $f$ is not flexible Lattès and $\lvert\rho_{f}(x)\rvert\in K$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$.
+To show this, we need to replace $A$ in the proof in §2.7 by $A=(A_{n}=(RL^{*}(f)_{n})\cap K)_{n}$, and then the original proof applies.
+Here, $A_{n}=(RL^{*}(f)_{n})\cap K$ is the sub-multiset of the multiset $RL^{*}(f)_{n}$ consisting of elements that also lie in $K$; see §2.6 and §2.7.
 Applying Theorem 2.1 to the case $K=\overline{\mathbb{Q}}$, we get Theorem 1.2.
-We will introduce the notions of Weil restriction, admissible subsets, and transcendental points in §2.1-2.3.
-Then we construct certain moduli spaces and study the length map on them in §2.4-2.6.
-After these preparations, we finish the proof of Theorem 2.1 in §2.7.
+We introduce the notions of Weil restriction, admissible subsets, and transcendental points in §2.12.3.
+Then we construct certain moduli spaces and study the length map on them in §2.42.6.
+After these preparations, we complete the proof of Theorem 2.1 in §2.7.
 2.1. Weil restriction
@@
 Lemma 2.3.
-[…]
+We have $K=L+iL$; in particular $[K:L]=2$.
 Proof of Lemma 2.3.
@@
 $u=\frac{u+\tau(u)}{2}+\frac{u-\tau(u)}{2i}i$
-and both $\frac{u+\tau(u)}{2}$ and $\frac{u-\tau(u)}{2i}$ are contained in $L$. This concludes the proof. ∎
-In order to prove Theorem 2.1, we need to analyze the length map respecting the structure of $\mathcal{M}_{d}(K)$ as an algebraic variety over $L$ using Weil restriction, due to the appearance of the absolute value in the condition.
+and both $\frac{u+\tau(u)}{2}$ and $\frac{u-\tau(u)}{2i}$ are contained in $L$. This completes the proof. ∎
+To prove Theorem 2.1, we need to analyze the length map while respecting the structure of $\mathcal{M}_{d}(K)$ as an algebraic variety over $L$ using Weil restriction, due to the appearance of the absolute value in the condition.
 Hence, the notion of Weil restriction, which we will briefly recall, is necessary.
-See [23, §4.6] and [4, §7.6] for more information.
+See [23, §4.6] and [4, §7.6] for more details.
 Denote by ${\rm Var}_{/K}$ (resp.
@@
 The functor $X\mapsto R(X)$ is called the Weil restriction.
-One has the canonical bijection $\psi_{K}:X(K)\to R(X)(L)$ between sets.
+There is a canonical bijection $\psi_{K}:X(K)\to R(X)(L)$ between sets.
 When $K=\mathbb{C}$, this map is a real analytic diffeomorphism.
@@
 A subset $Y$ of $X(K)$ is called $L$-algebraic if it is closed in the $L$-Zariski topology.
-When $K=\mathbb{C}$, the $L$-Zariski topology is exactly the real Zariski topology as in [12, §8.1.1].
-By the last statement of Proposition 2.6 below, the $L$-Zariski topology is stronger than the Zariski topology on $X(K)$.
+When $K=\mathbb{C}$, the $L$-Zariski topology coincides with the real Zariski topology as in [12, §8.1.1].
+By the last statement of Proposition 2.6 below, the $L$-Zariski topology is finer than the Zariski topology on $X(K)$.
 When $K=\mathbb{C}$, roughly speaking, the Weil restriction is just constructed by splitting a complex variable $z$ into two real variables $x,y$ via $z=x+iy$.
@@
 $\mathbb{B}[x_{1},y_{1},\dots,x_{N},y_{N}]=L[x_{1},y_{1},\dots,x_{N},y_{N}]\oplus IL[x_{1},y_{1},\dots,x_{N},y_{N}],$
-$F$ can be uniquely decomposed to
+$F$ can be uniquely decomposed as
 $F=r(f)+Ii(f)$
@@
 Proposition 2.6.
-Let $X,Y\in{\rm Var}_{/K}$, then we have the following properties:
+Let $X,Y\in{\rm Var}_{/K}$. The following properties hold:
   • if $X$ is irreducible, then $R(X)$ is irreducible;
@@
 Example 2.7.
-If $X$ is the subvariety of $\mathbb{A}^{N}_{K}={\rm Spec}K[z_{1},\dots,z_{N}]$ defined by the equations $\sum_{I}a_{i,I}z^{I}=0$, $i=1,\dots,s$.
-Then $X^{\tau}$ is the subvariety of $\mathbb{A}^{N}_{K}$ defined by $\sum_{I}\tau(a_{i,I})z^{I}=0$, $i=1,\dots,s$.
+If $X$ is the subvariety of $\mathbb{A}^{N}_{K}={\rm Spec}K[z_{1},\dots,z_{N}]$ defined by the equations $\sum_{I}a_{i,I}z^{I}=0$ ($i=1,\dots,s$), then $X^{\tau}$ is the subvariety of $\mathbb{A}^{N}_{K}$ defined by $\sum_{I}\tau(a_{i,I})z^{I}=0$ ($i=1,\dots,s$).
 The map $\tau:X=(X^{\tau})^{\tau}\to X^{\tau}$ sends a point $(z_{1},\dots,z_{N})\in X(K)$ to $(\tau(z_{1}),\dots,\tau(z_{N}))\in X^{\tau}(K)$.
 The following result due to Weil is useful for computing the Weil restriction.
-Proposition 2.8.
-[23, Exercise 4.7] We have a canonical isomorphism
+Proposition 2.8 ([23, Exercise 4.7]).
+There is a canonical isomorphism
 $R(X)\otimes_{L}K\simeq X\times X^{\tau}.$
 Under this isomorphism,
-$R(X)(L)=\{(z_{1},z_{2})\in X(K)\times X^{\tau}(K)|\,\,z_{2}=\tau(z_{1})\}$
+$R(X)(L)=\{(z_{1},z_{2})\in X(K)\times X^{\tau}(K)\mid z_{2}=\tau(z_{1})\}$
 and $\psi_{K}$ sends $z\in X(K)$ to $(z,\tau(z))\in R(X)(L)$.
 2.2. Admissible subsets
-In this subsection, we recall the notion of admissible subsets of real algebraic varieties introduced in [12], which satisfies the important decreasing chain condition.
+In this subsection, we recall the notion of admissible subsets of real algebraic varieties introduced in [12], which satisfies the important descending chain condition.
 Let $X$ be a variety over $\mathbb{R}$.
-Definition 2.9.
-[12, §8.2] A closed subset $V$ of $X(\mathbb{R})$ (with the real topology) is called admissible if there is a morphism $f:Y\to X$ of real algebraic varieties and a Zariski closed subset $V^{\prime}\subseteq Y$ such that $V=f(V^{\prime}(\mathbb{R}))$ and $f$ is étale at every point in $V^{\prime}(\mathbb{R})$.
+Definition 2.9 ([12, §8.2]).
+A closed subset $V$ of $X(\mathbb{R})$ (with the real topology) is called admissible if there is a morphism $f:Y\to X$ of real algebraic varieties and a Zariski closed subset $V^{\prime}\subseteq Y$ such that $V=f(V^{\prime}(\mathbb{R}))$ and $f$ is étale at every point in $V^{\prime}(\mathbb{R})$.
 In particular, every algebraic subset of $X(\mathbb{R})$ is admissible.
 Remark 2.10.
-Denote by $J$ the non-étale locus for $f$ in $Y$.
-We have $J\cap V^{\prime}(\mathbb{R})=\emptyset$.
+Denote by $J$ the non-étale locus of $f$ in $Y$.
+Then $J\cap V^{\prime}(\mathbb{R})=\emptyset$.
 Since we may replace $(Y,V^{\prime})$ by $(Y\setminus J,V^{\prime}\cap(Y\setminus J))$, in the above definition we may further assume that $f$ is étale.
-Proposition 2.11.
-[12, Remarks 8.14, 8.15, and Proposition 8.16] We have the following basic properties:
+Proposition 2.11 ([12, Remarks 8.14, 8.15, and Proposition 8.16]).
+The following basic properties hold:
   • (1)
@@
 The following theorem shows that admissible subsets satisfy the descending chain condition.
-Theorem 2.12.
-[12, Theorem 8.17] Let $(V_{n})_{n\geq 1}$ be a sequence of decreasing admissible subsets of $X(\mathbb{R})$. Then there is $N\geq 1$ such that $V_{n}=V_{N}$ for all $n\geq N$.
+Theorem 2.12 ([12, Theorem 8.17]).
+Let $(V_{n})_{n\geq 1}$ be a decreasing sequence of admissible subsets of $X(\mathbb{R})$. Then there exists $N\geq 1$ such that $V_{n}=V_{N}$ for all $n\geq N$.
 2.3. Transcendental points
-Let $X_{K}$ be a variety over $K$ and $X:=X_{K}\otimes_{K}\mathbb{C}$. We consider $X_{K}$ as a model of $X$ over $K$.
+Let $X_{K}$ be a variety over $K$ and $X:=X_{K}\otimes_{K}\mathbb{C}$. We regard $X_{K}$ as a model of $X$ over $K$.
 Denote by $\pi_{K}:X\to X_{K}$ the natural projection.
@@
 It is clear that $Z(x)$ is irreducible.
-We call $Z(x)$ the $\mathbb{C}/K$-closure of $x$ w.r.t the model $X_{K}$.
-We say that $x$ is transcendental if $\dim Z(x)\geq 1$ and call $\dim Z(x)$ the transcendental degree of $x$.
-The notion of transcendental points plays important role in the proof of some cases of the geometric Bombieri-Lang conjecture by Xie-Yuan [30, 31] and the proof of the dynamical André-Oort conjecture for curves by Ji-Xie [11].
+We call $Z(x)$ the $\mathbb{C}/K$-closure of $x$ with respect to the model $X_{K}$.
+We say that $x$ is transcendental if $\dim Z(x)\geq 1$ and call $\dim Z(x)$ the transcendence degree of $x$.
+The notion of transcendental points plays an important role in the proof of some cases of the geometric Bombieri-Lang conjecture by Xie-Yuan [30, 31] and in the proof of the dynamical André-Oort conjecture for curves by Ji-Xie [11].
 Roughly speaking, a very general point in $Z(x)$ satisfies the same algebraic properties as $x$.
-In this paper, we study lengths of periodic points in whose definition (see §1.1) we need the norm map $|\cdot|:\mathbb{C}\to\mathbb{R}_{\geq 0}$ which is not algebraic.
-However $|\cdot|^{2}:\mathbb{C}\to\mathbb{R}$ is real algebraic.
-For this reason, we need to generalize the notion of transcendental points to respect the structure of $X_{K}(K)$ as an algebraic variety over $L$ and the structure of $X(\mathbb{C})$ as a real algebraic variety, and the generalized notion is important in the proof of Theorem 2.1.
-The Weil restriction $R(X)$ of $X$ w.r.t.
-$\mathbb{C}/\mathbb{R}$ is an algebraic variety over $\mathbb{R}$ and the Weil restriction $R(X_{K})$ of $X_{K}$ w.r.t.
-$K/L$ is an algebraic variety over $L$.
+In this paper, we study lengths of periodic points, whose definition (see §1.1) involves the absolute value function $|\cdot|:\mathbb{C}\to\mathbb{R}_{\geq 0}$ which is not algebraic.
+However, the square $|\cdot|^{2}:\mathbb{C}\to\mathbb{R}$ is real algebraic.
+For this reason, we need to generalize the notion of transcendental points to respect the structure of $X_{K}(K)$ as an algebraic variety over $L$ and the structure of $X(\mathbb{C})$ as a real algebraic variety.
+This generalized notion is important in the proof of Theorem 2.1.
+The Weil restriction $R(X)$ of $X$ with respect to $\mathbb{C}/\mathbb{R}$ is an algebraic variety over $\mathbb{R}$, and the Weil restriction $R(X_{K})$ of $X_{K}$ with respect to $K/L$ is an algebraic variety over $L$.
 We have $R(X)=R(X_{K})\otimes_{L}\mathbb{R}$.
 Denote by $\pi_{L}:R(X)\to R(X_{K})$ the natural projection.
-For every $x\in X(\mathbb{C})$, let $Y(x)_{L}$ be the Zariski closure of $\pi_{L}(\psi_{\mathbb{C}}(x))$ in $R(X_{K})$ and $Y(x):=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, where $\psi_{\mathbb{C}}:X(\mathbb{C})\to R(X)(\mathbb{R})$ is the identification in §2.1.
-Set $Z^{\mathbb{R}}(x):=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$ which is a real Zariski closed subset of $X(\mathbb{C})$.
+For every $x\in X(\mathbb{C})$, let $Y(x)_{L}$ be the Zariski closure of $\pi_{L}(\psi_{\mathbb{C}}(x))$ in $R(X_{K})$ and $Y(x):=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, where $\psi_{\mathbb{C}}:X(\mathbb{C})\to R(X)(\mathbb{R})$ is the identification from §2.1.
+Set $Z^{\mathbb{R}}(x):=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$, which is a real Zariski closed subset of $X(\mathbb{C})$.
 We now give a more concrete description of $Z(x)$ and $Z^{\mathbb{R}}(x)$.
@@
 Then we have
-$Z(x)\cap U=\{y\in U|\,\,h(y)=0\text{ for every }h\in\mathcal{O}^{K}(U)\text{ with }h(x)=0\}$
+$Z(x)\cap U=\{y\in U\mid h(y)=0\text{ for every }h\in\mathcal{O}^{K}(U)\text{ with }h(x)=0\}$
 and $Z(x)$ is the Zariski closure of $Z(x)\cap U$.
@@
 Every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ induces a function $h\circ\psi_{\mathbb{C}}$ on $U(\mathbb{C})$.
-The functions of this form are exactly the $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$.
+Functions of this form are exactly the $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$.
 Denote by $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$ the $\mathbb{R}$-algebra of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$.
@@
 Let $\mathcal{O}^{L}(R(U)):=\pi_{L}^{*}(\mathcal{O}(R(U_{K})))$ be the set of algebraic functions defined over $L$ on $R(U)$.
-Let $\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$ the image of $\mathcal{O}^{L}(R(U))\otimes_{L}K$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$, which is the set of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$ defined over $L$.
+Let $\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$ be the image of $\mathcal{O}^{L}(R(U))\otimes_{L}K$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$, which is the set of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$ defined over $L$.
 It is clear that $\mathcal{O}^{K}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$.
@@
 We have
-$Z^{\mathbb{R}}(x)\cap U(\mathbb{C})=\{y\in U(\mathbb{C})|\,\,h(y)=0\text{ for every }h\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)\text{ with }h(x)=0\}$
+$Z^{\mathbb{R}}(x)\cap U(\mathbb{C})=\{y\in U(\mathbb{C})\mid h(y)=0\text{ for every }h\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)\text{ with }h(x)=0\}$
 and $Z^{\mathbb{R}}(x)$ is the real Zariski closure of $Z^{\mathbb{R}}(x)\cap U(\mathbb{C})$. This implies the following lemma.
@@
 Let $x^{\prime}\in X^{\prime}(\mathbb{C})$ and $x:=f(x^{\prime})\in X(\mathbb{C})$.
-Then we have $f(Z^{\mathbb{R}}(x^{\prime}))\subseteq Z^{\mathbb{R}}(x)$.
+Then $f(Z^{\mathbb{R}}(x^{\prime}))\subseteq Z^{\mathbb{R}}(x)$.
 Lemma 2.14.
-We have $Z^{\mathbb{R}}(x)\subseteq Z(x)$ and $Z^{\mathbb{R}}(x)$ is Zariski dense in $Z(x)$. In particular, if $x$ is transcendental, then $\dim_{\mathbb{R}}Z^{\mathbb{R}}(x)\geq 1$.
+We have $Z^{\mathbb{R}}(x)\subseteq Z(x)$ and $Z^{\mathbb{R}}(x)$ is Zariski-dense in $Z(x)$. In particular, if $x$ is transcendental, then $\dim_{\mathbb{R}}Z^{\mathbb{R}}(x)\geq 1$.
 Proof.
@@
 Then we have
-$h\otimes_{\mathbb{C}}1\in\mathcal{O}(X)\otimes_{\mathbb{C}}\tau(\mathcal{O}(X))=\mathcal{C}^{\mathbb{R}-alg}(X)=\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}.$
+$h\otimes_{\mathbb{C}}1\in\mathcal{O}(X)\otimes_{\mathbb{C}}\tau(\mathcal{O}(X))=\mathcal{C}^{\mathbb{R}-{\rm alg}}(X)=\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}.$
 View $\mathbb{C}$ as a vector space over $K$ and let $(e_{j})_{j\in J}$ be a $K$-basis of $\mathbb{C}$.
@@
 Write $h\otimes_{\mathbb{C}}1=\sum_{j\in J}g_{j}e_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}$, where $g_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)$.
-We have $g_{j}(x)=0$ for every $j\in J$.
+Then $g_{j}(x)=0$ for every $j\in J$.
 Let $(f_{n})_{n\in N}$ be a $K$-basis of $\mathcal{O}^{K}(X)$. We may assume that $0\in N$ and $1=f_{0}$. Write
 $g_{j}=\sum_{m,n\in N}b_{j,m,n}f_{m}\otimes_{K}\tau(f_{n}).$
-The we get
+Then we get
 $h\otimes_{\mathbb{C}}1=\sum_{j\in J,m,n\in N}b_{j,m,n}e_{j}f_{m}\otimes_{K}\tau(f_{n}).$
@@
 $\mathcal{C}^{\mathbb{R}-{\rm alg}}(X)=\mathcal{O}^{K}(X)\otimes_{K}\tau(\mathcal{O}^{K}(X))\otimes_{K}\mathbb{C},$
-and $h\otimes_{\mathbb{C}}1$ is contained in the image of $\mathcal{O}(X)=\mathcal{O}^{K}(X)\otimes_{K}\mathbb{C}$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(X).$ We have
+and $h\otimes_{\mathbb{C}}1$ is contained in the image of $\mathcal{O}(X)=\mathcal{O}^{K}(X)\otimes_{K}\mathbb{C}$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(X)$, we have
 $b_{j,m,n}=0$
-for every $n\neq 0$. So $g_{j}=\sum_{m\in N}b_{j,m,0}f_{m}\otimes 1\in\mathcal{O}^{K}(X)$. Since $g_{j}(x)=0$, $g_{j}|_{Z(x)}=0$. Then we have $h|_{Z(x)}=0$ which concludes the proof. ∎
+for every $n\neq 0$. So $g_{j}=\sum_{m\in N}b_{j,m,0}f_{m}\otimes 1\in\mathcal{O}^{K}(X)$. Since $g_{j}(x)=0$, $g_{j}\mid_{Z(x)}=0$. Hence $h\mid_{Z(x)}=0$, which completes the proof. ∎
 Lemma 2.15.
@@
 Write $g=\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i$, where $g_{1},g_{2}\in\mathcal{O}(R(X_{K}))$.
-Since $h(x)\in K$, we see that both $\pi_{L}^{*}(g_{1})(\psi_{\mathbb{C}}(x))$ and $\pi_{L}^{*}(g_{2})(\psi_{\mathbb{C}}(x))$ are in $L$.
-The map $\pi_{L}|_{\psi_{\mathbb{C}}(x)}:\psi_{\mathbb{C}}(x)\to Y(x)_{L}$ induces an embedding
-$(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}:\mathcal{O}(Y(x)_{L})\hookrightarrow\mathbb{R}.$
-We view elements of $L$ as constant functions contained $\mathcal{O}(Y(x)_{L})$ and note that $(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}(L)=L\subseteq\mathbb{R}.$ For $i\in\{1,2\}$, we have $g_{i}|_{Y(x)_{L}}\in\mathcal{O}(Y(x)_{L})$.
-As
-$(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}(g_{i}|_{Y(x)_{L}})=\pi_{L}^{*}(g_{i})(\psi_{\mathbb{C}}(x))\in L,$
-the injectivity of $(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}$ implies that $g_{i}|_{Y(x)_{L}}\in L$, for $i=1,2$.
-So both $\pi_{L}^{*}(g_{1})$ and $\pi_{L}^{*}(g_{2})$ are constants on $Y(x)=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, with values in $L$.
+Since $h(x)\in K$, both $\pi_{L}^{*}(g_{1})(\psi_{\mathbb{C}}(x))$ and $\pi_{L}^{*}(g_{2})(\psi_{\mathbb{C}}(x))$ are in $L$.
+The map $\pi_{L}\mid_{\psi_{\mathbb{C}}(x)}:\psi_{\mathbb{C}}(x)\to Y(x)_{L}$ induces an injection
+$(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}:\mathcal{O}(Y(x)_{L})\hookrightarrow\mathbb{R}.$
+We view elements of $L$ as constant functions in $\mathcal{O}(Y(x)_{L})$ and note that
+$(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}(L)=L\subseteq\mathbb{R}.$
+For $i\in\{1,2\}$, we have $g_{i}\mid_{Y(x)_{L}}\in\mathcal{O}(Y(x)_{L})$. Since
+$(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}(g_{i}\mid_{Y(x)_{L}})=\pi_{L}^{*}(g_{i})(\psi_{\mathbb{C}}(x))\in L,$
+the injectivity of $(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}$ implies that $g_{i}\mid_{Y(x)_{L}}\in L$ for $i=1,2$.
+So both $\pi_{L}^{*}(g_{1})$ and $\pi_{L}^{*}(g_{2})$ are constant on $Y(x)=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, with values in $L$.
 We conclude that $h=g\circ\psi_{\mathbb{C}}=(\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i)\circ\psi_{\mathbb{C}}$ is constant on $Z^{\mathbb{R}}(x)$, since $Z^{\mathbb{R}}(x)=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$.
@@
 For $d\geq 2$, let ${\rm Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$.
-It is a smooth irreducible affine variety of dimension $12d+$ 1 [25].
+It is a smooth irreducible affine variety of dimension $2d+1$ [25].
 The group ${\rm PGL}_{2}={\rm Aut}(\mathbb{P}^{1})$ acts on ${\rm Rat}_{d}$ by conjugacy.
@@
 $\mathcal{M}_{d}:={\rm Rat}_{d}/{\rm PGL}_{2}={\rm Spec}(\mathcal{O}({\rm Rat}_{d})^{{\rm PGL}_{2}})$
-is the (coarse) moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$, which is an irreducible affine variety of dimension $22d-$ 2, see [25, 24] for a detailed treatment. Let
+is the (coarse) moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$, which is an irreducible affine variety of dimension 2d22d-2, see [25, 24] for a detailed treatment. Let
 $\Psi:{\rm Rat}_{d}\to\mathcal{M}_{d}$
@@
 Let ${\rm FL}_{d}\subseteq{\rm Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}$.
-Set $[{\rm FL}_{d}]:=\Psi({\rm FL}_{d})$.
-The flexible-Lattès locus $[{\rm FL}_{d}]$ is non-empty if and only if $d$ is a square, and in this case $[{\rm FL}_{d}]$ is equidimensional of dimension $1$.
-The above construction works over any algebraically closed field of characteristic 00 and commutes with base changes.
-In the proof of Theorem 2.1, we need to consider only repelling periodic points of rational maps and the notion of moduli spaces with certain marked periodic points is necessary.
+Set ${\rm FL}_{d}]:=\Psi({\rm FL}_{d})$.
+The flexible-Lattès locus ${\rm FL}_{d}]$ is non-empty if and only if $d$ is a square, and in this case ${\rm FL}_{d}]$ is equidimensional of dimension $1$.
+The above construction works over any algebraically closed field of characteristic 0 and commutes with base changes.
+In the proof of Theorem 2.1, we need to consider only repelling periodic points of rational maps, and the notion of moduli spaces with certain marked periodic points is necessary.
 We introduce such moduli spaces here.
-See [11, §8] for more information.
+See [11, §8] for more details.
 For every $n\in\mathbb{Z}_{>0}$, we construct the space ${\rm Rat}_{d}[n]$ of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with a single marked $n$-periodic point as follows.
@@
 be the morphism given by the multiplier at the marked periodic point.
-Next we introduce the moduli space of endomorphisms of degree $d$ with a finite sequence of marked periodic points whose periods satisfy certain conditions.
+Next, we introduce the moduli space of endomorphisms of degree $d$ with a finite sequence of marked periodic points whose periods satisfy certain conditions.
 Let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$.
-We construct the space ${\rm R_{d}}(s_{1},\dots,s_{n})$ of rational maps of degree $d$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already $(n-1)!$-periodic, $\dots$, such that among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), and we always count with multiplicity, as follows.
-An $m$-periodic point of $f$ just means a fixed point of $f^{m}$ and we do not require its exact period to be $m$, hence an $m!$-periodic point is automatically a $k!$-periodic point for $k\geq m$.
+We construct the space ${\rm R_{d}}(s_{1},\dots,s_{n})$ of rational maps of degree $d$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already (n1)!(n-1)!-periodic, $\dots$, and among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), always counted with multiplicity, as follows.
+An $m$-periodic point of $f$ means a fixed point of $f^{m}$, and we do not require its exact period to be $m$; hence, an $m!$-periodic point is automatically $k!$-periodic for $k\geq m$.
 Consider the fiber product $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$ of $s_{n}$ copies of ${\rm Rat}_{d}[n!]$ over ${\rm Rat}_{d}$.
@@
 $\pi_{i,j}:({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}\to({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$
-be the projection to the $i,j$ coordinates.
+be the projection to the $i$-th and $j$-th coordinates.
 The diagonal $\Delta\subseteq({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$ is an irreducible component of $({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$.
@@
 in which the $s_{n}$ marked points are pairwise distinct.
-Let $U^{\prime}$ be the subset of $U$ of points $(f,x_{1},\dots,x_{s_{n}})$ satisfying $f^{m!}(x_{i})=x_{i}$ for all integers $1\leq m\leq n$ and $1\leq i\leq s_{m}$.
-It means that in $U^{\prime}$ the first $s_{m}$ points in the whole sequence of $s_{n}$ marked $n!$-periodic points are indeed $m!$-periodic, for all integers $1\leq m\leq n$.
+Let $U^{\prime}$ be the subset of $U$ consisting of points $(f,x_{1},\dots,x_{s_{n}})$ satisfying $f^{m!}(x_{i})=x_{i}$ for all integers $1\leq m\leq n$ and $1\leq i\leq s_{m}$.
+This means that in $U^{\prime}$, the first $s_{m}$ points in the whole sequence of $s_{n}$ marked $n!$-periodic points are indeed $m!$-periodic, for all integers $1\leq m\leq n$.
 The set $U^{\prime}$ is open and closed in $U$.
@@
 $\lambda_{s_{1},\dots,s_{n}}:{\rm R_{d}}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}:(f,x_{1},\dots,x_{s_{n}})\mapsto(df^{n!}(x_{1}),\dots,df^{n!}(x_{s_{n}})).$
-Recall that for a rational map $g(z)\in\mathbb{C}(z)$ of degree $\geq 2$ and a fixed point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ of $g$, the multiplicity of the fixed point $z_{0}$ of $g$ is the order of $g(z)-z=0$ at $z_{0}$ when $z_{0}\neq\infty$, and we can define the multiplicity using a conjugation to move $z_{0}$ to $\mathbb{C}$ when $z_{0}=\infty$.
-For every $k\geq 1$, it is clear that a fixed point $x$ of $f_{t}^{k}$ has multiplicity $1$ iff
+Recall that for a rational map $g(z)\in\mathbb{C}(z)$ of degree $\geq 2$ and a fixed point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ of $g$, the multiplicity of the fixed point $z_{0}$ of $g$ is the order of g(z)z=0g(z)-z=0 at $z_{0}$ when $z_{0}\neq\infty$, and we can define the multiplicity using a conjugation to move $z_{0}$ to $\mathbb{C}$ when $z_{0}=\infty$.
+For every $k\geq 1$, it is clear that a fixed point $x$ of $f_{t}^{k}$ has multiplicity $1$ if and only if
 (2.4) $\lambda_{k}(f_{t},x)=df_{t}^{k}(x)=\rho_{f^{k}}(x)\neq 1.$
@@
 $\mathcal{M}_{d}(s_{1},\dots,s_{n}):={\rm R_{d}}(s_{1},\dots,s_{n})/{\rm PGL}_{2}$
-to be the moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already $(n-1)!$-periodic, $\dots$, such that in these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), and we always count with multiplicity.
-The moduli space $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ is affine, in fact, so the construction is relatively easy.
+to be the moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already (n1)!(n-1)!-periodic, $\dots$, and among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), always counted with multiplicity.
+The moduli space $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ is affine, which makes the construction relatively easy.
 The morphisms $\phi_{n,m}$ and $\lambda_{s_{1},\dots,s_{n}}$ descend to
-$[\phi_{n,m}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d}(s_{1},\dots,s_{m})$
+$\phi_{n,m}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d}(s_{1},\dots,s_{m})$
 when $1\leq m\leq n$,
-$[\phi_{n,0}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d},$
+$\phi_{n,0}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d},$
 and
-$[\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}.$
-Then $[\phi_{n,0}]$ is étale at every point $x\in[\lambda_{s_{1},\dots,s_{n}}]^{-1}((\mathbb{A}^{1}\setminus\{1\})^{s_{n}})$.
-All the schemes ${\rm Rat}_{d}$, $\mathcal{M}_{d}$, ${\rm R_{d}}(s_{1},\dots,s_{n})$, and $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ which we consider are defined over $\mathbb{Q}$.
+$\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}.$
+Then $\phi_{n,0}]$ is étale at every point $x\in[\lambda_{s_{1},\dots,s_{n}}]^{-1}((\mathbb{A}^{1}\setminus\{1\})^{s_{n}})$.
+All the schemes ${\rm Rat}_{d}$, $\mathcal{M}_{d}$, ${\rm R_{d}}(s_{1},\dots,s_{n})$, and $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ considered here are defined over $\mathbb{Q}$.
 2.5. Length maps
-Recall that $K$ is an algebraically closed subfield of $\mathbb{C}$ which is invariant under the complex conjugation.
-We use the terminology and notations in §2.1-2.4.
-For $d\geq 2$, let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$.
-Let
+Recall that $K$ is an algebraically closed subfield of $\mathbb{C}$ which is invariant under complex conjugation. We use the terminology and notations from §2.12.4.
+For $d\geq 2$, let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. Let
 $|\lambda_{s_{1},\dots,s_{n}}|:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{R}_{\geq 0}^{s_{n}}$
 be the composition of
-$[\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{C}^{s_{n}}$
-and the norm map
+$\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{C}^{s_{n}}$
+and the absolute value map
 $(a_{1},\dots,a_{s_{n}})\in\mathbb{C}^{s_{n}}\mapsto(|a_{1}|,\dots,|a_{s_{n}}|)\in\mathbb{R}_{\geq 0}^{s_{n}}.$
@@
 In this subsection, we recall the rigidity of length spectrum proved by Ji-Xie [12].
-Let $f$ be an endomorphism of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$. As in [12, §8.3] the length spectrum
+Let $f$ be an endomorphism of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$. As in [12, §8.3], the length spectrum
 $L(f)=\{L(f)_{n},n\geq 1\}$
 of $f$ is a sequence of finite multisets [^1] A multiset is a set except allowing multiple instances for each of its elements.
-The number of the instances of an element is called the multiplicity.
-For example: $\{a,a,b,c,c,c\}$ is a multiset of cardinality $6$, the multiplicities for $a,b,c$ are 2,1,3, respectively., where $L(f)_{n}:=L_{n}(f)$ is the multiset of norms of multipliers (i.e., lengths) of all fixed points of $f^{n}$.
+The number of instances of an element is called its multiplicity.
+For example: $\{a,a,b,c,c,c\}$ is a multiset of cardinality $6$, with multiplicities $2,1,3$ for $a,b,c$, respectively., where $L(f)_{n}:=L_{n}(f)$ is the multiset of lengths of all fixed points of $f^{n}$.
 In particular, $L(f)_{n}$ is a multiset of non-negative real numbers of cardinality $d^{n}+1$.
-For every $n\geq 1$, let $RL(f)_{n}$ be the sub-multiset of $L(f)_{n}$ consisting of all elements $>1$.
+For every $n\geq 1$, let $RL(f)_{n}$ be the sub-multiset of $L(f)_{n}$ consisting of all elements >1>1.
 We call
@@
 $(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$
-is increasing, since every $n!$-periodic point is $(n+1)!$-periodic and the multiplicity of a repelling fixed point of $f^{n!}$ is $1$ by (2.4).
-Also, the sequence $(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$ is bounded from above, since every rational map on $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ has only finitely many periodic cycles whose multiplier has norm $\leq 1$ (see [19, Theorem 13.1]).
-As $L(f),RL(f)$, and $RL^{*}(f)$ are invariant under conjugacy, they descend on $\mathcal{M}_{d}(\mathbb{C})$.
-For every $[f]\in\mathcal{M}_{d}(\mathbb{C})$, define
+is non-decreasing, since every $n!$-periodic point is $(n+1)!$-periodic and the multiplicity of a repelling fixed point of $f^{n!}$ is $1$ by (2.4).
+Also, the sequence $(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$ is bounded from above, since every rational map on $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ has only finitely many periodic cycles whose multiplier has absolute value $\leq 1$ (see [18, Theorem 13.1]).
+Since $L(f),RL(f)$, and $RL^{*}(f)$ are invariant under conjugacy, they descend to $\mathcal{M}_{d}(\mathbb{C})$.
+For every $f]\in\mathcal{M}_{d}(\mathbb{C})$, define
 $L([f]):=L(f),RL([f]):=RL(f),RL^{*}([f]):=RL^{*}(f)$
 for any $f$ in the class $[f]$.
-Let $\Omega$ be the set of sequences $\{A_{n},n\geq 1\}$ of multisets consisting of real numbers of norm strictly larger than $1$ satisfying #$\#A_{n}\leq d^{n!}+1$ for $n\geq 1$, and such that for every $a\in A_{n}$ with multiplicity $m\geq 1$, we have $a^{n+1}\in A_{n+1}$ with multiplicity $\geq m$.
+Let $\Omega$ be the set of sequences $\{A_{n},n\geq 1\}$ of multisets consisting of real numbers of absolute values >1>1 satisfying $A_{n}\leq d^{n!}+1$ for $n\geq 1$, and such that for every $a\in A_{n}$ with multiplicity $m\geq 1$, we have $a^{n+1}\in A_{n+1}$ with multiplicity $\geq m$.
 For $A,B\in\Omega$, we write $A\subseteq B$ if $A_{n}\subseteq B_{n}$ as multisets for every $n\geq 1$.
@@
 For every endomorphism $f$ of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d$, we have $RL^{*}(f)\in\Omega$ and it is big.
-For a big $A\in\Omega$, Ji-Xie proved the following rigidity result on length spectrum, which is a main ingredient of the proof of Theorem 2.1.
+For a big $A\in\Omega$, Ji-Xie proved the following rigidity result on length spectrum, which is a main ingredient in the proof of Theorem 2.1.
 Theorem 2.16.
@@
 Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$.
-Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $|\rho_{f}(x)|\in K$.
-We want to show that $[f]\in\mathcal{M}_{d}(\mathbb{C})$ is not transcendental over $K$ for the model $(\mathcal{M}_{d})_{K}$.
-Now assume that $[f]$ is transcendental.
+Assume that $f$ is not a flexible Lattès map and that for every $x\in{\rm Per}(f)(\mathbb{C})$, $|\rho_{f}(x)|\in K$.
+We want to show that $f]\in\mathcal{M}_{d}(\mathbb{C})$ is not transcendental over $K$ for the model $(\mathcal{M}_{d})_{K}$.
+Now assume, for contradiction, that $[f]$ is transcendental.
 Set $A:=RL^{*}(f)\in\Omega$, which is big.
@@
 We may pick a sequence of periodic points $(x_{i})_{i=1}^{\infty}$ such that for every $n\geq 1$, the first $s_{n}$ points $x_{1},\dots,x_{s_{n}}$ are fixed by $f^{n!}$, and we have
-$A_{n}=\{|\rho(x_{i})|^{n!}:i=1,\dots,s_{n}\}.$
-Let $[f_{n}]\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$ be the point presented by $(f,x_{1},\dots,x_{s_{n}})$. It is clear that $[\phi_{n,0}]([f_{n}])=[f]$ for every $n\geq 1$.
-Since $[f]$ is transcendental, for every $n\geq 1$, $[f_{n}]$ is transcendental as well.
+$A_{n}=\left\{\left|\rho_{f^{n!}}(x_{i})\right|:i=1,\dots,s_{n}\right\}.$
+Let $f_{n}]\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$ be the point represented by $(f,x_{1},\dots,x_{s_{n}})$. It is clear that $\phi_{n,0}]([f_{n}])=[f]$ for every $n\geq 1$.
+Since $[f]$ is transcendental, for every $n\geq 1$, $f_{n}]$ is also transcendental.
 By Lemma 2.14, we get $\dim_{\mathbb{R}}(Z^{\mathbb{R}}([f_{n}]))\geq 1$ for every $n\geq 1$.
@@
 Then from the last paragraph of §2.5, we see that $|\lambda_{s_{1},\dots,s_{n}}|$ is constant on $Z^{\mathbb{R}}(f_{n})$.
-As $|\lambda_{s_{1},\dots,s_{n}}|([f_{n}])\in(1,+\infty)^{s_{n}}$, the morphism $[\phi_{n,0}]$ is étale in a neighborhood of $Z^{\mathbb{R}}(f_{n})$.
-Since $[\phi_{n,0}]$ is a finite morphism, the subset
+As $|\lambda_{s_{1},\dots,s_{n}}|([f_{n}])\in(1,+\infty)^{s_{n}}$, the morphism $\phi_{n,0}]$ is étale in a neighborhood of $Z^{\mathbb{R}}([f_{n}])$.
+Since $\phi_{n,0}]$ is a finite morphism, the subset
 $V_{n}:=[\phi_{n,0}](Z^{\mathbb{R}}([f_{n}]))$
@@
 Then for every $g\in V_{N}$, we have $A\subseteq RL^{*}([g])$.
-Since $[f]\not\in[{\rm FL}_{d}]$, the set $Z^{\mathbb{R}}([f_{N}])$ is real irreducible with $\dim_{\mathbb{R}}Z^{\mathbb{R}}([f_{N}])\geq 1$, hence $V_{N}\cap(\mathcal{M}_{d}(\mathbb{C})\setminus[{\rm FL}_{d}])$ is infinite.
-This contradicts to Theorem 2.16, which concludes the proof.
+Since $f]\not\in[{\rm FL}_{d}]$, the set $Z^{\mathbb{R}}([f_{N}])$ is real irreducible with $\dim_{\mathbb{R}}Z^{\mathbb{R}}([f_{N}])\geq 1$, so $V_{N}\cap(\mathcal{M}_{d}(\mathbb{C})\setminus[{\rm FL}_{d}])$ is infinite.
+This contradicts Theorem 2.16, which completes the proof.
 ∎
 3. An equidistribution theorem
-Results of equidistribution-type play an important role in the study of arithmetic geometry and arithmetic dynamics in past decades.
+Equidistribution-type results have played a significant role in arithmetic geometry and arithmetic dynamics over the past decades.
 In this section, we state and prove a reformulation (Theorem 3.1) of Yuan’s equidistribution theorem [32, Theorem 3.1].
-The reason that we present the theorem is that the equidistribution theorems in the literature (for example, [32]) are only dealing with the Galois orbit of a single small point at one time, which is not sufficient for our proof of Theorem 1.5.
-Our statement is slightly stronger than [32, Theorem 3.1] as we work with a sequence of (weighted) countable Galois invariant subsets, and this is useful for applications in dynamics.
-The presented proof of Theorem 3.1 is by applying Yuan’s original equidistribution theorem [32, Theorem 3.1].
+We present this theorem because existing equidistribution theorems in the literature (e.g., [32]) typically handle the Galois orbit of a single small point at a time, which is insufficient for our proof of Theorem 1.5.
+Our statement is slightly stronger than [32, Theorem 3.1] as we work with sequences of (weighted) countable Galois-invariant subsets, which is particularly useful for applications in dynamics.
+The proof of Theorem 3.1 is based on applying Yuan’s original equidistribution theorem [32, Theorem 3.1].
 We follow the terminology in [32].
-For the convenience of readers, we present a short review of some basic notions in Arakelov geometry. See [36, 32, 20, 33] for more details.
-Let $K$ be a number field, $X$ a projective variety of dimension $n-1$ over $K$, and $\mathcal{L}$ a line bundle on $X$.
+For the convenience of readers, we briefly review some basic notions in Arakelov geometry. More details can be found in [36, 32, 20, 33].
+Let $K$ be a number field, $X$ a projective variety of dimension n1n-1 over $K$, and $\mathcal{L}$ a line bundle on $X$.
 We fix an embedding $\overline{K}\to\mathbb{C}_{v}$ for each place $v\in\mathcal{M}_{K}$ of $K$.
-An adelic metric on $\mathcal{L}$ is a $\mathbb{C}_{v}$-norm $\|\cdot\|_{v}$ on the fiber $\mathcal{L}_{\mathbb{C}_{v}}(x)$ for every $x\in X(\overline{K})$ and $v\in\mathcal{M}_{K}$, satisfying certain continuity and coherence conditions, see [36, §1] for a precise definition.
-An adelic metric on $\mathcal{L}$ is called an algebraic metric (or a model metric) if it is induced by an $\mathcal{O}_{K}$-model $(\mathcal{X},\tilde{\mathcal{L}})$ of $(X,\mathcal{L}^{e})$ for some $e\in\mathbb{Z}_{>0}$ (i.e., $\mathcal{X}$ is an integral scheme projective and flat over $\mathcal{O}_{K}$ and $\tilde{\mathcal{L}}$ is a Hermitian line bundle over $\mathcal{X}$ such that the generic fiber of $(\mathcal{X},\tilde{\mathcal{L}})$ is $(X,\mathcal{L}^{e})$).
-Here the induced metric at an archimedean place is induced by the Hermitian metric of $\tilde{\mathcal{L}}$, and for a non-archimedean place $v$ and $x\in X(\overline{K})$, the induced metric on $\mathcal{L}_{\mathbb{C}_{v}}(x)$ is determined by the lattice $(\tilde{x}^{*}\tilde{\mathcal{L}}_{\mathcal{O}_{\mathbb{C}_{v}}})^{1/e}$ in $\mathcal{L}_{\mathbb{C}_{v}}(x)$, where $\tilde{x}:{\rm Spec}(\mathcal{O}_{\mathbb{C}_{v}})\to\mathcal{X}_{\mathcal{O}_{\mathbb{C}_{v}}}$ is the extension of $x$.
-The metric induced by such a model always satisfies the conditions of an adelic metric, see [36, §1].
-The algebraic metric is called semipositive if $\tilde{\mathcal{L}}$ has semipositive curvature at all archimedean places and non-negative degree on every vertical curve of $\mathcal{X}$.
-An (adelic) metric on $\mathcal{L}$ is called semipositive if it is the uniform limit of some sequence of semipositive algebraic metrics on $\mathcal{L}$.
-Assume that $\mathcal{L}$ is ample and equip it with a semipositive (adelic) metric (such a metric always exists).
+An adelic metric on $\mathcal{L}$ is a $\mathbb{C}_{v}$-norm $\|\cdot\|_{v}$ on the fiber $\mathcal{L}_{\mathbb{C}_{v}}(x)$ for every $x\in X(\overline{K})$ and $v\in\mathcal{M}_{K}$, satisfying certain continuity and coherence conditions; see [36, §1] for a precise definition.
+[…]
+An (adelic) metric on $\mathcal{L}$ is called semipositive if it is the uniform limit of a sequence of semipositive algebraic metrics on $\mathcal{L}$.
+Assume that $\mathcal{L}$ is ample and equipped with a semipositive (adelic) metric (which always exists).
 The resulting metrized line bundle is denoted by $\overline{\mathcal{L}}$.
@@
 $h_{\overline{\mathcal{L}}}(Y)=\frac{\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}}{(\dim(Y)+1)\deg_{\mathcal{L}}(\overline{Y})},$
-where $\overline{Y}$ is the image of $Y\to X_{\overline{K}}\to X$ (i.e., the closed subvariety of $X$ corresponding to the Galois orbit of $Y$), and $\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}$ is the arithmetic self-intersection number of ¯$\overline{\mathcal{L}}|_{\overline{Y}}$ on $\overline{Y}$.
-In particular, for every algebraic point $x\in X(\overline{K})$, we get its height $h_{\overline{\mathcal{L}}}(x)$ and the function $h_{\overline{\mathcal{L}}}:{X(\overline{K})}\to\mathbb{R}$ is a Weil height associated to $\mathcal{L}$ (up to a normalizing factor $[K:\mathbb{Q}]$).
+where $\overline{Y}$ is the image of $Y\to X_{\overline{K}}\to X$ (i.e., the closed subvariety of $X$ corresponding to the Galois orbit of $Y$), and $\hat{c}_{1}(\overline{\mathcal{L}}|_{\overline{Y}})^{\dim(Y)+1}$ is the arithmetic self-intersection number of $\overline{\mathcal{L}}|_{\overline{Y}}$ on $\overline{Y}$.
+In particular, for every algebraic point $x\in X(\overline{K})$, we get its height $h_{\overline{\mathcal{L}}}(x)$, and the function $h_{\overline{\mathcal{L}}}:{X(\overline{K})}\to\mathbb{R}$ is a Weil height associated to $\mathcal{L}$ (up to a normalizing factor $K:\mathbb{Q}]$).
 For an algebraic point $x\in X(\overline{K})$, we have
@@
 For $n\in\mathbb{Z}_{>0}$, let $S_{n}$ be a countable subset of $X(\overline{\mathbb{Q}})$ which is ${\rm Gal}(\overline{\mathbb{Q}}/K)$-invariant.
-For every $n\geq 1$ and $y\in S_{n}$, fix a real number $a_{n,y}>0$ such that $\sum_{y\in S_{n}}a_{n,y}=1$ and $a_{n,y}=a_{n,\sigma(y)}$, for all $n\geq 1$, $y\in S_{n}$, and $\sigma\in{\rm Gal}(\overline{\mathbb{Q}}/K)$.
+For every $n\geq 1$ and $y\in S_{n}$, fix a real number $a_{n,y}>0$ such that $\sum_{y\in S_{n}}a_{n,y}=1$ and $a_{n,y}=a_{n,\sigma(y)}$ for all $n\geq 1$, $y\in S_{n}$, and $\sigma\in{\rm Gal}(\overline{\mathbb{Q}}/K)$.
 Assume that the following two conditions hold:
 (1) (small) $\sum_{y\in S_{n}}a_{n,y}h_{\overline{\mathcal{L}}}(y)\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$;
-(2) (generic) for every proper subvariety $V\subsetneqq X$ of $X$, $\sum_{y\in S_{n}\cap V}a_{n,y}\to 0$ as $n\to+\infty$.
-Then the measure $\mu_{n}:=\sum_{y\in S_{n}}a_{n,y}\delta_{y}$ converges weakly to $\mu$ on $X(\mathbb{C})$ as $n\to+\infty$, where $\delta_{y}$ denotes the Dirac measure at the point $y$, i.e., for all continuous function $g$ on $X(\mathbb{C})$ (with the complex topology), we have
-[…]
+(2) (generic) for every proper subvariety $V\subsetneqq X$, $\sum_{y\in S_{n}\cap V}a_{n,y}\to 0$ as $n\to+\infty$.
+Then the measure $\mu_{n}:=\sum_{y\in S_{n}}a_{n,y}\delta_{y}$ converges weakly to $\mu$ on $X(\mathbb{C})$ as $n\to+\infty$, where $\delta_{y}$ denotes the Dirac measure at the point $y$; i.e., for every continuous function $g$ on $X(\mathbb{C})$ (with the complex topology), we have
+(3.5) $\lim\limits_{n\to\infty}\sum_{y\in S_{n}}a_{n,y}g(y)=\int_{X(\mathbb{C})}gd\mu.$
 Proof.
 For every $n\geq 1$, write $S_{n}=\sqcup_{i=1}^{m_{n}}O_{n,i}$, where $O_{n,i}$ are the distinct Galois orbits contained in $S_{n}$ and $m_{n}\in\mathbb{Z}_{>0}\cup\{\infty\}$.
-In [32, Theorem 3.1], Yuan prove the theorem for the case where every $S_{n}$ ($n\geq 1$) is consisting of a single Galois orbit, i.e., $m_{n}=1$ for every $n\geq 1$.
-[…]
-[…]
+In [32, Theorem 3.1], Yuan proved the theorem for the case where each $S_{n}$ ($n\geq 1$) consists of a single Galois orbit, i.e., $m_{n}=1$ for every $n\geq 1$.
+By contradiction, after passing to a subsequence of $(S_{n})$, we may assume that there is a continuous (real-valued) function $g$ on $X(\mathbb{C})$ and a constant $c\in\mathbb{R}_{>0}$ such that $\int_{X(\mathbb{C})}gd\mu=0$ and
+(3.6) $\int_{X(\mathbb{C})}gd\mu_{n}\geq c>0$
 for every $n\geq 1$.
-As $a_{n,y}$ is constant for $y$ in its Galois orbit, we can define $a_{n,i}:=a_{n,y}$ where $y$ is an arbitrary point in $O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
+As $a_{n,y}$ is constant on Galois orbits, we can define $a_{n,i}:=a_{n,y}$ for an arbitrary $y\in O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
 For all $n\geq 1$ and $1\leq i\leq m_{n}$, set $\mu_{n,i}:=\sum_{y\in O_{n,i}}\delta_{y}$.
 Then we have
-$\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})=1\text{ and }\mu_{n}=\sum_{i=1}^{m_{n}}a_{n,i}\mu_{n,i}$
+$\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})=1\quad\text{and}\quad\mu_{n}=\sum_{i=1}^{m_{n}}a_{n,i}\mu_{n,i}$
 for every $n\geq 1$.
-As $h_{\overline{\mathcal{L}}}(y)$ is constant for $y$ in its Galois orbit, we can define $h_{n,i}:=h_{\overline{\mathcal{L}}}(y)$ where $y$ is an arbitrary point in $O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
-The small condition $(1)$ asserts that $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$. Taking a suitable subsequence, we may assume that
+As $h_{\overline{\mathcal{L}}}(y)$ is constant on Galois orbits, we can define $h_{n,i}:=h_{\overline{\mathcal{L}}}(y)$ for an arbitrary $y\in O_{n,i}$, for all $n\geq 1$ and $1\leq i\leq m_{n}$.
+The small condition (1) implies that $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\to h_{\overline{\mathcal{L}}}(X)$ as $n\to+\infty$. After taking a suitable subsequence, we may assume that
 (3.7) $\sum_{i=1}^{m_{n}}a_{n,i}(\#O_{n,i})h_{n,i}\leq h_{\overline{\mathcal{L}}}(X)+\frac{1}{n^{2}}$
 for every $n\geq 1$.
-As $X$ has only countably many subvarieties defined over $\overline{\mathbb{Q}}$, there is a sequence $V_{1}\subseteq V_{2}\subseteq\cdots\subseteq V_{n}\subseteq\cdots\subsetneqq X$ of increasing proper subvarieties of $X$ defined over $K$, such that for every proper subvariety $V$ of $X$ defined over $\overline{\mathbb{Q}}$, there is $N=N(V)\in\mathbb{Z}_{>0}$ such that $V\subseteq V_{n}$ for all integers $n\geq N$.
-(By considering the Galois orbits of subvarieties, we can assume that the $V_{n}$’s are defined over $K$.) By Zhang’s theorem of successive minimum [36, Theorem 1.10], we may further assume that
+As $X$ has only countably many subvarieties defined over $\overline{\mathbb{Q}}$, there is an increasing sequence $V_{1}\subseteq V_{2}\subseteq\cdots\subseteq V_{n}\subseteq\cdots\subsetneqq X$ of proper subvarieties of $X$ defined over $K$ such that for every proper subvariety $V$ of $X$ defined over $\overline{\mathbb{Q}}$, there exists $N=N(V)\in\mathbb{Z}_{>0}$ with $V\subseteq V_{n}$ for all integers $n\geq N$.
+(By considering the Galois orbits of subvarieties, we may assume that the $V_{n}$ are defined over $K$.) By Zhang’s theorem of successive minima [36, Theorem 1.10], we may further assume that
 (3.8) $h_{\overline{\mathcal{L}}}(y)\geq h_{\overline{\mathcal{L}}}(X)-\frac{1}{n^{2}}$
@@
 for $n\geq 1$. (Note that for every subvariety $V$ of $X$ defined over $K$, either $O_{n,i}\cap V=\emptyset$ or $O_{n,i}\subseteq V$.)
-The generic condition $(2)$ shows that after replacing $(S_{n})$ by a suitable subsequence, we may assume that
+The generic condition (2) implies that after passing to a subsequence of $(S_{n})$, we may assume that
 $1-\frac{1}{n^{2}}\leq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})\leq 1,$
-for every $n\geq 1$. Equivalently, we have
+for every $n\geq 1$. Equivalently,
 (3.9) $0\leq\sum_{i\in B_{n}}a_{n,i}(\#O_{n,i})\leq\frac{1}{n^{2}},$
@@
 Set $H_{n}:=\{i\in G_{n}:h_{n,i}\leq h_{\overline{\mathcal{L}}}(X)+1/n\}$ and $I_{n}:=G_{n}\setminus H_{n}$ for $n\geq 1$. By (3.8), we have
-$\frac{1+M}{n^{2}}$ $\geq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})(h_{n,i}-h_{\overline{\mathcal{L}}}(X))\geq\frac{-\sum_{i\in H_{n}}a_{n,i}(\#O_{n,i})}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}$
+$\frac{1+M}{n^{2}}$ $\geq\sum_{i\in G_{n}}a_{n,i}(\#O_{n,i})(h_{n,i}-h_{\overline{\mathcal{L}}}(X))\geq-\frac{\sum_{i\in H_{n}}a_{n,i}(\#O_{n,i})}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}$
 $\geq-\frac{1}{n^{2}}+\frac{\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})}{n}.$
-Therefore, we get
+Therefore,
 $\sum_{i\in I_{n}}a_{n,i}(\#O_{n,i})\leq\frac{2+M}{n}$
@@
 as $n\to+\infty$.
-[…]
+Set $\mu_{n}^{\prime}:=\sum_{i\in H_{n}}a_{n,i}\mu_{n,i}$ for $n\geq 1$.
+Then $\mu_{n}-\mu_{n}^{\prime}$ is a (positive) measure on $X(\mathbb{C})$ whose total mass tends to 0 as $n\to+\infty$.
+After removing finitely many terms, we may assume that $\int_{X(\mathbb{C})}gd\mu_{n}^{\prime}\geq c/2$ for every $n\geq 1$, since $g$ is bounded on the compact set $X(\mathbb{C})$.
+Hence there is $i_{n}\in H_{n}$ such that $\int_{X(\mathbb{C})}gd\mu_{n,i_{n}}\geq c/2$; fix such an $i_{n}$ for each $n\geq 1$.
+Since $O_{n,i_{n}}\cap V_{n}=\emptyset$, the sequence $(O_{n,i_{n}})_{n}$ of Galois orbits satisfies the generic condition (2) by construction of $(V_{n})$.
+By (3.8) and the definition of $H_{n}$, we have
 $h_{\overline{\mathcal{L}}}(X)-\frac{1}{n^{2}}\leq h_{n,i_{n}}\leq h_{\overline{\mathcal{L}}}(X)+\frac{1}{n}$
 for every $n\geq 1$, i.e., the small condition (1) holds for $(O_{n,i_{n}})_{n}$.
-Applying Yuan’s equidistribution Theorem [32, Theorem 3.1] to $(O_{n,i_{n}})_{n}$, we have $\mu_{n,i_{n}}\to\mu$ as $n\to+\infty$.
-Thus we have
-[…]
+Applying Yuan’s equidistribution theorem [32, Theorem 3.1] to $(O_{n,i_{n}})_{n}$, we get $\mu_{n,i_{n}}\to\mu$ as $n\to+\infty$.
+Thus,
+$0<\frac{c}{2}\leq\lim_{n\to\infty}\int_{X(\mathbb{C})}gd\mu_{n,i_{n}}=\int_{X(\mathbb{C})}gd\mu=0,$
 which is a contradiction. ∎
 Remark 3.2.
-For a place $v\in\mathcal{M}_{K}$, we can consider two canonical analytic $v$-spaces of $X$, i.e., the $\mathbb{C}_{v}$-analytic space $X^{an}_{\mathbb{C}_{v}}$ and the $K_{v}$-analytic space $X^{an}_{K_{v}}$.
-We call the former the geometric case and the latter the algebraic case, see [32, §3.1].
+For a place $v\in\mathcal{M}_{K}$, we can consider two canonical analytic $v$-spaces of $X$: the $\mathbb{C}_{v}$-analytic space $X^{an}_{\mathbb{C}_{v}}$ and the $K_{v}$-analytic space $X^{an}_{K_{v}}$.
+We refer to the former as the geometric case and the latter as the algebraic case; see [32, §3.1].
 Theorem 3.1 deals with the geometric case at an archimedean place.
-The same idea also applies for a non-archimedean place or the algebraic case, which gives the full analogy of [32, Theorems 3.1 and 3.2].
-In order to check the “generic” condition in Theorem 3.1, we need the following lemma.
-The proof uses ergodic theory with respect to the constructible topology (on algebraic varieties) introduced by Xie [28].
+The same idea applies to non-archimedean places or the algebraic case, which gives the full analogy of [32, Theorems 3.1 and 3.2].
+To verify the “generic” condition in Theorem 3.1, we need the following lemma. The proof uses ergodic theory with respect to the constructible topology (on algebraic varieties) introduced by Xie [28].
 Lemma 3.3.
 Let $K$ be a number field and $X$ be a projective variety over $K$.
-Given a dominant endomorphism $f:X\to X$, and a sequence $(x_{n})_{n=1}^{\infty}$ of $f$-periodic points in $X(\overline{K})$ which are pairwise distinct.
+Given a dominant endomorphism $f:X\to X$ and a sequence $(x_{n})_{n=1}^{\infty}$ of pairwise distinct $f$-periodic points in $X(\overline{K})$.
 Assume that $(x_{n})_{n=1}^{\infty}$ is generic in $X$, i.e., every proper Zariski closed subset $Z\subsetneqq X$ contains only finitely many $x_{n}$.
 Then for every proper subvariety $V\subsetneqq X$, we have
-(3.11) #$\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}\to 0,\ \text{as}\ n\to+\infty,$
+(3.11) $\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}\to 0,\text{ as }n\to+\infty,$
 where $O_{f}(x_{n})$ is the (forward) orbit of $x_{n}$ under $f$.
@@
 Clearly, it suffices to show that for every subsequence $(n_{k})_{k}$ of $(n)_{n=1}^{\infty}$, there exists a subsubsequence $(n_{k_{l}})_{l}$ such that
-#$\frac{\#(V\cap O_{f}(x_{n_{k_{l}}}))}{\#O_{f}(x_{n_{k_{l}}})}\to 0,\ \text{as}\ l\to+\infty.$
-Fix a proper subvariety $V$ in $X$.
-Let $\lvert X\rvert$ be $X$ equipped with the constructible topology (i.e., the topology of $X$ generated by all its Zariski closed and open subsets) and $\mathcal{M}^{1}(\lvert X\rvert)$ be the space of all probability Radon measures on $\lvert X\rvert$ with the topology of weak convergence relative to all continuous functions on $\lvert X\rvert$.
-Then $\mathcal{M}^{1}(\lvert X\rvert)$ is sequentially compact (cf.
-[28, Corollary 1.14]).
-For $n\in\mathbb{Z}_{>0}$, set
+$\frac{\#(V\cap O_{f}(x_{n_{k_{l}}}))}{\#O_{f}(x_{n_{k_{l}}})}\to 0,\text{ as }l\to+\infty.$
+[…]
 $m_{n}=(\#O_{f}(x_{n}))^{-1}\sum_{z\in O_{f}(x_{n})}\delta_{z}.$
-Denote the characteristic function of $V\subsetneqq\lvert X\rvert$ by $1_{V}$, then $1_{V}$ is continuous with respect to the constructible topology.
-By the sequential compactness of $\mathcal{M}^{1}(\lvert X\rvert)$, it suffices to show that for every subsequence $(n_{k})_{k}$ of $(n)_{n=1}^{\infty}$ with $m_{n_{k}}\to m\ \text{as}\ k\to+\infty$ in $\mathcal{M}^{1}(\lvert X\rvert)$ for some $m\in\mathcal{M}^{1}(\lvert X\rvert)$, we have
 […]
-Without loss of generality, we may assume that $(m_{n})$ itself converges to a measure $m\in\mathcal{M}^{1}(\lvert X\rvert)$; and it suffices to show (3.12) in this case.
-As $f_{*}m_{n}=m_{n}$ for all $n\geq 1$, we see that $f_{*}m=m$ as well.
-Then according to [28, Lemma 5.3], $m$ must have the form $m=\sum_{y\in S}a_{y}\delta_{O_{f}(y)}$, where $S$ is a countable set of $f$-periodic points in $\lvert X\rvert$ (note that the points in $S$ may not be closed with respect to the Zariski topology of $X$), $a_{y}\in\mathbb{R}_{>0}$ with $\sum_{y\in S}a_{y}=1$, and $\delta_{O_{f}(y)}=(\#O_{f}(y))^{-1}\sum_{z\in O_{f}(y)}\delta_{z}$ for $y\in S$.
-As $m_{n}\to m$, we get
+(3.12) $\int 1_{V}dm=0.$
 […]
+$\frac{\#(V\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}=\int 1_{V}dm_{n}\to\int 1_{V}dm,\text{ as }n\to+\infty.$
 Suppose that (3.12) fails.
-Then there must be a scheme point $y\in S$ with $a_{y}>0$ and $V\cap O_{f}(y)\neq\emptyset$.
-Denote the exact period of $y$ under $f$ by $k$.
-Let $Y$ be the Zariski closure of $\{y\}$.
-Then $Y\subseteq\cup_{j=0}^{k-1}f^{j}(V)$, hence $Y$ is also a proper Zariski closed subset of $X$.
+Then there must be a point $y\in S$ (in the scheme-theoretic sense) with $a_{y}>0$ and $V\cap O_{f}(y)\neq\emptyset$.
+Let $k$ be the exact period of $y$ under $f$, and let $Y$ be the Zariski closure of $\{y\}$.
+Then $Y\subseteq\cup_{j=0}^{k-1}f^{j}(V)$, so $Y$ is a proper Zariski closed subset of $X$.
 Note that
-[…]
-Hence for every sufficiently large integer $n\gg 1$, we have
+$\frac{\#(Y\cap O_{f}(x_{n}))}{\#O_{f}(x_{n})}=\int 1_{Y}dm_{n}\to\int 1_{Y}dm\geq\frac{a_{y}}{k}>0,\text{ as }n\to+\infty.$
+Hence for all sufficiently large $n\gg 1$, we have
 (3.13) $x_{n}\in\cup_{j=0}^{\infty}f^{j}(Y)=\cup_{j=0}^{k-1}f^{j}(Y),$
-where the last equality follows from $f^{k}(y)=y$ and $Y=\overline{\{y\}}^{{\rm Zar}}$. Note that
+where the last equality follows from $f^{k}(y)=y$ and $Y=\overline{\{y\}}^{{\rm Zar}}$. Since
 $\dim(\cup_{j=0}^{k-1}f^{j}(Y))=\dim(Y)<\dim(X),$
-hence $\cup_{j=0}^{k-1}f^{j}(Y)$ contains only finitely many $x_{n}$ because $(x_{n})$ is generic, contradicting to (3.13). ∎
+the set $\cup_{j=0}^{k-1}f^{j}(Y)$ contains only finitely many $x_{n}$ because $(x_{n})$ is generic, contradicting (3.13). ∎
 4. Proofs of Theorem 1.5 and the defined over $\overline{\mathbb{Q}}$ case of Theorem 1.4
 Proof of Theorem 1.5.
-Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $d\geq 2$ and there is a number field $K$ such that for every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n}\in K$ for some integer $n=n(z_{0})\geq 1$.
-We need to show that such a $f$ must be exceptional (Definition 1.1).
+Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $d\geq 2$, and there is a number field $K$ such that for every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert^{n}\in K$ for some integer $n=n(z_{0})\geq 1$.
+We need to show that such an $f$ must be exceptional (Definition 1.1).
 Assume by contradiction that $f$ is not exceptional.
-In particularly, $f$ is not a flexible Lattès map.
-For every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert\in\overline{\mathbb{Q}}$, thus $f$ is defined over $\overline{\mathbb{Q}}$ (hence over some number field) by Theorem 1.2.
-After replacing $K$ by a suitable finite extension of $K$, we may assume that both $f$ and $\overline{f}$ are defined over $K$.
-Here we denote by $\overline{f}$ the rational map obtained from $f$ via replacing the coefficients by their complex conjugations.
-According to [10, Theorem 9 and Lemma 11], which are essentially proved by Zdunik [35], there exists a sequence $(x_{n})_{n=1}^{\infty}$ of distinct points in ${\rm Per}^{*}(f)$ such that
+In particular, $f$ is not a flexible Lattès map.
+For every $z_{0}\in{\rm Per}(f)$, we have $\lvert\rho_{f}(z_{0})\rvert\in\overline{\mathbb{Q}}$, so $f$ is defined over $\overline{\mathbb{Q}}$ (hence over some number field) by Theorem 1.2.
+After replacing $K$ by a suitable finite extension, we may assume that both $f$ and $\overline{f}$ are defined over $K$, where $\overline{f}$ denotes the rational map obtained by conjugating the coefficients of $f$.
+According to [10, Theorem 7 and Lemma 9], which are essentially due to Zdunik [35], there exists a sequence $(x_{n})_{n=1}^{\infty}$ of distinct points in ${\rm Per}^{*}(f)$ such that
 $a:=\lim\limits_{n\to\infty}\chi_{f}(x_{n})>\mathcal{L}_{f},$
 where the limit exists and is finite.
-Clearly, $\mathcal{L}_{f}=\mathcal{L}_{\overline{f}}$. For an arbitrary $x\in{\rm Per}(f)$, we have
-$\overline{x}\in{\rm Per}(\overline{f}),n_{f}(x)=n_{\overline{f}}(\overline{x}),\text{ and }\rho_{f}(x)=\overline{\rho_{\overline{f}}(\overline{x})},$
-hence $\chi_{f}(x)=\chi_{\overline{f}}(\overline{x})$.
+Clearly, $\mathcal{L}_{f}=\mathcal{L}_{\overline{f}}$. For any $x\in{\rm Per}(f)$, we have
+$\overline{x}\in{\rm Per}(\overline{f}),\quad n_{f}(x)=n_{\overline{f}}(\overline{x}),\quad\text{and}\quad\rho_{f}(x)=\overline{\rho_{\overline{f}}(\overline{x})},$
+so $\chi_{f}(x)=\chi_{\overline{f}}(\overline{x})$.
 Consider the morphism
@@
 For $n\in\mathbb{Z}_{>0}$, set $p_{n}=(x_{n},\overline{x_{n}})\in{\rm Per}(F)$.
-Let $\Gamma$ be the Zariski closure of $\{p_{n}\mid n\in\mathbb{Z}_{>0}\}$ in $\mathbb{P}^{1}\times\mathbb{P}^{1}$, which is defined over $\overline{\mathbb{Q}}$ since the $p_{n}$ are over $\overline{\mathbb{Q}}$.
+Let $\Gamma$ be the Zariski closure of $\{p_{n}\colon n\in\mathbb{Z}_{>0}\}$ in $\mathbb{P}^{1}\times\mathbb{P}^{1}$, which is defined over $\overline{\mathbb{Q}}$ since the $p_{n}$ are defined over $\overline{\mathbb{Q}}$.
 As the points $p_{n}$ are pairwise distinct, we have $\dim(\Gamma)\geq 1$.
-After replacing $(p_{n})$ by a suitable subsequence, we may assume that $\Gamma$ is irreducible and that $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$, i.e., no proper subvariety of $\Gamma$ contains an infinite number of $p_{n}$.
-Let $\pi_{j}:\Gamma\to\mathbb{P}^{1}$ be the $j$-th projection on $\Gamma\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$ to the $j$-th factor of $\mathbb{P}^{1}\times\mathbb{P}^{1}$, for $j=1,2$.
+After passing to a subsequence of $(p_{n})$, we may assume that $\Gamma$ is irreducible and that $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$, i.e., no proper subvariety of $\Gamma$ contains infinitely many $p_{n}$.
+Let $\pi_{j}:\Gamma\to\mathbb{P}^{1}$ be the $j$-th projection from $\Gamma\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$ to the $j$-th factor, for $j=1,2$.
 It is clear that $\pi_{1}\circ F=f\circ\pi_{1}$ and $\pi_{2}\circ F=\overline{f}\circ\pi_{2}$.
-There are 2 cases to consider, i.e., $\dim(\Gamma)=2$ or $\dim(\Gamma)=1$.
-When $\dim(\Gamma)=2$, then $\Gamma=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and the canonical $F$-invariant probability measure on the projective complex manifold $\Gamma(\mathbb{C})=\mathbb{P}^{1}(\mathbb{C})\times\mathbb{P}^{1}(\mathbb{C})$ is $\mu:=\mu_{f}\times\mu_{\overline{f}}$, where $\mu_{f}$ and $\mu_{\overline{f}}$ are the canonical probability measures on $\mathbb{P}^{1}(\mathbb{C})$ which are $f$-invariant and $\overline{f}$-invariant, respectively.
+There are 2 cases to consider, i.e., $\dim(\Gamma)=2$ and $\dim(\Gamma)=1$.
+If $\dim(\Gamma)=2$, then $\Gamma=\mathbb{P}^{1}\times\mathbb{P}^{1}$, and the canonical $F$-invariant probability measure on the projective complex manifold $\Gamma(\mathbb{C})=\mathbb{P}^{1}(\mathbb{C})\times\mathbb{P}^{1}(\mathbb{C})$ is $\mu:=\mu_{f}\times\mu_{\overline{f}}$, where $\mu_{f}$ and $\mu_{\overline{f}}$ are the canonical probability measures on $\mathbb{P}^{1}(\mathbb{C})$ which are $f$-invariant and $\overline{f}$-invariant, respectively.
 (Equivalently, $\mu_{f}$ and $\mu_{\overline{f}}$ are the unique maximal entropy measures of $f$ and $\overline{f}$, respectively.)
-When $\dim(\Gamma)=1$, by the solution [8] of the dynamical Manin-Mumford problem for $F$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$, the irreducible curve $\Gamma$ is periodic under $F$.
-After replacing $f$ by $f^{m}$ for some suitable $m\in\mathbb{Z}_{>0}$, we may assume that the irreducible curve $\Gamma$ is $F$-invariant.
+If $\dim(\Gamma)=1$, by the solution [8] of the dynamical Manin-Mumford problem for $F$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$, the irreducible curve $\Gamma$ is periodic under $F$.
+After replacing $f$ by $f^{m}$ for a suitable $m\in\mathbb{Z}_{>0}$, we may assume that the irreducible curve $\Gamma$ is $F$-invariant.
 Still denote by $\mu$ the canonical $F$-invariant probability measure on the projective complex analytic space $\Gamma(\mathbb{C})$.
 Recall that if $g:X\to X$ is a polarized surjective endomorphism on a projective variety $X$ over $\mathbb{C}$ (hence $g$ is finite), then the canonical $g$-invariant probability measure $\mu_{g}$ on the projective complex analytic space $X(\mathbb{C})$ is exactly the unique probability measure with continuous potentials on $X(\mathbb{C})$ such that $g_{*}\mu_{g}=\mu_{g}$ and $g^{*}\mu_{g}=\deg(g)^{\dim(X)}\mu_{g}$.
-Since $f\circ\pi_{1}=\pi_{1}\circ F$ and $\overline{f}\circ\pi_{2}=\pi_{2}\circ F$ are commuting finite morphisms, via the Tate’s limiting construction of canonical measures, it is easy to deduce that
+Since $f\circ\pi_{1}=\pi_{1}\circ F$ and $\overline{f}\circ\pi_{2}=\pi_{2}\circ F$ are commuting finite morphisms, Tate’s limiting construction of canonical measures implies that
 $\deg(\pi_{1})\mu=\pi_{1}^{*}(\mu_{f})\text{ and }\deg(\pi_{2})\mu=\pi_{2}^{*}(\mu_{\overline{f}}).$
-[…]
-For $n\in\mathbb{Z}_{>0}$, we consider the probability measure
+Now we deal with both cases ($\dim(\Gamma)=2$ and $\dim(\Gamma)=1$) simultaneously. We will show that $2\mathcal{L}_{f}<\int\log\lvert\det(dF)\rvert d\mu$.
+For $n\in\mathbb{Z}_{>0}$, consider the probability measure
 $\nu_{n}=\frac{1}{n_{f}(x_{n})[K_{n}:K]}\sum_{j=0}^{n_{f}(x_{n})-1}\sum_{\tau\in{\rm Gal}(K_{n}/K)}\delta_{F^{j}(\tau(p_{n}))},$
@@
 Then $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$.
-By [36, §2], there exists a unique semipositive metric over $\mathcal{L}$ making $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$ an isometry; denote $\mathcal{L}$ with this metric by $\overline{\mathcal{L}}$.
-The canonical measure $\mu_{\overline{\mathcal{L}}}$ associated with the metrized line bundle $\overline{\mathcal{L}}$ on $\Gamma(\mathbb{C})$ is exactly the canonical $F$-invariant probability measure $\mu$ on $\Gamma(\mathbb{C})$ considered before, since the metric of ${\overline{\mathcal{L}}}$ can be obtained by Tate’s limit process like $\mu$, see [36, §2] and [32, §3.5].
-We need to check the conditions (1) (small) and (2) (generic) in Theorem 3.1.
-The condition (1) is trivial, since $h_{\overline{\mathcal{L}}}(\Gamma)=0$ and the height of every periodic algebraic point relative to $\overline{\mathcal{L}}$ is zero, in the setting of a polarized algebraic dynamical system (cf.
+By [36, §2], there exists a unique semipositive metric on $\mathcal{L}$ making $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$ an isometry; denote this metrized line bundle by $\overline{\mathcal{L}}$.
+The canonical measure $\mu_{\overline{\mathcal{L}}}$ associated with the metrized line bundle $\overline{\mathcal{L}}$ on $\Gamma(\mathbb{C})$ is exactly the canonical $F$-invariant probability measure $\mu$ on $\Gamma(\mathbb{C})$, since the metric of $\overline{\mathcal{L}}$} can be obtained by Tate’s limit process, similar to $\mu$; see [36, §2] and [32, §3.5].
+We need to verify the conditions (1) (small) and (2) (generic) in Theorem 3.1.
+Condition (1) is trivial because $h_{\overline{\mathcal{L}}}(\Gamma)=0$ and the height of every periodic algebraic point relative to $\overline{\mathcal{L}}$ is zero in the setting of a polarized algebraic dynamical system (cf.
 [32]).
-For the condition (2), let $V$ be an arbitrary proper subvariety of $\Gamma$ and fix $V$.
+For the condition (2), let $V$ be an arbitrary proper subvariety of $\Gamma$.
 After replacing $V$ by its Galois orbit, we may assume that $V$ is defined over $K$.
-Then the generic condition (2) follows from Lemma 3.3 immediately.
-Thus the claim is true.
+Then the generic condition (2) follows immediately from Lemma 3.3.
+Thus the claim holds.
 Let $n\in\mathbb{Z}_{>0}$, take $m\in\mathbb{Z}_{>0}$ such that $\lvert\rho_{f}(x_{n})\rvert^{m}\in K$ by the assumption, and write $l=n_{f}(x_{n})$.
-For every $\tau\in{\rm Gal}(\overline{\mathbb{Q}}/K)$ and $10\leq j\leq l-$ 1, we have
+For every $\tau\in{\rm Gal}(\overline{\mathbb{Q}}/K)$ and $0\leq j\leq l-1$, we have
 $\det(dF^{l}(F^{j}(\tau(p_{n}))))^{m}=\det(dF^{l}(\tau(p_{n})))^{m}$
@@
 $=$ $\lvert\rho_{f}(x_{n})\rvert^{2m},$
-where the first equality holds because both $\tau(p_{n})$ and $F^{j}(\tau(p_{n}))$ are in the same $F$-periodic orbit of length $l$ and the appearance $dF^{l}=d(F^{l})$ in the expressions, hence $\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert=\lvert\rho_{f}(x_{n})\rvert^{2}$.
+where the first equality holds because both $\tau(p_{n})$ and $F^{j}(\tau(p_{n}))$ belong to the same $F$-periodic orbit of length $l$, and $dF^{l}=d(F^{l})$ in the expressions.
+Hence $\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert=\lvert\rho_{f}(x_{n})\rvert^{2}$.
 Then by the definition of $\nu_{n}$, we have
-[…]
+$\int\log\lvert\det(dF)\rvert d\nu_{n}=\frac{1}{l}\int\log\lvert\det(dF^{l})\rvert d\nu_{n}$
 $=$ $\frac{1}{l^{2}[K_{n}:K]}\sum_{j=0}^{l-1}\sum_{\tau\in{\rm Gal}(K_{n}/K)}\log\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert$
@@
 For every $A\in\mathbb{R}$, since the function $\max\{\log\lvert\det(dF)\rvert,A\}$ is continuous, we have
-[…]
-[…]
-[…]
-Let $A\to-\infty$, by the monotone convergence theorem, we have
-[…]
-We will show that (4.14) leads to a contradiction, by considering the case where $\dim(\Gamma)=2$ and the case where $\dim(\Gamma)=1$ separately.
-When $\dim(\Gamma)=2$, it is clear that
-[…]
+$2a=2\lim\limits_{n\to\infty}\chi_{f}(x_{n})=\lim\limits_{n\to\infty}\int\log\lvert\det(dF)\rvert d\nu_{n}$
+$\leq$ $\lim\limits_{n\to\infty}\int\max\{\log\lvert\det(dF)\rvert,A\}d\nu_{n}$
+$=$ $\int\max\{\log\lvert\det(dF)\rvert,A\}d\mu.$
+Letting $A\to$-$\infty$, by the monotone convergence theorem, we have
+(4.14) $2\mathcal{L}_{f}<2a\leq\int\log\lvert\det(dF)\rvert d\mu.$
+We now show that (4.14) leads to a contradiction by considering the two cases separately.
+If $\dim(\Gamma)=2$, then
+$\int\log\lvert\det(dF)\rvert d\mu=\int_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\log\lvert\det(d(f\times\overline{f}))\rvert d(\mu_{f}\times\mu_{\overline{f}})=\mathcal{L}_{f}+\mathcal{L}_{\overline{f}}=2\mathcal{L}_{f},$
 contradicting (4.14).
-When $\dim(\Gamma)=1$, then
-[…]
-[…]
-[…]
-[…]
+If $\dim(\Gamma)=1$, then
+$\int\log\lvert\det(dF)\rvert d\mu=\int_{\Gamma}\log\lvert\det(d(f\times\overline{f}))\rvert d\mu$
+$=$ $\int_{\Gamma}\log\lvert\det(df)\circ\pi_{1}\rvert d\mu+\int_{\Gamma}\log\lvert\det(d\overline{f})\circ\pi_{2}\rvert d\mu$
+$=$ $\int_{\Gamma}\log\lvert\det(df)\circ\pi_{1}\rvert d\frac{\pi_{1}^{*}(\mu_{f})}{\deg(\pi_{1})}+\int_{\Gamma}\log\lvert\det(d\overline{f})\circ\pi_{2}\rvert d\frac{\pi_{2}^{*}(\mu_{\overline{f}})}{\deg(\pi_{2})}$
+$=$ $\int_{\mathbb{P}^{1}}\log\lvert\det(df)\rvert d\mu_{f}+\int_{\mathbb{P}^{1}}\log\lvert\det(d\overline{f})\rvert d\mu_{\overline{f}}$
 $=$ $\mathcal{L}_{f}+\mathcal{L}_{\overline{f}}=2\mathcal{L}_{f},$
 contradicting (4.14).
-Therefore, $f$ must be exceptional. We have finished the proof. ∎
+Therefore, $f$ must be exceptional. This completes the proof. ∎
 Proof of Theorem 1.4 when $f$ is defined over $\overline{\mathbb{Q}}$.
-Assume that $f$ is defined over $\overline{\mathbb{Q}}$, hence $f$ is defined over a number field $K$.
-Suppose that the Theorem 1.4 does not hold for $f$.
-Let $V$ be the $\mathbb{Q}$-span of $\chi_{f}({\rm Per}^{*}(f))$ in $\mathbb{R}$, then $\dim_{\mathbb{Q}}V<\infty$.
+Assume that $f$ is defined over $\overline{\mathbb{Q}}$, hence $f$ is defined over some number field $K$.
+Suppose that Theorem 1.4 fails for $f$.
+Let $V$ be the $\mathbb{Q}$-span of $\chi_{f}({\rm Per}^{*}(f))$ in $\mathbb{R}$; then $\dim_{\mathbb{Q}}V<\infty$.
 We can take $M\in\mathbb{Z}_{>0}$ and $x_{1},\dots,x_{M}\in{\rm Per}^{*}(f)$ such that $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ generate $V$ over $\mathbb{Q}$.
 By enlarging $K$, we may assume that $\lvert\rho_{f}(x_{1})\rvert,\dots,\lvert\rho_{f}(x_{M})\rvert\in K$.
-Then for every $z_{0}\in{\rm Per}^{*}(f)$, its characteristic exponent $\chi_{f}(z_{0})$ is a linear combination of $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ over $\mathbb{Q}$; in other words, there exist $n\in\mathbb{Z}_{>0}$ and $n_{1},\dots,n_{M}\in\mathbb{Z}$ such that
+Then for every $z_{0}\in{\rm Per}^{*}(f)$, its characteristic exponent $\chi_{f}(z_{0})$ is a $\mathbb{Q}$-linear combination of $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$; in other words, there exist $n\in\mathbb{Z}_{>0}$ and $n_{1},\dots,n_{M}\in\mathbb{Z}$ such that
 $\lvert\rho_{f}(z_{0})\rvert^{n}=\lvert\rho_{f}(x_{1})\rvert^{n_{1}}\cdots\lvert\rho_{f}(x_{M})\rvert^{n_{M}}\in K,$
-which contradicts Theorem 1.5 since $f$ is not exceptional by the assumption. ∎
+which contradicts Theorem 1.5 since $f$ is not exceptional by assumption. ∎
 Remark 4.1.
-We will finish the whole proof of Theorem 1.4 in §6.3.
-According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps.
+We will complete the whole proof of Theorem 1.4 in §6.3.
+According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ within the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps.
 Therefore, the above proof of Theorem 1.4 for $f$ defined over $\overline{\mathbb{Q}}$ includes the special case where $f$ is PCF (and non-exceptional).
@@
 For $f,g\in{\rm PHom}(V,W)$, we define $f+g$ to be the pair $(V_{f}\cap V_{g},f|_{V_{f}\cap V_{g}}+g|_{V_{f}\cap V_{g}})$.
-Then ${\rm PHom}(V,W)$ is a commutative semigroup with $+$ as the operation.
-We denote by 00 the pair $(V,0)$.
+We denote by 0 the pair $(V,0)$.
 We have $0+f=f$ for all $f\in{\rm PHom}(V,W)$.
+Then ${\rm PHom}(V,W)$ is a commutative monoid under the operation ++ with the zero element 0.
 For every $a\in\mathbb{R}$, we define $af$ to be the pair $(V_{f},af)$.
-We note that $f+(-f)=(V_{f},0)$, which is not 00 if $V_{f}\neq V$.
+Note that $f+(-f)=(V_{f},0)$, which is not 0 if $V_{f}\neq V$.
 We have a natural embedding ${\rm Hom}(V,W)\hookrightarrow{\rm PHom}(V,W)$.
@@
 Definition 5.1.
-A sequence $(f_{i})_{i=1}^{\infty}$ in ${\rm PHom}(V,\mathbb{R})$ is said to be an $O$-sequence if the following conditions are satisfied:
-(i) $f_{i}(O)\subseteq\mathbb{R}_{\geq 0}\cup\{\infty\}$ for $i\geq 1$; (ii) for every $\lambda\in O$, the set $\{i\geq 1:f_{i}(\lambda)\neq 0\}$ is finite.
-Clearly, an infinite subsequence of an $O$-sequence is still an $O$-sequence.
+A sequence $(f_{i})_{i=1}^{\infty}$ in ${\rm PHom}(V,\mathbb{R})$ is called an $O$-sequence if the following conditions hold:
+  1. (i)
+    $f_{i}(O)\subseteq\mathbb{R}_{\geq 0}\cup\{\infty\}$ for $i\geq 1$;
+  2. (ii)
+    for every $\lambda\in O$, the set $\{i\geq 1:f_{i}(\lambda)\neq 0\}$ is finite.
+Clearly, any infinite subsequence of an $O$-sequence is also an $O$-sequence.
 Definition 5.2.
 Let $(\lambda_{i})_{i=1}^{\infty}$ be a sequence in $O$ and $(f_{i})_{i=1}^{\infty}$ be a sequence in ${\rm PHom}(V,\mathbb{R})$.
-We say that $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ is an upper triangle $O$-system (resp.
-weak upper triangle $O$-system) if the following conditions hold:
-(i) $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence;
-(ii) $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}$ (resp. $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}\cup\{\infty\})$ for $i\geq 1$;
-(iii) $f_{j}(\lambda_{i})=0$ for $j>i\geq 1$.
-Clearly, an upper triangle $O$-system is a weak upper triangle $O$-system.
+We say that $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ is an upper triangular $O$-system (resp.
+weak upper triangular $O$-system) if the following conditions hold:
+  1. (i)
+    $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence;
+  2. (ii)
+    $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}$ (resp. $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}\cup\{\infty\}$) for $i\geq 1$;
+  3. (iii)
+    $f_{j}(\lambda_{i})=0$ for $j>i\geq 1$.
+Clearly, an upper triangular $O$-system is a weak upper triangular $O$-system.
 Lemma 5.3.
-Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be a weak upper triangle $O$-system. Then $(\lambda_{i})_{i=1}^{\infty}$ are linearly independent over $\mathbb{R}$.
+Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be a weak upper triangular $O$-system. Then $(\lambda_{i})_{i=1}^{\infty}$ are linearly independent over $\mathbb{R}$.
 Proof.
-Since $f_{1}(\lambda_{1})\neq 0$, we see that $\lambda_{1}\neq 0$.
-Then we only need to show that for all $l\geq 2$, $\lambda_{l}$ is not contained in ${\rm span}_{\mathbb{R}}\{\lambda_{i}:i\leq l-1\}$.
-Otherwise, $\lambda_{l}=\sum_{i=1}^{l-1}a_{i}\lambda_{i}$ for some $l\geq 2,a_{i}\in\mathbb{R},1\leq i\leq l-1$.
-Then $f_{l}(\lambda_{l})=\sum_{i=1}^{l-1}a_{i}f_{l}(\lambda_{i})=0$, which contradicts to our assumption.
+Since $f_{1}(\lambda_{1})\neq 0$, we have $\lambda_{1}\neq 0$.
+It remains to show that for all $l\geq 2$, $\lambda_{l}$ is not contained in ${\rm span}_{\mathbb{R}}\{\lambda_{i}:i\leq l-1\}$.
+Suppose otherwise that $\lambda_{l}=\sum_{i=1}^{l-1}a_{i}\lambda_{i}$ for some $l\geq 2$ and $a_{i}\in\mathbb{R}$, $1\leq i\leq l-1$.
+Then $f_{l}(\lambda_{l})=\sum_{i=1}^{l-1}a_{i}f_{l}(\lambda_{i})=0$, contradicting our assumption.
 ∎
-Let $\tau:V\to V$ be an involution (here we just means that $\tau^{2}=\mathrm{id}$ and $\tau$ may be the identity on $V$).
+Let $\tau:V\to V$ be an involution (here we only assume that that $\tau^{2}=\mathrm{Id}$; $\tau$ may be the identity on $V$).
 Lemma 5.4.
 Assume that $\tau(O)\subseteq O$.
-Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangle $O$-system.
-Then there exists a strictly increasing sequence $(m_{i})_{i=1}^{\infty}$ in $\mathbb{Z}_{>0}$ such that the pair $((\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))_{i=1}^{\infty},(f_{m_{i}})_{i=1}^{\infty})$ is a weak upper triangle $O^{\prime}$-system, where $O^{\prime}=\{\lambda_{m_{i}}+\tau(\lambda_{m_{i}}):i\in\mathbb{Z}_{>0}\}$.
+Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangular $O$-system.
+Then there exists a strictly increasing sequence $(m_{i})_{i=1}^{\infty}$ in $\mathbb{Z}_{>0}$ such that the pair $((\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))_{i=1}^{\infty},(f_{m_{i}})_{i=1}^{\infty})$ is a weak upper triangular $O^{\prime}$-system, where $O^{\prime}=\{\lambda_{m_{i}}+\tau(\lambda_{m_{i}}):i\in\mathbb{Z}_{>0}\}$.
 Proof.
@@
 $f_{m_{1}}(\lambda_{m_{1}}+\tau(\lambda_{m_{1}}))=f_{1}(\lambda_{1})+f_{1}(\tau(\lambda_{1}))\in\mathbb{R}_{>0}\cup\{\infty\}.$
-Assume that we have constructed $m_{1},\dots,m_{l}$ satisfying the conditions for weak upper triangle $O^{\prime}$-systems.
-Since $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence and $\tau(\lambda_{1}),\cdots,\tau(\lambda_{l})\in O$, there exists $m_{l+1}>m_{l}$ such that $f_{m_{l+1}}(\tau(\lambda_{i}))=0$ for all $i=1,\dots,l$.
+Assume that we have constructed $m_{1},\dots,m_{l}$ satisfying the conditions for weak upper triangular $O^{\prime}$-systems.
+Since $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence and $\tau(\lambda_{1}),\dots,\tau(\lambda_{l})\in O$, there exists $m_{l+1}>m_{l}$ such that $f_{m_{l+1}}(\tau(\lambda_{m_{i}}))=0$ for all $1\leq i\leq l$.
 Then for all $i=1,\dots,l$, we have
 $f_{m_{l+1}}(\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))=f_{m_{l+1}}(\lambda_{m_{i}})+f_{m_{l+1}}(\tau(\lambda_{m_{i}}))=0.$
-Also,
+Moreover,
 $f_{m_{l+1}}(\lambda_{m_{l+1}}+\tau(\lambda_{m_{l+1}}))=f_{m_{l+1}}(\lambda_{m_{l+1}})+f_{m_{l+1}}(\tau(\lambda_{m_{l+1}}))\in\mathbb{R}_{>0}\cup\{\infty\}$
@@
 $f_{m_{i}}(\lambda_{m_{l+1}}+\tau(\lambda_{m_{l+1}}))=f_{m_{i}}(\lambda_{m_{l+1}})+f_{m_{i}}(\tau(\lambda_{m_{l+1}}))\in\mathbb{R}_{\geq 0}\cup\{\infty\}.$
-We conclude the proof. ∎
-By Lemma 5.4 and Lemma 5.3 we get the following result.
+This completes the proof. ∎
+Combining Lemma 5.4 and Lemma 5.3, we obtain:
 Corollary 5.5.
 Assume that $\tau(O)\subseteq O$.
-Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangle $O$-system.
+Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangular $O$-system.
 Then $\dim_{\mathbb{R}}{\rm span}_{\mathbb{R}}\{2^{-1}(\lambda_{i}+\tau(\lambda_{i})):i\geq 1\}=\infty$.
 Remark 5.6.
-Note that the discussion in this subsection also applies with $\mathbb{R}$ replaced by any ordered field $F$ (for example, $\mathbb{Q}$ with the usual order).
+The discussion in this subsection also applies when $\mathbb{R}$ is replaced by any ordered field $F$ (for example, $\mathbb{Q}$ with the usual order).
 5.2. Linear algebra for multiplication
-For every field $k$ of characteristic 00, denote by $\mu_{k}$ the subgroup of roots of unity in $k$. Denote by
+For every field $k$ of characteristic 0, denote by $\mu_{k}$ the subgroup of roots of unity in $k$. Let
 ${\rm rog}:k^{*}\to\mathbb{D}(k):=k^{*}/\mu_{k}$
-the quotient map.
-Extend ${\rm rog}$ to a map ${\rm rog}:k\to\mathbb{D}(k)\cup\{\infty\}$ by sending 00 to $\infty$.
-Here we use the notation “${\rm rog}$” since it is an analogy of the classical $\log$ function to some extent.
-The embedding $k\hookrightarrow\overline{k}$ gives a natural embedding $\mathbb{D}(k)\hookrightarrow\mathbb{D}(\overline{k})$ as multiplicative abelian groups.
-Set $\mathbb{D}(k)_{\mathbb{Q}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{Q}$, where $\mathbb{D}(k)$ is viewed as a multiplicative abelian group, hence a $\mathbb{Z}$-module.
+be the quotient map.
+Extend $\rm rog$} to a map ${\rm rog}:k\to\mathbb{D}(k)\cup\{\infty\}$ by sending 0 to $\infty$.
+(The notation $\rm rog$}” is used by analogy with the classical $\log$ function.) The embedding $k\hookrightarrow\overline{k}$ induces a natural embedding $\mathbb{D}(k)\hookrightarrow\mathbb{D}(\overline{k})$ of multiplicative abelian groups.
+Set $\mathbb{D}(k)_{\mathbb{Q}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{Q}$, where $\mathbb{D}(k)$ is viewed as a multiplicative abelian group (hence a $\mathbb{Z}$-module).
 Then $\mathbb{D}(k)_{\mathbb{Q}}$ is the subspace of $\mathbb{D}(\overline{k})$ spanned by $\mathbb{D}(k)$ over $\mathbb{Q}$.
@@
 5.3. Norms
-Let $k$ be a field of characteristic 00. For every finite field extension $\widetilde{k}$ of $k$, denote by $N_{\widetilde{k}/k}:\widetilde{k}\to k$ the norm map. The norm map induces a map
-$\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{Q}}\to\mathbb{D}(k)_{\mathbb{Q}},{\rm rog}(x)\mapsto[l:k]^{-1}{\rm rog}(N_{l/k}(x)),$
-where $l$ is any finite field extension of $k$ containing $x$.
-One can check that $\mathbf{n}_{k}$ is well defined and is $\mathbb{Q}$-linear.
+Let $k$ be a field of characteristic 0. For every finite extension $\widetilde{k}$ of $k$, denote by $N_{\widetilde{k}/k}:\widetilde{k}\to k$ the norm map. The norm map induces a map
+$\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{Q}}\to\mathbb{D}(k)_{\mathbb{Q}},{\rm rog}(x)\mapsto[k^{\prime}:k]^{-1}{\rm rog}(N_{k^{\prime}/k}(x)),$
+where $k^{\prime}$ is any finite extension of $k$ containing $x$.
+One can check that $\mathbf{n}_{k}$ is well-defined and $\mathbb{Q}$-linear.
 We also denote by $\mathbf{n}_{k}$ its $\mathbb{R}$-linear extension $\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{R}}\to\mathbb{D}(k)_{\mathbb{R}}$.
-When the field $k$ is clear, we also write $\mathbf{n}$ for $\mathbf{n}_{k}$.
+When the field $k$ is clear, we write $\mathbf{n}$ for $\mathbf{n}_{k}$.
 5.4. Valuations
@@
 Denote by $\mathcal{M}_{K}$ the set of all places of $K$.
-For every $v\in\mathcal{M}_{K}$, we define a map $\phi_{v}:\mathbb{D}(K)_{\mathbb{R}}\to\mathbb{R}$, by $\phi_{v}({\rm rog}(x))=-\log(\lvert x\rvert_{v})$ for $x\in K^{*}$ and extending $\mathbb{R}$-linearly, where $\left|\cdot\right|_{v}$ is a fixed multiplicative norm on $K$ corresponding to $v$.
+For every $v\in\mathcal{M}_{K}$, we define a map $\phi_{v}:\mathbb{D}(K)_{\mathbb{R}}\to\mathbb{R}$ by $\phi_{v}({\rm rog}(x))=-\log(\lvert x\rvert_{v})$ for $x\in K^{*}$ and extending $\mathbb{R}$-linearly, where $\left|\cdot\right|_{v}$ is a fixed multiplicative norm on $K$ corresponding to $v$.
 It is easy to check that the map $\phi_{v}$ is well-defined and $\mathbb{R}$-linear.
@@
 For every $a\in\mathbb{D}(K)_{\mathbb{R}}$, it is clear that the set $\{v\in\mathcal{M}_{K}:\phi_{v}(a)\neq 0\}$ is finite.
-Let $S$ be a finite subset of $\mathcal{M}_{K}$ containing all the archimedean places.
+Let $S$ be a finite subset of $\mathcal{M}_{K}$ containing all archimedean places.
 Let $\mathcal{O}_{K,S}$ be the ring of $S$-integers in $K$.
@@
 $\mathbf{n}_{\mathbb{R}}:\mathbb{D}(\mathbb{C})\to\mathbb{D}(\mathbb{R}),{\rm rog}(x)\mapsto\frac{{\rm rog}(x)+{\rm rog}(\tau(x))}{2}.$
-Let $\mathbf{k}$ be an algebraically closed subfield of $\mathbb{C}$ stable under the complex conjugation $\tau$.
-Still denote by $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ the restriction of the complex conjugation on $\mathbf{k}$, which is an involution of $\mathbf{k}$.
-Denote by $\mathbf{k}^{\tau}=\mathbf{k}\cap\mathbb{R}$ the $\tau$-invariant subfield of $\mathbf{k}$.
+Let $\mathbf{k}$ be an algebraically closed subfield of $\mathbb{C}$ stable under complex conjugation $\tau$.
+Still denote by $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ the restriction of complex conjugation on $\mathbf{k}$, which is an involution of $\mathbf{k}$.
+Let $\mathbf{k}^{\tau}=\mathbf{k}\cap\mathbb{R}$ be the $\tau$-invariant subfield of $\mathbf{k}$.
 Then the restriction of the absolute value $\mathbf{n}_{\mathbb{R}}$ on $\mathbf{k}$ is
 $\mathbf{n}_{\mathbf{k}^{\tau}}:\mathbb{D}(\mathbf{k})\to\mathbb{D}(\mathbf{k}^{\tau}),{\rm rog}(x)\mapsto\frac{{\rm rog}(x)+{\rm rog}(\tau(x))}{2}.$
-We shall prove the following theorem involving an involution $\tau$ (which may not be the complex conjugation and may be the identity) in the next section.
+We shall prove the following theorem involving an involution $\tau$ (which may not be complex conjugation and may be the identity) in the next section.
 Theorem 5.7.
-Let $\mathbf{k}$ be an algebraically closed field of characteristic 00.
+Let $\mathbf{k}$ be an algebraically closed field of characteristic 0.
 Let $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ be an element with $\tau^{2}=\mathrm{Id}$.
-If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace in $\mathbb{D}(\mathbf{k}^{\tau})_{\mathbb{Q}}$ spanned by $\{\mathbf{n}_{\mathbf{k}^{\tau}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
-Applying Theorem 5.7 to the case where $\mathbf{k}=\mathbb{C}$ and $\tau$ is the complex conjugation, we deduce Theorem 1.4 in the case where $f$ is not PCF.
+If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbf{k}^{\tau})_{\mathbb{Q}}$ spanned by $\{\mathbf{n}_{\mathbf{k}^{\tau}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
+Applying Theorem 5.7 to the case $\mathbf{k}=\mathbb{C}$ and $\tau$ being complex conjugation, we deduce Theorem 1.4 in the case where $f$ is not PCF.
 Remark 5.8.
-Setting $\tau=\mathrm{id}$, then from Theorem 5.7 we get the following result:
-Assume that $\mathbf{k}$ is an algebraically closed field of characteristic 00.
-If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace in $\mathbb{D}(\mathbf{k})_{\mathbb{Q}}$ spanned by $\{{\rm rog}(\rho_{f}(x)):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
+By setting $\tau=\mathrm{Id}$, Theorem 5.7 yields the following result:
+Assume that $\mathbf{k}$ is an algebraically closed field of characteristic 0.
+If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbf{k})_{\mathbb{Q}}$ spanned by $\{{\rm rog}(\rho_{f}(x)):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
 6. Proofs of Theorem 5.7 and Theorem 1.4
@@
 Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ such that the (forward) $f$-orbit $O_{f}(o)$ of $o$ is infinite.
-We fix such a critical point $o$.
+Fix such a critical point $o$.
 Let $X$ be the union of all (forward) orbits of periodic critical points of $f$.
@@
 Denote by $\mathcal{M}_{K}$ the set of all places of $K$.
-Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all the archimedean places, satisfying $\tau(B)=B$, such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$.
-Then we have $\tau(\mathcal{O}_{K,B})=\mathcal{O}_{K,B}$.
+Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all archimedean places, satisfying $\tau(B)=B$, and such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$.
+Then $\tau(\mathcal{O}_{K,B})=\mathcal{O}_{K,B}$.
 For $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$, set $\lambda(x)=(n_{f}(x))^{-1}{\rm rog}(\rho_{f}(x))\in\mathbb{D}(\overline{\mathbb{Q}})_{\mathbb{R}}\cup\{\infty\}$.
@@
 Observe that for every $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$, we have $\sigma(\lambda(x))=\lambda(\sigma(x))$.
-For every $v\in\mathcal{M}_{K}\setminus B$ and $x\in\mathbb{P}^{1}(\mathbb{C}_{v})$, denote by $\tilde{x}\in\mathbb{P}^{1}(\overline{\widetilde{K_{v}}})$ the reduction of $x$ in the special fiber at $v$ and $f_{v}:\mathbb{P}^{1}_{\overline{\widetilde{K_{v}}}}\to\mathbb{P}^{1}_{\overline{\widetilde{K_{v}}}}$ the reduction of $f$.
-(Here, $\overline{\widetilde{K_{v}}}$ is a fixed algebraic closure of the residue field $\widetilde{K_{v}}$ of $K_{v}$.) After enlarging $B$, we can assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$, where $X_{v}$ is the reduction of $X$ contained in $\mathbb{P}^{1}(\widetilde{K_{v}})\subseteq\mathbb{P}^{1}(\overline{\widetilde{K_{v}}})$.
+[…]
 Observe that for every $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$ of exact period $n\geq 1$ and every $v\in\mathcal{M}_{K}\setminus B$, we have $\phi_{v}(\mathbf{n}_{K}(\lambda(x)))\geq 0$ by [24, Corollary 2.23 (a)].
@@
 (ii) there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ such that $(f^{n}_{v})^{\prime}(\widetilde{\sigma(x)})=0$;
-(iii) there exist an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$, $q\in\mathcal{C}_{f}$, and $m\in\mathbb{Z}_{\geq 0}$, such that $\widetilde{\sigma(q)}$ is $f_{v}$-periodic with $\widetilde{\sigma(x)}=f_{v}^{m}(\widetilde{\sigma(q)})$.
+[…]
 For every $v\in\mathcal{M}_{K}\setminus B$, denote by $P_{v}$ the union of all $f_{v}$-orbits of periodic critical points of $f_{v}$.
@@
 For every $v\in\mathcal{M}_{K}\setminus B$ and $q\in P_{v}$, by Hensel’s lemma there exists a unique periodic point $y=y(q)\in{\rm Per}(f)(\mathbb{C}_{v}\cap\overline{\mathbb{Q}})$ such that $\tilde{y}=q$.
-Then there exists a unique ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit $O(q)$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ such that for some (then every) $x\in O(q)$, there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ satisfying $\widetilde{\sigma(x)}=q$ (here $O(q)$ is the ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit of $y=y(q)$).
+Then there exists a unique ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit $O(q)$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ such that for some (hence every) $x\in O(q)$, there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ satisfying $\widetilde{\sigma(x)}=q$ (here $O(q)$ is the ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit of $y=y(q)$).
 In particular, we have $X_{v}\subseteq P_{v}$ and $\cup_{q\in X_{v}}O(q)=X$.
@@
 Proof.
-Recall that $o\in\mathcal{C}_{f}$ is a critical point of $f$ which is not $f$-preperiodic and we assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$.
-By [3, Lemma 4.1], there are infinitely may $v\in\mathcal{M}_{K}\setminus B$, such that there exists $n\in\mathbb{Z}_{>0}$ with $f_{v}^{n}(\tilde{o})=\tilde{o}$.
-For such a place $v$, we have $\tilde{o}\in P_{v}\setminus X_{v}$, which concludes the proof.
+Recall that $o\in\mathcal{C}_{f}$ is a critical point of $f$ that is not $f$-preperiodic, and we assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$.
+By [3, Lemma 4.1], there are infinitely many $v\in\mathcal{M}_{K}\setminus B$ such that there exists $n\in\mathbb{Z}_{>0}$ with $f_{v}^{n}(\tilde{o})=\tilde{o}$.
+For such a place $v$, we have $\tilde{o}\in P_{v}\setminus X_{v}$, which proves the lemma.
 ∎
@@
 There exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{i})))>0$ for all $i\geq 1$ and $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))=0$ for $j\neq i$.
-In particular, $((\mathbf{n}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangle $O$-system for $\mathbb{D}(K)_{\mathbb{Q}}$.
+In particular, $((\mathbf{n}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangular $O$-system for $\mathbb{D}(K)_{\mathbb{Q}}$.
 Proof.
-We construct these two sequences recursively.
-By Lemma 6.1, $S$ is infinite.
-Pick $v_{1}\in S$, then there exists $x_{1}\in Q_{v_{1}}\setminus X\subseteq$ ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$.
+We construct these sequences recursively.
+By Lemma 6.1, $S$ is infinite. Pick $v_{1}\in S$, then there exists
+$x_{1}\in Q_{v_{1}}\setminus X\subseteq{\rm Per}^{*}(f)(\overline{\mathbb{Q}}).$
 We have $\phi_{v_{1}}(\mathbf{n}(\lambda(x_{1})))>0$.
-Assume that we have constructs $x_{1},\dots,x_{m}\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and $v_{1},\dots,v_{m}\in$ $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))\geq 0$ and the equality holds if and only if $j\neq i$.
+Assume that we have constructed $x_{1},\dots,x_{m}\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and $v_{1},\dots,v_{m}\in\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))\geq 0$ and equality holds if and only if $j\neq i$.
 The set $\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ is finite.
@@
 By Lemma 6.1, we can take a place $v_{m+1}\in S\setminus(\{v_{1},\dots,v_{m}\}\cup T_{m})$.
-Then we have $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{i})))=0$ for all $1\leq i\leq m$.
+Then $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{i})))=0$ for all $1\leq i\leq m$.
 Pick $x_{m+1}\in Q_{v_{m+1}}\setminus X$.
@@
 ∎
-Then we conclude the proof by Corollary 5.5 and Remark 5.6.∎
+The conclusion now follows from Corollary 5.5 and Remark 5.6.∎
 6.2. Proof of Theorem 5.7: the general case
@@
 Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ which is not $f$-preperiodic.
-We fix such a critical point $o$.
-Fix a subfield $K$ of $\mathbf{k}$ such that $K/\mathbb{Q}$ is finite generated, $\tau(K)=K$, and $o,f$ are defined over $K$.
+Fix such a critical point $o$.
+Fix a subfield $K$ of $\mathbf{k}$ such that $K/\mathbb{Q}$ is finitely generated, $\tau(K)=K$, and $o,f$ are defined over $K$.
 Without loss of generality, we may assume that $\mathbf{k}=\overline{K}$.
-Take a finite generated $\mathbb{Z}$-subalgebra $A$ of $K$ with ${\rm Frac}(A)=K$ and $\tau(A)=A$.
+Take a finitely generated $\mathbb{Z}$-subalgebra $A$ of $K$ with ${\rm Frac}(A)=K$ and $\tau(A)=A$.
 After shrinking ${\rm Spec}(A)$, we may assume that there exists an endomorphism $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ over $A$ whose restriction $f_{K}:\mathbb{P}^{1}_{K}\to\mathbb{P}^{1}_{K}$ on the generic fiber $\mathbb{P}^{1}_{K}$ satisfies $f=f_{K}\otimes_{K}\mathbf{k}$.
@@
 By [9, Lemma 3.3], we can take $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$ such that the $f_{c}$-orbit of $o_{c}$ is infinite, i.e., $o_{c}$ is not $f_{c}$-preperiodic.
-In particular, the specialization map $f_{c}$ is not PCF as well.
-We take a number field $L\subseteq\overline{K}$ such that $c$ is defined over $L$ and $\tau(L)=L$.
+In particular, the specialization map $f_{c}$ is not PCF.
+Take a number field $L\subseteq\overline{K}$ such that $c$ is defined over $L$ and $\tau(L)=L$.
 Denote by $A_{1}$ the $\mathbb{Z}$-algebra generated by $A$, $\mathcal{O}_{L}$, and $\tau(\mathcal{O}_{L})$ in $A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}}$.
-We may replace ${\rm Spec}(A)$ by a suitable non-empty affine Zariski open set of ${\rm Spec}(A_{1})$ such that $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ is still everywhere well-defined and $c\in{\rm Spec}(A\otimes_{\mathcal{O}_{L}}L)(L)$, after the replacement.
+We may replace ${\rm Spec}(A)$ by a suitable non-empty affine Zariski open subset of ${\rm Spec}(A_{1})$ such that $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ remains everywhere well-defined and $c\in{\rm Spec}(A\otimes_{\mathcal{O}_{L}}L)(L)$ after the replacement.
 We view ${\rm Spec}(A)$ as an $\mathcal{O}_{L}$-scheme.
@@
 is also finite. Similarly, $W_{\tau(c)}:=\{y\in{\rm Per}(f)(\overline{K}):s_{\tau(p)}(\mathbf{n}_{K}(\lambda(y)))=\infty\}$ is finite.
-By Lemma 6.2, there exist $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}(f_{c})(\overline{L})$ and $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{L}\setminus S$ such that $((\mathbf{n}_{L}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangle $(\mathbb{D}(\mathcal{O}_{L,S})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(L)_{\mathbb{Q}}$.
+By Lemma 6.2, there exist sequences $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}(f_{c})(\overline{L})$ and $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{L}\setminus S$ such that $((\mathbf{n}_{L}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangular $(\mathbb{D}(\mathcal{O}_{L,S})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(L)_{\mathbb{Q}}$.
 For every $i\in\mathbb{Z}_{>0}$, we can take $y_{i}\in{\rm Per}(f)(\overline{K})$ such that $x_{i}\in(y_{i})_{c}$, and hence
 $s_{p}(\mathbf{n}_{K}(\lambda(y_{i})))=\mathbf{n}_{L}(\lambda(x_{i})).$
-After removing finitely many terms, we can assume that $y_{i}\notin W_{c}\cup W_{\tau(c)}$ for all $i\geq 1$.
+After removing finitely many terms, we may assume that $y_{i}\notin W_{c}\cup W_{\tau(c)}$ for all $i\geq 1$.
 It follows that $\mathbf{n}_{K}(\lambda(y_{i}))\in{\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ for $i\geq 1$.
@@
 $((\mathbf{n}_{K}(\lambda(y_{i})))_{i=1}^{\infty},(\phi_{v_{i}}\circ s_{p})_{i=1}^{\infty})$
-is an upper triangle $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(K)_{\mathbb{Q}}$. Since
+is an upper triangular $({\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(K)_{\mathbb{Q}}$. Since
 ${\rm rog}(A\setminus(p\cup\tau(p)))\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$
@@
 1.
-The case $f$ is PCF.
-According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps.
-We conclude that $f$ is defined over $\overline{\mathbb{Q}}$ since it is non-exceptional, and Theorem 1.4 was already proved in §4.
-2. The case $f$ is not PCF. Then Theorem 1.4 is a direct consequence of Theorem 5.7.
-We finish the proof of Theorem 1.4.∎
+The case where $f$ is PCF.
+According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ within the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps.
+Since $f$ is non-exceptional, it is defined over $\overline{\mathbb{Q}}$, and Theorem 1.4 was already proved in §4.
+2. The case where $f$ is not PCF. Then Theorem 1.4 is a direct consequence of Theorem 5.7.
+This completes the proof of Theorem 1.4.∎
 7. Proofs of the Applications
 7.1. Proof of Theorem 1.12
-Recall that our goal is to prove the Zariski-dense orbit conjecture for $f=f_{1}\times\cdots\times f_{N}:X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}\to X$ where $f_{1},\dots,f_{N}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ are non-constant endomorphisms on the projective line over an algebraically closed field $k$ of characteristic 00.
+Recall that our goal is to prove the Zariski-dense orbit conjecture for $f=f_{1}\times\cdots\times f_{N}:X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}\to X$, where $f_{1},\dots,f_{N}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ are non-constant endomorphisms on the projective line over an algebraically closed field $k$ of characteristic 0.
 Without loss of generality, we may assume that $k$ is of finite transcendence degree over $\mathbb{Q}$.
@@
 We view $f$ as an endomorphism on $X$ defined over $\mathbb{C}$.
-According to [29, Theorem 3.34], we may assume that all $f_{j}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ has degree at least $2$ for $1\leq j\leq N$.
+According to [29, Theorem 3.34], we may assume that all $f_{j}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ have degree at least $2$ for $1\leq j\leq N$.
 Assume first that all $f_{j}$ are not exceptional, $1\leq j\leq N$.
-Corollary 1.7 implies that we can take $x_{j}\in{\rm Per}^{*}(f_{j})(\mathbb{C})$ for $1\leq j\leq N$ such that the multipliers $\rho_{f_{1}}(x_{1}),\cdots,\rho_{f_{N}}(x_{N})$ are multiplicatively independent in $\mathbb{C}$.
-After replacing $f$ by an iterate, we may assume that $f_{j}(x_{j})=x_{j}$ for $1\leq j\leq N$, and the multipliers $(\rho_{f_{j}}(x_{j})=f_{j}^{\prime}(x_{j}))_{j=1}^{N}$ are still multiplicatively independent.
-Denote $x=(x_{1},\dots,x_{N})\in X(k)$.
+Corollary 1.7 implies that we can take $x_{j}\in{\rm Per}^{*}(f_{j})(\mathbb{C})$ for $1\leq j\leq N$ such that the multipliers $\rho_{f_{1}}(x_{1}),\dots,\rho_{f_{N}}(x_{N})$ are multiplicatively independent in $\mathbb{C}$.
+After replacing $f$ by an iterate, we may assume that $f_{j}(x_{j})=x_{j}$ for $1\leq j\leq N$, and the multipliers $(\rho_{f_{j}}(x_{j})=f_{j}^{\prime}(x_{j}))_{j=1}^{N}$ remain multiplicatively independent.
+Let $x=(x_{1},\dots,x_{N})\in X(k)$.
 Then $x$ is a fixed point of $f$ (smooth in the fixed locus of $f$) such that the eigenvalues of $df|_{x}$ are nonzero and multiplicatively independent.
-Then the conclusion follows from [1].
-Assume that all $f_{j}$ are exceptional, $1\leq j\leq N$. This case is easy, and we just refer to the proof in the first several paragraphs of [29, §9.3].
-We may assume that $0\leq s\leq N$ such that $f_{1},\cdots,f_{s}$ are not exceptional and $f_{s+1},\cdots,f_{N}$ are exceptional.
+The conclusion then follows from [1].
+If all $f_{j}$ are exceptional, $1\leq j\leq N$, this case is relatively easy, and we refer to [29, Theorem 4.1, Theorem 1.14, and Lemma 3.30] (see also [29, §9.3]).
+Now assume that $0\leq s\leq N$ such that $f_{1},\dots,f_{s}$ are not exceptional and $f_{s+1},\dots,f_{N}$ are exceptional.
 Let $l(f)=\min\{s,N-s\}\geq 0$.
-Then we have done in the case $l(f)=0$.
-Then an induction on $l(f)\in\mathbb{Z}_{\geq 0}$ proves Theorem 1.12, as shown in the last several paragraphs of [29, §9.3].∎
+The case $l(f)=0$ has just been treated.
+An induction on $l(f)N\in\mathbb{Z}_{\geq 0}$ then proves Theorem 1.12, as shown in the last several paragraphs of [29, §9.3].∎
 7.2. Proof of Theorem 1.14
 Recall that $f:\mathbb{P}^{1}(\mathbb{C})\to\mathbb{P}^{1}(\mathbb{C})$ is a rational map of degree $\geq 2$.
-We aim to study the condition $(1)$ $f$ is PCF.
+We aim to study the condition: (1) $f$ is PCF.
 Our goal is to prove that (1) is equivalent to each of the conditions (2) and (3) in Theorem 1.14, which are determined by the multiplier spectrum and the length spectrum of $f$, respectively.
@@
 $(2)^{\prime}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x)))$ for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
-$(3)^{\prime}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $\mathbf{n}_{\mathbb{Q}}({\rm rog}(\lvert\rho_{f}(x)\lvert))$ for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
-Now we prove that the three conditions $(1),(2)^{\prime}$, and $(3)^{\prime}$ are equivalent.
+$(3)^{\prime}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$ generated by $)\mathbf{n}_{\mathbb{Q}}({\rm rog}(\lvert\rho_{f}(x)\lvert$)) for $x\in{\rm Per}^{*}(f)(\mathbb{C})$ is of finite dimension over $\mathbb{Q}$.
+We now prove that the three conditions (1), $(2)^{\prime}$, and $(3)^{\prime}$ are equivalent.
 (1) $\Rightarrow(2)^{\prime}$ and $(3)^{\prime}$:
@@
 By [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in $\mathcal{M}_{d}$, except for the family of flexible Lattès maps.
-If $f$ is flexible Lattès, then according to [18, Lemma 5.6], we have $\rho_{f}(x)\in\mathbb{Z}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
+If $f$ is flexible Lattès, then according to [19, Lemma 5.6], we have $\rho_{f}(x)\in\mathbb{Z}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
 If $f$ is defined over $\overline{\mathbb{Q}}$, then clearly $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})={\rm Per}(f)(\mathbb{Q})$.
-Thus, we always have $\rho_{f}(x),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
-Suppose that $(2)^{\prime}$ is false, then
+Thus, in both cases, $\rho_{f}(x),\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all $x\in{\rm Per}(f)(\mathbb{C})$.
+Suppose that $(2)^{\prime}$ is false, i.e.,
 $\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}=\infty.$
-By [18, Corollary 3.9], $f$ cannot be a flexible Lattès map, hence $f$ is defined over $\overline{\mathbb{Q}}$, and over a number field $K$.
-We use the notation and ideas from §6.1 in the case where $\tau=\mathrm{id}$.
-Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all the archimedean places such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$.
-For every $v\in\mathcal{M}_{K}\setminus B$, the reduction $f_{v}$ is still PCF and its critical orbits come from those of $f$.
+By [19, Corollary 3.9], $f$ cannot be a flexible Lattès map, so $f$ is defined over $\overline{\mathbb{Q}}$, and hence over some number field $K$.
+We use the notation and ideas from §6.1 with $\tau=\mathrm{Id}$.
+Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all archimedean places such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$.
+For every $v\in\mathcal{M}_{K}\setminus B$, the reduction $f_{v}$ is still PCF, and its critical orbits come from those of $f$.
 Arguing similarly as in §6.1, it is easy to see that the set
@@
 is co-finite in ${\rm Per}^{*}(f)(\mathbb{C})$, i.e., ${\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}$ is finite.
-It is well-known that ${\rm rank}(\mathcal{O}_{K,B}^{\times})=|B|-1<\infty$ (cf.
+It is well-known that ${\rm rank}(\mathcal{O}_{K,B}^{\times})=(\#B)-1<\infty$ (cf.
 [21, Theorem 3.12]).
@@
 for all $x\in{\rm Per}^{*}(f)(\mathbb{C})$ by [24, Corollary 2.23 (a)]. We deduce that
-$\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{K}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}\leq|{\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W}|+|B|-1<\infty,$
-which implies $\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}<\infty$, contradicting the assumption.
+$\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{K}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}\leq\#({\rm Per}^{*}(f)(\mathbb{C})\setminus\mathcal{W})+(\#B)-1$
+is finite, which implies $\dim_{\mathbb{Q}}{\rm span}_{\mathbb{Q}}\{\mathbf{n}_{\mathbb{Q}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbb{C})\}<\infty$, contradicting the assumption.
 Thus $(2)^{\prime}$ must hold.
-$(3)^{\prime}$ follows from an argument similar to the above paragraph, corresponding to the case where $\tau$ is the complex conjugation of §6.1.
-$(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$:
-Let $\tau$ be the identity on $\mathbb{C}$ in the case $(2)^{\prime}$ and be the complex conjugation in the case $(3)^{\prime}$.
-We will show $(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$ simultaneously.
-Suppose that $f$ is not PCF.
-In particularly, $f$ is not flexible Lattès.
-By $(2)^{\prime}$ (or $(3)^{\prime}$) and Theorem 1.2, the map $f$ is defined over $\overline{\mathbb{Q}}$, hence over a number field $K$.
+$(3)^{\prime}$ follows from a similar argument corresponding to the case where $\tau$ is complex conjugation in §6.1.
+$(2)^{\prime}$ or $(3)^{\prime}\Rightarrow$ (1):
+Let $\tau$ be the identity on $\mathbb{C}$ in the case $(2)^{\prime}$ and complex conjugation in the case $(3)^{\prime}$.
+We prove $(2)^{\prime}$ $\Rightarrow(1)$ and $(3)^{\prime}$ $\Rightarrow(1)$ simultaneously.
+Suppose $f$ is not PCF.
+In particular, $f$ is not a flexible Lattès map.
+By $(2)^{\prime}$ (or $(3)^{\prime}$) and Theorem 1.2, the map $f$ is defined over $\overline{\mathbb{Q}}$, hence over some number field $K$.
 We use the notation and ideas from §6.1 for $\tau$.
-Arguing similarly as in §6.1, after enlarging the number field $K$, we can take a finite set $B\subseteq\mathcal{M}_{K}$ of places containing all the archimedean places satisfying the following conditions:
-  • $K/\mathbb{Q}$ is Galois and $\tau(K)=K$, hence $\tau|_{K}\in{\rm Gal}(K/\mathbb{Q})$;
+Arguing similarly as in §6.1, after enlarging the number field $K$, we can choose a finite set $B\subseteq\mathcal{M}_{K}$ of places containing all archimedean places satisfying the following conditions:
+  • $K/\mathbb{Q}$ is Galois and $\tau(K)=K$, so $\tau|_{K}\in{\rm Gal}(K/\mathbb{Q})$;
   • $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$;
@@
 (Here $o$ and all points in $X$ are assumed to be defined over $K$.)
-A small modification of the proof of Lemma 6.2 shows that there exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ satisfying the following conditions:
+A slight modification of the proof of Lemma 6.2 shows that there exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ satisfying:
 (7.15) $\phi_{v_{i}}(\mathbf{n}_{K}(\lambda(x_{i})))>0\text{ for all }i\geq 1;$
 (7.16) $\phi_{\sigma(v_{j})}(\mathbf{n}_{K}(\lambda(x_{i})))=0\text{ for all }i\neq j\text{ and }\sigma\in{\rm Gal}(K/\mathbb{Q}).$
-For $i\geq 1$, let $p_{i}$ be the rational prime below $v_{i}$ and $\widetilde{B}\subseteq\mathcal{M}_{\mathbb{Q}}$ be the set of places of $\mathbb{Q}$ lying below the places in $B$.
-From (7.15) and (7.16), it is easy to see that the pair $((\mathbf{n}_{\mathbb{Q}}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{p_{i}})_{i=1}^{\infty})$ is an upper triangle $(\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$.
+For $i\geq 1$, let $p_{i}$ be the rational prime below $v_{i}$ and $\widetilde{B}\subseteq\mathcal{M}_{\mathbb{Q}}$ the set of places of $\mathbb{Q}$ lying below the places in $B$.
+From (7.15) and (7.16), it is easy to see that the pair $((\mathbf{n}_{\mathbb{Q}}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{p_{i}})_{i=1}^{\infty})$ is an upper triangular $(\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0})$-system for $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$.
 Note that $\mathbb{D}(\mathcal{O}_{\mathbb{Q},\widetilde{B}})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ is $\tau$-invariant, where $\tau$ descends to $\mathbb{D}(\mathbb{Q})$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(\mathbb{Q})_{\mathbb{Q}}$.
@@
 Thus, $f$ must be PCF.
+∎
 Acknowledgement
@@
 The first-named author would like to thank Beijing International Center for Mathematical Research in Peking University for the invitation.
-The second and third-named authors Junyi Xie and Geng-Rui Zhang are supported by NSFC Grant (No.12271007).
+The first-named author, Zhuchao Ji, is supported by National Key R&D Program of China (No.2025YFA1018300), NSFC Grant (No.12401106), and ZPNSF grant (No.XHD24A0201).
+The second- and third-named authors, Junyi Xie and Geng-Rui Zhang, are supported by NSFC Grant (No.12271007).
 Declarations